### v1

# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

Abstract
Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can be used as a prior for guiding the reconstruction of an undersampled subsequent contrast. To this end, several learning-based guided reconstruction methods have been proposed. However, two key challenges remain – (a) the requirement of large paired training datasets and (b) the lack of intuitive understanding of the model’s internal representation and utilization of the shared information. We propose a modular two-stage approach for guided reconstruction, addressing these challenges. A content/style model of two-contrast image data is learned in a largely unpaired manner and is subsequently applied as a plug-and-play operator in iterative reconstruction. The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors. Based on this, incorporating prior information into the reconstruction reduces to simply replacing the aliased reconstruction content with clean content derived from the reference scan. We name this novel approach PnP-MUNIT. Various aspects like interpretability and convergence are explored via simulations. Furthermore, its practicality is demonstrated on the NYU fastMRI DICOM dataset and two in-house raw datasets, obtaining up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM. In a radiological task, PnP-MUNIT allowed 33.3% more acceleration over clinical reconstruction at diagnostic quality.
I Introduction
Magnetic resonance imaging (MRI) is an invaluable medical imaging modality due to the high-quality scans it delivers, the variety of complementary information it can capture, and its lack of radiation-related risks, leading to its wide usage in clinical practice. However, its central limitation is the inherently slow data acquisition process. The raw sensor data is acquired in the frequency domain, or k-space, from which the image is reconstructed. Over the last 25 years, advancements such as parallel imaging [1, 2], compressed sensing (CS) [3], and deep learning reconstruction [4] have enabled considerable speedups by allowing sub-Nyquist k-space sampling and relying on computationally sophisticated reconstruction. These techniques have subsequently been implemented on commercial MRI systems and have demonstrated (potential) improvements in clinical workflow [5].
A clinical MRI session typically involves the acquisition of multiple scans of the same anatomy through the application of different MR pulse sequences and thus exhibiting different contrasts. Because these scans are different reflections of the same underlying reality, they share a high degree of shared structure. However, currently deployed clinical protocols acquire and reconstruct each scan as an independent measurement, not leveraging the information redundancy across scans. There is, therefore, an opportunity to further optimize MRI sessions by exploiting this shared information. On the reconstruction side, multi-contrast methods have addressed this problem by introducing the shared information into the reconstruction phase to allow higher levels of k-space undersampling. In the simplest case of two contrasts, multi-contrast reconstruction can be classified into two types – (a) guided reconstruction, where an existing high-quality reference scan is used to guide the reconstruction of an undersampled second scan [6, 7, 8] and (b) joint reconstruction, where both contrasts are undersampled and are reconstructed simultaneously [9, 10, 11]. In this work, we consider the problem of guided reconstruction, assuming no inter-scan motion between the reference scan and the reconstruction.
Figure 1: Our two-stage approach to guided reconstruction. (a) The first stage learns a content/style model of two-contrast MR image data in a largely unpaired manner. (b) The reconstruction stage applies this model as a content restoration operator (bottom) within an ISTA-based iterative algorithm, namely PnP-MUNIT. Guidance from the reference is introduced into the reconstruction by simply replacing its aliased content with clean content $c$ derived from the reference. CR denotes a content refinement update, which iteratively corrects for inconsistencies between the reference content and the measured k-space data, improving the effectiveness of the content restoration operator. DC denotes data consistency. Wavelet-domain soft-thresholding (top) and CNN-based denoising (middle) used in CS and PnP reconstructions, respectively, are shown for comparison.
The guided reconstruction problem entails using the spatial structure of the reference scan as a prior to complement the undersampled k-space measurements of the target scan. This problem has been formulated in different ways, ranging from conventional CS [6, 7] to end-to-end learning with unrolled networks [8, 12, 13] and, more recently, diffusion model-based Bayesian maximum a posteriori estimation [14]. Most learning-based approaches, although more powerful than earlier hand-crafted ones, suffer from two main limitations: (a) requirement of large paired training dataset, limiting the effective usage of real-world clinical MR data, and (b) lack of insight into the representation and utilization of the shared underlying structure. We address both issues, leveraging ideas from image-to-image modeling, content/style decomposition, and plug-and-play reconstruction. To the best of our knowledge, ours is the first work that deeply explores content/style modeling for multi-contrast MRI reconstruction.
In recent years, image-to-image translation has found application in the direct estimation of one MR contrast from another [15, 16, 17]. Viewed in the light of MR physics, cross-contrast translation takes an extreme stance by completely ignoring fundamental differences in contrast-specific sensitivities and relying solely on the prior contrast. Cross-contrast translation, by itself, is not a practically reliable solution because data acquisition is necessary for obtaining new information about the anatomy. That said, literature on unpaired image translation provides a repertoire of useful tools such as joint generative modeling of two-domain image data [18, 19], which can be adopted to complement image reconstruction. We observe that methods such as MUNIT [19] can be applied to learn semantically meaningful representations of contrast-independent and contrast-specific information as content and style, respectively, without the need for spatially aligned paired training data.
Plug-and-play (PnP) methods are an emerging paradigm for solving inverse problems in computational imaging. The main research line [20, 21] has focused on learning CNN-based denoising models and applying them as functions replacing proximal operators in iterative algorithms like ISTA and ADMM, demonstrating improved image recovery. An advantage of this approach is the decoupling of the learning problem of image modeling from the inverse problem of image reconstruction, thereby simplifying model training and improving generalizability across different acceleration factors, undersampling patterns, etc. With this design pattern in mind, we combine content/style image modeling and iterative reconstruction in a PnP-like framework.
We propose a two-stage approach for guided reconstruction. We first leverage semantic content/style modeling to learn explicit representations of contrast-independent and contrast-specific components from two-contrast MR image data. This learned model is subsequently embedded into an iterative reconstruction framework resulting in a content/style-based PnP-like algorithm [22]. An overview of our approach is shown in Fig. 1. Specifically, our contributions are four-fold:
  1. 1.
    We show that unpaired training can be used to disentangle contrast-independent and contrast-specific information as content and style, respectively, followed by a fine-tuning strategy to refine the content representation using a modest amount of paired data.
  2. 2.
    With this content/style model as the basis, we define a content restoration operator capable of removing severe undersampling artifacts from the reconstruction image given the corresponding reference image.
  3. 3.
    Developing this idea further, we propose PnP-MUNIT, a modular algorithm for guided reconstruction, combining the flexibility of the plug-and-play approach with the semantic interpretability of content/style decomposition.
  4. 4.
    Through comprehensive experimentation, we shed light on several properties of PnP-MUNIT such as interpretability and convergence as well as demonstrating its applicability on real-world raw data and its potential clinical utility for a specific radiological task.
II Related Work
II-A Reconstruction Methods for Accelerated MRI
Techniques for accelerating MRI by k-space undersampling go back to compressed sensing (CS) [3], which combines random sampling with sparsity-based iterative denoising, most commonly implemented based on the ISTA [23] or ADMM [24] family of algorithms. Most modern deep learning-based reconstruction methods focus on improving the denoising part. Plug-and-play (PnP) methods [20, 21] replace the proximal operator in ISTA and ADMM with off-the-shelf denoisers such as a learned convolutional denoising model. Unrolled networks [25, 26, 4] extend this idea by casting the entire iterative algorithm into one large network, trained end-to-end. This makes them more adaptive to factors such as sampling pattern and acceleration, although at a cost of generalizability [20].
One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke [6] introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on edge features in the reference scan. On the other hand, unrolled network-based guided methods [8, 12, 27, 28] supply the reference scan as an additional input to the reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proved to be more effective than conventional algorithms. However, they suffer from two issues limiting their practicality. First, they require large paired training datasets, which is too strong a constraint when working with retrospectively collected clinical data which reflects the natural inconsistencies of routine practice such as missing contrasts, differences in spatial resolutions, and the existence of inter-scan motion. Second, most of these methods neither explicitly represent the prior information nor provide a clear mechanism by which it is incorporated into the reconstruction. On the other hand, a conventional CS approach like STV precisely defines the prior information and the guidance mechanism. A reconstruction method combining this form of transparency with the effectiveness of deep learning-based methods, while maintaining lenient data requirement, is still needed.
II-B Unpaired Image-to-Image Modeling
Image-to-image modeling is the general problem of learning a mapping between two image domains and was first addressed by Pix2Pix [29] and CycleGAN [30] in paired and unpaired settings, respectively. Another line of unpaired image translation methods, the first of which was UNIT [18], assumes a shared latent space underlying the two domains to explicitly represent shared information, thereby making it more interpretable than CycleGAN which directly models the image-to-image mapping. However, both assume a deterministic one-to-one mapping between the two domains, ignoring the fact that an image in one domain can have multiple valid renderings in the other. Deterministic image translation has been widely applied to MRI contrast-to-contrast synthesis [15, 31, 16]. However, they fundamentally do not account for the variability of the scanning setup which influence the realized contrast level and the differential visibility of pathologies in the target image. Denck et al. [17] partially address this problem by proposing contrast-aware MR image translation where the acquisition sequence parameters are fed into the network to control the output’s contrast level. However, this model is too restrictive since it assumes a single pre-defined mode of variability (i.e. global contrast level) in the data, for which the labels (i.e. sequence parameters) must be available.
MUNIT [19] extended UNIT by modeling domain-specific variability in addition to domain-independent structure, overcoming the deterministic mapping limitation of UNIT. The result was a stochastic image translation model which, given an input image, generates a distribution of synthetic images sharing the same “content” but differing in “style”. Clinical MR images of a given protocol contain multiple modes of variability, many of which are not known a priori. We argue that MUNIT, due to the minimal assumptions it makes, can be generally applied to learn domain-independent anatomical structure and domain-specific variability from an unpaired multi-contrast dataset as content and style, respectively.
II-C Combining Image Translation with Reconstruction
Acknowledging the fundamental limitations of MR contrast prediction, some prior work has attempted combining it with multi-contrast reconstruction. A naive form of guided reconstruction involves generating a synthetic image from the reference scan via deterministic image translation and using it in L2-regularized classical reconstruction [32]. More recent work by Xuan et al. [33] proposes a joint image translation and reconstruction method based on optimal transport which additionally accounts for misalignment between the reference and reconstruction images. Levac et al. [14] formulate guided reconstruction as a highly general Bayesian maximum a posteriori estimation problem and solve it iteratively via Langevin update steps, using an image-domain diffusion model as the score function of the prior distribution. While these are promising directions, an alternative general approach leveraging semantic content/style modeling in multi-contrast reconstruction could potentially provide a connection with the well-established MR physics concepts of quantitative tissue maps, acquisition parameters, and physical models. Such an interpretable and unifying approach remains to be found.
III Methods
III-A Reconstruction Problem
III-A1 Undersampled MRI reconstruction
Given a set of $P$ acquired k-space samples $y\in\mathbb{C}^{P}$ and the MRI forward operator $A\in\mathbb{C}^{P\times Q}$, CS reconstruction of the image $x\in\mathbb{C}^{Q}$ with $Q$ voxels is given as
$\min_{x}||Ax-y||_{2}^{2}+\lambda||\Psi x||_{1},$ (1)
where $\Psi$ is some sparsifying transform (e.g. wavelet) and $\lambda$ is the regularization strength. A commonly used algorithm to solve this optimization problem is ISTA, which iteratively applies the following two update steps:
$r^{k}\leftarrow\Psi^{H}\mathrm{soft}(\Psi x^{k-1};\lambda),$ (2)
$x^{k}\leftarrow r^{k}-\eta A^{H}(Ar^{k}-y).$ (3)
Eq. (2) performs soft-thresholding in the transform domain, thereby reducing the incoherent undersampling artifacts in image $x^{k-1}$, whereas (3) enforces soft data consistency on image $r^{k}$ by taking a single gradient descent step over the least-squares term, controlled by step size $\eta$.
III-A2 Plug-and-play denoiser
Plug-and-play methods replace the analytical operation of (2) with off-the-shelf denoisers. A CNN-based denoiser is of special interest as it incorporates a learning-based component into iterative reconstruction. Given a CNN model $f_{\theta}$ trained to remove i.i.d. Gaussian noise from an image, PnP-CNN [20] modifies Eq. (2) to
$r^{k}\leftarrow f_{\theta}(x^{k-1}).$ (4)
III-A3 Plug-and-play content restoration operator
In guided reconstruction, a spatially aligned reference $x^{\mathrm{ref}}_{1}$ is available, which captures the same underlying semantic content as the reconstruction image. Inspired by the PnP design, we cast the problem of incorporating prior information from $x^{\mathrm{ref}}_{1}$ into reconstruction $x_{2}^{k-1}$ as the restoration of its semantic content, which has been corrupted by undersampling. We propose a content restoration operator $g_{M}(.;c)$ such that (2) becomes
$r_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c),$ (5)
where a content/style model $M$ is used to decompose $x_{2}^{k-1}$ into content and style, followed by a replacement of this content with clean content $c$ derived from $x^{\mathrm{ref}}_{1}$ and composing the image $r_{2}^{k}$ from it. Before formally defining this operator and developing the reconstruction algorithm, we discuss the steps required to learn the content/style model $M$.
III-B Content/Style Modeling
Given an image dataset of two MR contrasts we make the general assumption that there exists an underlying contrast-independent spatial structure which, influenced by arbitrary contrast-specific factors, is rendered as the contrast images.
Figure 2: A trained content/style model $M$ and the content restoration operator $g_{M}(.;c)$ based on it. (a) T1W/T2W pair and corresponding content maps. Despite general agreement between their content, there is notable discrepancy. (b) Synthetic T2W images generated from 16 style codes grid-sampled from $\mathcal{S}_{2}$, showing a smooth variation in image contrast and a roughly constant anatomical structure. (c) $g_{M}(.;c)$ is applied on 3 cases where $x_{2}^{*}$ was corrupted with $R$=4 undersampling with different center fractions. Sampling more low-frequency lines leads to more accurate style estimate $\hat{s}_{2}$ and thus better image $x_{2}^{\mathrm{restored}}$. Contours in $\mathcal{S}_{2}$ indicate the mean absolute error of the synthetic images from that region.
III-B1 Unpaired pre-training of MUNIT
We formulate our content/style model based on the MUNIT framework [19]. We define “domain” as the set $\mathcal{X}_{i}$ of images of a certain contrast comprising the dataset, where $i\in\{1,2\}$. “Content” $c\in\mathcal{C}$ is defined here as the underlying contrast-independent structure and is represented as a set of spatial maps, whereas “style” $s_{i}\in\mathcal{S}_{i}$ corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g. global effects of acquisition settings, contrast-specific tissue features, etc, and is thus represented as a low-dimensional vector. MUNIT posits the existence of functions $G_{i}^{*}:\mathcal{C}\times\mathcal{S}_{i}\rightarrow\mathcal{X}_{i}$ and their inverses $E_{i}^{*}=(G_{i}^{*})^{-1}$, and learns them jointly via unpaired training given samples from marginal distributions $p(x_{i})$. In practice, the encoder $E_{i}$ is split into content encoder $E_{i}^{c}$ and style encoder $E_{i}^{s}$. Thus, the content/style model is specified as $M=\{E_{1}^{c},E_{2}^{c},E_{1}^{s},E_{2}^{s},G_{1},G_{2}\}$.
The MUNIT loss function comprises of 4 terms:
$\mathcal{L_{\mathrm{MUNIT}}}=\mathcal{L}_{\mathrm{GAN}}+\alpha_{1}\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}+\alpha_{2}\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}+\alpha_{3}\mathcal{L}_{\mathrm{style}}^{\mathrm{self}},$ (6)
where
$\begin{split}\mathcal{L}_{\mathrm{GAN}}=\>&\mathbb{E}_{x_{1}\sim p(x_{1}),s_{2}\sim q(s_{2})}[(1-D_{2}(G_{2}(E_{1}^{c}(x_{1}),s_{2})))^{2}]+\\ &\mathbb{E}_{x_{2}\sim p(x_{2}),s_{1}\sim q(s_{1})}[(1-D_{1}(G_{1}(E_{2}^{c}(x_{2}),s_{1})))^{2}],\end{split}$ (7)
$\begin{split}\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}=\>&\mathbb{E}_{x_{1}\sim p(x_{1})}[||x_{1}-G_{1}(E_{1}^{c}(x_{1}),E_{1}^{s}(x_{1}))||_{1}]+\\ &\mathbb{E}_{x_{2}\sim p(x_{2})}[||x_{2}-G_{2}(E_{2}^{c}(x_{2}),E_{2}^{s}(x_{2}))||_{1}],\end{split}$ (8)
$\begin{split}\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}=\>&\mathbb{E}_{c_{1}\sim p(c_{1}),s_{2}\sim q(s_{2})}[||c_{1}-E_{1}^{c}(G_{2}(c_{1},s_{2}))||_{1}]+\\ &\mathbb{E}_{c_{2}\sim p(c_{2}),s_{1}\sim q(s_{1})}[||c_{2}-E_{2}^{c}(G_{1}(c_{2},s_{1}))||_{1}],\end{split}$ (9)
$\begin{split}\mathcal{L}_{\mathrm{style}}^{\mathrm{self}}=\>&\mathbb{E}_{c_{1}\sim p(c_{1}),s_{2}\sim q(s_{2})}[||s_{2}-E_{2}^{s}(G_{2}(c_{1},s_{2}))||_{1}]+\\ &\mathbb{E}_{c_{2}\sim p(c_{2}),s_{1}\sim q(s_{1})}[||s_{1}-E_{1}^{s}(G_{1}(c_{2},s_{1}))||_{1}].\end{split}$ (10)
$\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}$ is the autoencoder image recovery loss, which promotes preservation of input information in the latent space. $\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}$ and $\mathcal{L}_{\mathrm{style}}^{\mathrm{self}}$ are the content and style recovery losses, respectively, which specialize parts of the latent code into content and style components. $\mathcal{L}_{\mathrm{GAN}}$ is the adversarial loss, which enables unpaired training by enforcing distribution-level similarity between synthetic and real images via discriminators $D_{1}$ and $D_{2}$. $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are hyperparameters.
III-B2 Paired fine-tuning
While the pre-training stage learns useful content/style representations, the model can be adapted to our task by improving its content preservation. To this end, we propose a paired fine-tuning (PFT) stage, leveraging a modest amount of paired data. Given samples from the joint distribution $p(x_{1},x_{2})$, our fine-tuning objective is given as
$\mathcal{L}_{\mathrm{PFT}}=\mathcal{L}_{\mathrm{GAN}}+\beta_{1}\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}+\beta_{2}\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}+\beta_{3}\mathcal{L}_{\mathrm{content}}^{\mathrm{cross}},$ (11)
where
$\begin{split}\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}=\>&\mathbb{E}_{\{x_{1},x_{2}\}\sim p(x_{1},x_{2})}[\\ &||x_{2}-G_{2}(E_{1}^{c}(x_{1}),E_{2}^{s}(x_{2}))||_{1}+\\ &||x_{1}-G_{1}(E_{2}^{c}(x_{2}),E_{1}^{s}(x_{1}))||_{1}],\end{split}$ (12)
$\begin{split}\mathcal{L}_{\mathrm{content}}^{\mathrm{cross}}=\mathbb{E}_{\{x_{1},x_{2}\}\sim p(x_{1},x_{2})}[||E_{1}^{c}(x_{1})-E_{2}^{c}(x_{2})||_{1}].\end{split}$ (13)
$\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}$ is a pixel-wise image translation loss, which provides image-level supervision, and $\mathcal{L}_{\mathrm{content}}^{\mathrm{cross}}$ is a paired content loss, which penalizes discrepancy between the contents. $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are an additional set of hyperparameters.
III-B3 Network architecture and locality bias
Following the original paper [19], our content encoders $E_{i}^{c}$ consist of an input convolutional layer potentially followed by strided downsampling convolutions, and finally a series of residual blocks. Style encoders $E_{i}^{s}$ consist of input and downsampling convolutions followed by adaptive average pooling and a fully-connected layer that outputs the latent vector. Decoders $G_{i}$ follow a similar structure as the content encoders except in reverse. Style is introduced into the decoder via AdaIN operations which modulate the activation maps derived from the content. We note that in this architecture, the content-size-to-image-size ratio reflects the level of spatial structure one expects to be shared between the two domains. This (relative) content size is thus an inductive bias, known as the locality bias [34], built into the model. A model with high locality bias has a large content size, allowing for a richer content that strongly influences the output’s spatial structure, while simultaneously restricting style to learn low-level global features.
III-C Iterative Reconstruction with Content Consistency
Given the content/style model $M$, we consider $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ as reference and target domains, respectively and define the content restoration operator $g_{M}(.;c)$ as
$x^{\mathrm{restored}}_{2}=g_{M}(x^{\mathrm{corrupt}}_{2};c_{1}^{\mathrm{ref}}):=G_{2}(c_{1}^{\mathrm{ref}},E^{s}_{2}(x^{\mathrm{corrupt}}_{2})),$ (14)
where $x^{\mathrm{corrupt}}_{2}$ is the reconstruction corrupted with undersampling artifacts, $x^{\mathrm{restored}}_{2}$ is the restored image, and content $c_{1}^{\mathrm{ref}}=E^{c}_{1}(x^{\mathrm{ref}}_{1})$. This operation restores $x^{\mathrm{corrupt}}_{2}$ by simply replacing its aliased content with the clean content from $x^{\mathrm{ref}}_{1}$, a rule which will later be softened with (15).
Note that this is a radical operation as it discards all structure contained in $x^{\mathrm{corrupt}}_{2}$, retaining only a compact style code $\hat{s}_{2}=E_{2}^{s}(x^{\mathrm{corrupt}}_{2})$. Let $x_{2}^{*}$ be the ground truth reconstruction with content $c_{2}^{*}$ and style $s_{2}^{*}$. The success of the content restoration operation depends on two conditions – (a) $c_{1}^{\mathrm{ref}}$ is close to $c_{2}^{*}$ and (b) $\hat{s}_{2}$ is close to $s_{2}^{*}$. The first condition is roughly satisfied because the model is explicitly trained to minimize content discrepancy, which can be seen in Fig. 2a, but more on this later. On the other hand, it is not obvious that the second condition should hold too and hence deserves a closer look. Assuming a high degree of shared spatial structure between the two domains, the optimal model has high locality bias (based on Section III-B3). Thus, style represents low-level global image features, e.g. contrast variations, as observed in Fig. 2b. It is a well known fact that image contrast is contained prominently in the center of the k-space. Hence, the estimate $\hat{s}_{2}$ can be made arbitrarily close to $s_{2}^{*}$ by sufficiently sampling the k-space center, as shown empirically in Fig. 2c.
Algorithm 1 PnP-MUNIT
1:$y$, $A$, $x^{\mathrm{ref}}_{1}$, $M$, $\eta$, $\gamma$triangleright where $M=\{E_{1}^{c},E_{1}^{s},E_{2}^{c},E_{2}^{s},G_{1},G_{2}\}$
2:$k\leftarrow 0$
3:$x_{2}^{k}\leftarrow A^{H}y$
4:$c^{k}\leftarrow E_{1}^{c}(x^{\mathrm{ref}}_{1})$
5:repeat
6:     $k\leftarrow k+1$
7:     $r_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c^{k-1})$triangleright Content consistency
8:     $x_{2}^{k}\leftarrow r_{2}^{k}-\eta A^{H}(Ar_{2}^{k}-y)$triangleright Data consistency
9:     $c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},E_{2}^{s}(x_{2}^{k}))-y||_{2}^{2}$triangleright CR
10:until convergence
11:return $x_{2}^{k}$
Applying data consistency update (3) following $g_{M}(.;c)$ in repetition yields an iterative scheme, where the latter can be viewed as a hard content consistency update. Fig. 2a shows that although $c_{1}^{\mathrm{ref}}$ and $c_{2}^{*}$ are roughly similar, there is notable discrepancy between them. There are at least two possible sources of content discrepancy, which make this phenomenon inevitable – (a) model-related, e.g. practical issues like sub-optimal training or fundamental limits such as irreducible error, and (b) reference image-related, e.g. presence of artifacts independent of the reconstruction. During reconstruction, a $c_{1}^{\mathrm{ref}}$ that conflicts with $c_{2}^{*}$, and with k-space $y$ by extension, would hamper convergence. While model-related discrepancy is partly tackled by the PFT strategy (Section III-B2), we propose a content refinement (CR) step to correct for the remaining discrepancy in the reconstruction stage. The content state is initialized with $c_{1}^{\mathrm{ref}}$ and gradually refined to be consistent with the acquired k-space $y$ as
$c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},E_{2}^{s}(x_{2}^{k}))-y||_{2}^{2},$ (15)
where $\gamma$ is the step size. This update equation mirrors the form of the data consistency update (3), except here the gradient is taken with respect to content through the (non-linear) augmented forward operator $AG_{2}(.)$, which relates the latent content directly to the measured k-space. Hence, by aligning content with k-space explicitly, CR aligns content consistency updates with data consistency updates, correcting for any conflicts between them caused due to content discrepancy. We refer to the resulting algorithm as PnP-MUNIT (Algorithm 1).
IV Experiments
Without loss of generality, we consider the special case of reconstructing T2W scans using T1W references and focus on head applications. In IV-A, we present an analysis of locality bias, convergence, and tuning using BrainWeb-based simulated MR datasets [35]. In IV-B, we benchmark against ablations and closely related methods on NYU fastMRI brain dataset [36, 37] and two in-house multi-coil raw datasets of the brain. Finally, in IV-C, we present a radiological evaluation comparing single- and multi-contrast reconstructions. All software was implemented in PyTorch, and training was performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
IV-A Experiments on Simulated MR Dataset
BrainWeb provides 20 anatomical models of normal brain, each comprised of fuzzy segmentation maps of 12 tissue types. The 20 volumes were first split in 18:1:1 ratio for model training, validation, and reconstruction testing, respectively. T1W/T2W spin-echo scans were simulated using TE/TR values randomly sampled from realistic ranges. For testing, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling.
IV-A1 Interpretation of content size
TABLE I: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices) at $R$=4. Optimal locality bias is shown in bold. PSNR worse than L1-wavelet CS (16.83$\pm$0.07) is marked as .
Model Config Data Config
RefSize-1 RefSize-2 RefSize-4 RefSize-8
ContentSize-1 27.90$\pm$0.08 21.17$\pm$0.05 19.11$\pm$0.07 16.43$\pm$0.07
ContentSize-2 26.19$\pm$0.08 22.80$\pm$0.08 20.05$\pm$0.08 16.57$\pm$0.08
ContentSize-4 23.98$\pm$0.08 20.88$\pm$0.10 19.44$\pm$0.09 16.88$\pm$0.06
The goal of this experiment was to (a) determine the relationship between the locality bias of the content/style model and the relative spatial resolution of the dataset and (b) explain the effect of reducing the reference image resolution on the reconstruction in terms of locality bias. We simulated 4 datasets, denoted as RefSize-$n$, where $n\in\{1,2,4,8\}$ represents the reference domain downsampling factor. In the $n$=1 case, T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower $1/n$ frequency components. For each dataset, we trained 3 models (for 200k iterations) with locality bias denoted as ContentSize-$m$, where $m\in\{1,2,4\}$ is the content downsampling factor. Table I compares PnP-MUNIT reconstruction quality across these configurations. PFT and CR were disabled here for simplicity. We observe that (a) the optimal locality bias decreases with reference resolution and (b) the reconstruction quality decreases as well, eventually dropping below non-guided reconstruction. Both are explained by the fact that with decrease in reference contrast resolution, the amount of effective content in the dataset decreases, thereby limiting the value of guided reconstruction.
IV-A2 Convergence
Figure 3: (a) Convergence of PnP-MUNIT and the baseline versions at $R$=4. (b) Evolution of the reconstruction shown with its error map.
We explore here the convergence of PnP-MUNIT by comparing it with two variants. The first uses the content $c_{2}^{*}$ of the ground truth T2W image, thus assuming zero content discrepancy and representing an upper-bound of PnP-MUNIT. The second uses the reference content $c_{1}^{\mathrm{ref}}$ but with CR step disabled, representing a lower-bound of PnP-MUNIT where the non-zero content discrepancy is not corrected. Here as well as in the following simulation experiment, we used the RefSize-1 dataset and the MUNIT-1 model additionally fine-tuned on 2 training volumes for 50k iterations. Fig. 3 shows the convergence curves and intermediate reconstructions. The upper-bound version, given $c_{2}^{*}$, converged in a single iteration, while with $c_{1}^{\mathrm{ref}}$, the ablated version converged equally fast but to a sub-optimal solution. Enabling the CR step closed the gap with the upper-bound, although costing convergence rate.
Figure 4: Reconstruction PSNR as a function of acceleration $R$ and k-space noise level $\sigma$ for different values of the $\gamma$ parameter. $\gamma$ controls the consistency of the content with the measured k-space.
IV-A3 Tuning the CR step size
The hyperparameter $\gamma$ in Algorithm 1 controls the CR strength. Fig. 4 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal $\gamma$ decreases with the amount and quality of k-space data. This makes sense as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
Figure 5: Evaluation plots over recon-test subset of NYU, LUMC-TRA, and LUMC-COR datasets, consisting of 550, 1366, and 200 slices, respectively.
IV-B Benchmark and Ablation
For benchmarking, we used L1-wavelet CS (CS-WT) and PnP-CNN [20] as our single-contrast baselines. The denoising model of PnP-CNN was defined as a 5-layer network with convolution and ReLU operations and a residual connection after the final layer. We additionally compared against deterministic image translation based on MUNIT defined as $x_{2}^{\mathrm{syn}}:=\mathbb{E}_{s_{2}\sim q(s_{2})}[G_{2}(E_{1}^{c}(x^{\mathrm{ref}}_{1}),s_{2})]$. As guided reconstruction baselines, we used STV-based CS (CS-STV) [6] and image translation combined with L2-regularized reconstruction (Translation + L2 recon) [32] given as $\mathop{\min}_{x_{2}}||Ax_{2}-y||_{2}^{2}+\lambda||x_{2}-x_{2}^{\mathrm{syn}}||_{2}^{2}$. For PnP-MUNIT, we ablated PFT and CR to assess their contribution, and finally, as an upper-bound representing zero content discrepancy, we disabled PFT and CR and used the ideal content $c_{2}^{*}$.
TABLE II: Data split for benchmarking. Splits were performed at subject-level, each subject belonging to exactly one subset.
Dataset Split Subjects Sessions Scans (T1W / T2W)
NYU model-train 200 200 200 / 200
model-val 27 27 27 / 27
recon-val 50 50 50 / 50
recon-test 50 50 50 / 50
Total 327 327 327 / 327
LUMC-TRA model-train 295 418 360 / 415
model-val 16 17 17 / 17
recon-val 18 21 21 / 21
recon-test 20 31 31 / 31
Total 339 487 429 / 484
LUMC-COR model-train 242 277 269 / 272
model-val 18 18 18 / 18
recon-val 15 18 18 / 18
recon-test 16 17 17 / 17
Total 291 330 322 / 325
The NYU brain dataset includes raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR. Following [33], a DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans. Single-coil T2W k-space data was simulated via Fourier transform with 1D Cartesian random sampling of $R\in\{2,3,4,5\}$ and added Gaussian noise of $\sigma=0.01\mathrm{max}(x_{2}^{*})$.
Our in-house datasets consisted of brain scans of patients from LUMC, the use of which was approved for research purpose by the institutional review board. A total of 1669 brain scans were obtained from 817 clinical MR examinations of 630 patients acquired on 3T Philips Ingenia scanners. We focused on accelerating two T2W sequences – (a) 2D T2W TSE transversal (FA=90°, TR=4000-5000ms, TE=80-100ms, ETL=14-18, median voxel size=0.37$\times$0.5mm2, median FOV=238$\times$190mm2, median slice thickness=3mm, median slices=50), and (b) 2D T2W TSE coronal (FA=90°, TR=2000-3500ms, TE=90-100ms, ETL=17-19, median voxel size=0.39$\times$0.47mm2, median FOV=130$\times$197mm2, median slice thickness=2mm, median slices=15). For guidance, two corresponding T1W sequences were used – (a) 3D T1W TFE transversal (FA=8°, TR=9.8-9.9ms, TE=4.6ms, ETL=200, median voxel size=0.98$\times$0.99$\times$0.91mm3, median FOV=238$\times$191$\times$218mm3), and (b) 3D T1W TSE coronal (FA=80-90°, TR=500-800ms, TE=6.5-16ms, ETL=10-13, median voxel size=0.59$\times$0.62$\times$1.19mm3, median FOV=130$\times$238$\times$38mm3). The transversal and coronal protocols were considered as two separate datasets, namely LUMC-TRA and LUMC-COR. The multi-coil T2W raw data comprised of 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps. This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of $R$=1.8-2. We further undersampled it retrospectively to higher accelerations $R\in\{4,6,8,10\}$ by dropping subsets of the acquired lines.
Each dataset was split into 4 subsets – (a) model-train, (b) model-val, (c) recon-val, and (d) recon-test. The former two were used to train and tune the plug-and-play models, recon-val to tune the reconstruction algorithms, and recon-test as the held-out test set. An overview of the data split is shown in Table II. Additionally, 20 subjects from model-train were designated for PFT. Spatially aligned images required by all except the unpaired training set were obtained via rigid registration using Elastix [38]. We empirically found the ContentSize-1 and ContentSize-2 configurations as optimal for the NYU and LUMC datasets, respectively, which reflects the lower effective content and hence greater difficulty of the LUMC datasets. For each dataset, the model was pre-trained and fine-tuned for 400k and 50k iterations, respectively. We used 3 perceptual metrics for evaluation: SSIM, HaarPSI, and DISTS. While SSIM is commonly used, HaarPSI and DISTS are known to correlate better with visual judgment of image quality [39]. All three metrics are bounded to $[0,1]$ with $1$ representing perfect image quality. We additionally conducted paired Wilcoxon signed-rank tests to measure statistical significance when comparing pairs of algorithms.
Fig. 5 plots the benchmark metrics, which can be summarized in the following three trends. First, pure image translation was worse compared to single-contrast reconstruction, especially at lower acceleration factors (comparing with PnP-CNN, $p$$<$0.05 for all metrics and datasets). Combining it with L2-regularized reconstruction improved SSIM and HaarPSI ($p$$<$0.05 for all datasets and accelerations except SSIM at $R$=3 in NYU), but not necessarily DISTS, suggesting that the the available complementary information was not fully utilized. PnP-MUNIT was consistently better than both image translation and naive guidance ($p$$<$0.05 throughout for both cases). It also outperformed the conventional guided CS-STV ($p$$<$0.05 throughout, except SSIM at $R$=4 in LUMC-COR). Second, both PnP-CNN and PnP-MUNIT performed similarly at lower acceleration. At higher acceleration, PnP-MUNIT outperformed PnP-CNN ($p$$<$0.05 for all metrics and datasets except SSIM at $R$=8 in LUMC-COR). Third, introducing PFT and CR into the ablated PnP-MUNIT improved the reconstructions ($p$$<$0.05 throughout in both cases except in DISTS for the latter at $R$=10 in LUMC-TRA). The contribution of PFT was almost constant across $R$, whereas that of CR decreased with $R$, which was expected since CR depends on the measured k-space data to refine the content.
Figure 6: Sample slice from the LUMC-COR recon-test set comparing PnP-MUNIT with PnP-CNN and pure image translation.
Fig. 6 shows an example from LUMC-COR representative of the first two trends. Image translation produced severe anatomical defects and predicted structure in the lower region of the image where the ground truth contains low signal. At $R$=4, PnP-MUNIT was visually similar to PnP-CNN, with the main difference being a blur effect in PnP-CNN and mild texture artifacts in PnP-MUNIT. At $R$=10, PnP-CNN reconstruction was non-viable, containing severe blur and artifacts (green arrows). On the other hand, PnP-MUNIT did not degrade much from the $R$=4 case and most fine structures were sharply resolved (green arrows). However, an interesting failure mode of PnP-MUNIT at high acceleration was the subtle localized distortions in the anatomy (pink arrows).
IV-C Radiological Evaluation
Figure 7: Averaged results of the radiological evaluation.
We conducted the radiological evaluation on a small sample of LUMC-TRA recon-test set. We selected 3 cases that showed brain metastases in corresponding gadolinium-enhanced T1W scans. The evaluation comprised two parts, namely visual quality and pathology. Four visual quality criteria were used, namely sharpness, noise, artifacts, and contrast between gray and white matter and CSF. For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases. The images were scored using a five-point Likert scale [40] – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality. We considered PnP-MUNIT and PnP-CNN at 4 clinically realistic accelerations $R\in\{3,4,5,6\}$. In total, 9 images per patient, including the reconstructions and the clinical ground truth of $R$=2, were presented to a senior neuroradiologist who scored each image individually, blinded to the reconstruction method and $R$. Figure 7 plots the evaluation result. At $R$=3, both algorithms produced at least “fairly diagnostic” reconstructions and allowed the detection of more than 90% of metastases. PnP-CNN was better than PnP-MUNIT at $R\in\{3,4\}$ in terms of sharpness, contrast, and the pathology criteria. At $R$=5, PnP-MUNIT matched or exceeded PnP-CNN in terms of visual quality, although being worse in terms of pathology criteria. At $R$=6, PnP-CNN sharply dropped to non-diagnostic level, whereas PnP-MUNIT was better, especially in terms of sharpness and contrast as well as in overall diagnostic quality.
V Discussion
We explored the interpretability aspect of PnP-MUNIT in terms of the effective content in a two-contrast dataset (Section IV-A1) and obtained a notion of limit to guided reconstruction in terms of the reference image resolution. This limit would indeed depend on more fundamental factors such as MR sequence types, which we did not investigate here. Our data-driven definition of content and style is loosely analogous to MR physics-based descriptions of quantitative tissue maps and acquisition-related factors. Hence, a further level of interpretability could be achieved by constraining the learned content to represent physically meaningful tissue parameters. Content discrepancy is characterized as a limiting factor in guided reconstruction (Section III-C) and plays a major role throughout this work. Viewed as a set of quasi tissue maps, estimating content from a single image is challenging, ideally requiring a time series of images, which captures the magnetization dynamics. Content discrepancy is, therefore, inevitable. We proposed two complementary ways to minimize its effect on the reconstruction. While paired fine-tuning improves the model using a small amount of paired data, content refinement (CR) corrects for the remaining discrepancy adaptively during reconstruction using the measured k-space. A drawback of CR, however, is slower convergence (Section IV-A2). This could be mitigated by partially adopting an unrolled network design. The resulting tradeoff between model latency and generalizability would be interesting to explore.
Unpaired pre-training of the content/style model boosts the practical applicability of PnP-MUNIT, e.g. in the case of LUMC-TRA where the T1W/T2W data imbalance is considerable. In this work, we limited our baselines to those algorithms that closely resemble ours in terms of design or data requirement. As future work, a more comprehensive study could be conducted with a broader range of methods and evaluation criteria, e.g. comparing generalizability across accelerations and sampling patterns with guided unrolled networks, given a fixed budget of paired training data. At reconstruction time, we assumed the reference and reconstruction images to be spatially aligned. The CR update, which can correct minor registration errors implicitly, would likely break down at the typically observed levels of patient motion. A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further relaxing the overall data constraints.
In terms of reconstruction quality, we observed (Section IV-B) that using both measured k-space and reference scan via PnP-MUNIT was more beneficial than (a) using either one of them and (b) combining both in a naive way or based on hand-crafted prior, suggesting that our approach maximally exploits the complementary information. Compared to PnP-CNN, PnP-MUNIT allowed up to 32.6% more acceleration for a given SSIM. [^1] Based on linear interpolation of LUMC-TRA SSIM, PnP-MUNIT and PnP-CNN allowed $R$=10 and $R$=6.7, respectively, the maximum difference in $R$, at median SSIM=0.906. Hence, the maximum time saving 32.6%. In the radiological task (Section IV-C), PnP-MUNIT produced diagnostic-quality reconstructions at $R$=3, a difference of 33.3% over the clinical reconstructions of $R$=2.222We define diagnostic quality as visibility of at least 90% of the metastases and scoring at least “fairly diagnostic quality” on all other criteria. In clinical practice, while 10% of the metastases may not be detected in the reconstructed T2W scan, they would be prominently visible in the corresponding gadolinium-enhanced T1W scan, and hence would not go undetected. However, it was somewhat worse than PnP-CNN, while theoretically, the additional guidance should result in at least equivalent image quality. We argue that the likely cause is the limitations of the dataset and the specific content/style model, not of the proposed general approach. First, in LUMC-TRA, T1W scans had 2-3 times lower resolution than T2W scans, limiting the value of guidance (Section IV-A1). In the long run, this may be solved by co-designing a higher-resolution reference sequence with the target sequence to optimize guidance. Second, and more importantly, correcting for content discrepancy using CR relies on computing gradients through the decoder $G_{2}$. If $G_{2}$ is highly non-linear and the reference content is not sufficiently close to the optimum, CR may converge to a sub-optimal point leading to texture artifacts in the reconstruction, as seen in Fig. 6 at $R$=4. In the radiological study, precisely this phenomenon is likely to have impacted the visual and diagnostic quality of PnP-MUNIT. Hence, to improve clinical utility, the content/style model must be improved using specialized architecture and regularization.
VI Conclusion
In this work, we introduced PnP-MUNIT, a novel approach to multi-contrast reconstruction combining content/style modeling with iterative reconstruction, which is interpretable in that it allows the two stages to be analyzed separately. By exploring the interpretability aspect, two limiting factors of guided reconstruction were identified – the amount of effective content in a two-contrast dataset and content discrepancy. On real-world clinical data, PnP-MUNIT provided up to 32.6% more acceleration over PnP-CNN for a given SSIM, enabling sharper reconstructions at high accelerations. Given the radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-MUNIT produced diagnostic-quality images at $R$=3, enabling 33.3% more acceleration over clinical reconstructions. To progress towards a practical implementation, future work will focus on improving the content/style model and reconstruction latency and incorporating online registration into the reconstruction.

### v2

@@
 To this end, several learning-based guided reconstruction methods have been proposed.
-However, two key challenges remain – (a) the requirement of large paired training datasets and (b) the lack of intuitive understanding of the model’s internal representation and utilization of the shared information.
-We propose a modular two-stage approach for guided reconstruction, addressing these challenges.
-A content/style model of two-contrast image data is learned in a largely unpaired manner and is subsequently applied as a plug-and-play operator in iterative reconstruction.
+However, a key challenge is the requirement of large paired training datasets comprising raw data and aligned reference images.
+We propose a modular two-stage approach for guided reconstruction addressing this issue, which additionally provides an explanatory framework for the multi-contrast problem in terms of the shared and non-shared generative factors underlying two given contrasts.
+A content/style model of two-contrast image data is learned from a largely unpaired image-domain dataset and is subsequently applied as a plug-and-play operator in iterative reconstruction.
 The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors.
-Based on this, incorporating prior information into the reconstruction reduces to simply replacing the aliased reconstruction content with clean content derived from the reference scan.
+Based on this, incorporating prior information into the reconstruction reduces to simply replacing the aliased content of the image estimate with high-quality content derived from the reference scan.
+Combining this component with a data consistency step and introducing a general corrective process for the content yields an iterative scheme.
 We name this novel approach PnP-MUNIT.
 Various aspects like interpretability and convergence are explored via simulations.
-Furthermore, its practicality is demonstrated on the NYU fastMRI DICOM dataset and two in-house raw datasets, obtaining up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM.
+Furthermore, its practicality is demonstrated on the NYU fastMRI DICOM dataset and two in-house multi-coil raw datasets, obtaining up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM.
 In a radiological task, PnP-MUNIT allowed 33.3% more acceleration over clinical reconstruction at diagnostic quality.
@@
 The raw sensor data is acquired in the frequency domain, or k-space, from which the image is reconstructed.
-Over the last 25 years, advancements such as parallel imaging [1, 2], compressed sensing (CS) [3], and deep learning reconstruction [4] have enabled considerable speedups by allowing sub-Nyquist k-space sampling and relying on computationally sophisticated reconstruction.
+Over the last 25 years, advancements such as parallel imaging [1, 2], compressed sensing (CS) [3], and deep learning reconstruction [4, 5] have enabled considerable speedups by allowing sub-Nyquist k-space sampling and relying on computationally sophisticated reconstruction.
 These techniques have subsequently been implemented on commercial MRI systems and have demonstrated (potential) improvements in clinical workflow [6].
-A clinical MRI session typically involves the acquisition of multiple scans of the same anatomy through the application of different MR pulse sequences and thus exhibiting different contrasts.
+A clinical MRI session typically involves the acquisition of multiple scans of the same anatomy through the application of different MR pulse sequences, exhibiting different contrasts.
 Because these scans are different reflections of the same underlying reality, they share a high degree of shared structure.
@@
 In the simplest case of two contrasts, multi-contrast reconstruction can be classified into two types – (a) guided reconstruction, where an existing high-quality reference scan is used to guide the reconstruction of an undersampled second scan [7, 8, 9] and (b) joint reconstruction, where both contrasts are undersampled and are reconstructed simultaneously [10, 11, 12].
-In this work, we consider the problem of guided reconstruction, assuming no inter-scan motion between the reference scan and the reconstruction.
+In this work, we consider the problem of guided reconstruction, assuming no inter-scan motion between reference and target scans.
 Figure 1: Our two-stage approach to guided reconstruction.
-(a) The first stage learns a content/style model of two-contrast MR image data in a largely unpaired manner.
-(b) The reconstruction stage applies this model as a content restoration operator (bottom) within an ISTA-based iterative algorithm, namely PnP-MUNIT.
-Guidance from the reference is introduced into the reconstruction by simply replacing its aliased content with clean content $c$ derived from the reference.
-CR denotes a content refinement update, which iteratively corrects for inconsistencies between the reference content and the measured k-space data, improving the effectiveness of the content restoration operator.
+(a) The first stage learns a content/style model of two-contrast MR image data.
+The two image domains $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ are decomposed into a shared content domain $\mathcal{C}$ and separate style domains $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$.
+This illustration only provides an intuition about the information encoded in the content and style domains, and does not represent the actual training process, which is given in detail in Section III-B.
+(b) The reconstruction stage applies this model as a content consistency operator (bottom) within an ISTA-based iterative algorithm.
+Given an aligned reference image, guidance is introduced into the reconstruction by simply replacing its aliased content with content derived from the reference.
+CR denotes a content refinement update, which iteratively corrects for inconsistencies between the reference content and the measured k-space data, improving the effectiveness of the content consistency operator.
 DC denotes data consistency.
-Wavelet-domain soft-thresholding (top) and CNN-based denoising (middle) used in CS and PnP reconstructions, respectively, are shown for comparison.
-The guided reconstruction problem entails using the spatial structure of the reference scan as a prior to complement the undersampled k-space measurements of the target scan.
-This problem has been formulated in different ways, ranging from conventional CS [6, 7] to end-to-end learning with unrolled networks [8, 12, 13] and, more recently, diffusion model-based Bayesian maximum a posteriori estimation [14].
-Most learning-based approaches, although more powerful than earlier hand-crafted ones, suffer from two main limitations: (a) requirement of large paired training dataset, limiting the effective usage of real-world clinical MR data, and (b) lack of insight into the representation and utilization of the shared underlying structure.
-We address both issues, leveraging ideas from image-to-image modeling, content/style decomposition, and plug-and-play reconstruction.
-To the best of our knowledge, ours is the first work that deeply explores content/style modeling for multi-contrast MRI reconstruction.
-In recent years, image-to-image translation has found application in the direct estimation of one MR contrast from another [15, 16, 17].
-Viewed in the light of MR physics, cross-contrast translation takes an extreme stance by completely ignoring fundamental differences in contrast-specific sensitivities and relying solely on the prior contrast.
-Cross-contrast translation, by itself, is not a practically reliable solution because data acquisition is necessary for obtaining new information about the anatomy.
-That said, literature on unpaired image translation provides a repertoire of useful tools such as joint generative modeling of two-domain image data [18, 19], which can be adopted to complement image reconstruction.
-We observe that methods such as MUNIT [19] can be applied to learn semantically meaningful representations of contrast-independent and contrast-specific information as content and style, respectively, without the need for spatially aligned paired training data.
+Wavelet-domain soft-thresholding (top) and CNN-based denoising (middle) used in CS-WT and PnP-CNN reconstruction, respectively, are shown for comparison.
+The guided reconstruction problem entails using the local structure of the reference scan as a prior to complement the undersampled k-space measurements of the target scan.
+This problem has been formulated in different ways, ranging from conventional CS [7, 8] to end-to-end learning with unrolled networks [9, 13, 14, 15] and, more recently, diffusion model-based Bayesian maximum a posteriori estimation [16].
+Most end-to-end approaches, although more powerful than earlier hand-crafted ones, suffer from the main drawback of requiring large paired training datasets consisting of the target image and its k-space together with an aligned reference image, thereby limiting their application on real-world MR data.
+We address this issue by proposing a plug-and-play reconstruction method that splits the problem into a purely image-domain learning sub-problem and an iterative reconstruction sub-problem.
+The learning problem leverages ideas from content/style decomposition, thereby offering a degree of interpretability to our approach.
+To the best of our knowledge, ours is the first work that thoroughly explores content/style-based generative modeling for multi-contrast MRI reconstruction.
+In recent years, image-to-image translation has found application in the direct estimation of one MR contrast from another [17, 18, 19, 20].
+While these methods are attractive due to their lack of dependence on k-space data, viewed in the light of MR physics, cross-contrast translation takes an extreme stance by not explicitly taking into account contrast-specific sensitivities and relying solely on the prior contrast.
+Hence, in the context of MR image formation, cross-contrast translation, by itself, can only provide a part of the information about the target image as data acquisition via MR sequences is necessary to obtain new information about the anatomy.
+That being said, literature on unpaired image translation provides a repertoire of useful tools such as joint generative modeling of two-domain image data [21, 22], which can be adopted to complement image reconstruction.
+We observe that methods such as MUNIT [22] can be applied to learn semantically meaningful representations of contrast-independent and contrast-specific information as content and style, respectively, without the need for paired image-domain training data.
 Plug-and-play (PnP) methods are an emerging paradigm for solving inverse problems in computational imaging.
-The main research line [20, 21] has focused on learning CNN-based denoising models and applying them as functions replacing proximal operators in iterative algorithms like ISTA and ADMM, demonstrating improved image recovery.
+The main research line [23, 24] has focused on learning CNN-based denoising models on image-domain data and applying them as functions replacing proximal operators in iterative algorithms like ISTA and ADMM, demonstrating improved image recovery.
 An advantage of this approach is the decoupling of the learning problem of image modeling from the inverse problem of image reconstruction, thereby simplifying model training and improving generalizability across different acceleration factors, undersampling patterns, etc.
 With this design pattern in mind, we combine content/style image modeling and iterative reconstruction in a PnP-like framework.
-We propose a two-stage approach for guided reconstruction.
 We first leverage semantic content/style modeling to learn explicit representations of contrast-independent and contrast-specific components from two-contrast MR image data.
-This learned model is subsequently embedded into an iterative reconstruction framework resulting in a content/style-based PnP-like algorithm [22].
+This training process is independent of the reconstruction problem and can be performed using unpaired images.
+We then make the interesting observation that in multi-contrast MR images, the style information tends to localize in the center of the k-space, allowing for an accurate style estimation from an undersampled image.
+Since the content of the reconstructing image, which is the remaining piece of information and is contrast-independent, is supplied by the reference image, one can compose a de-aliased estimate of the reconstruction from the undersampled image in a single step.
+We term this the content consistency operation, which forms the basis of our iterative reconstruction algorithm PnP-MUNIT [25].
 An overview of our approach is shown in Fig.
 1.
+While the PnP-based decoupling of image modeling and image reconstruction simplifies the training process and allows the two stages to be analyzed separately, a further level of decoupling offered by content/style disentanglement offers additional modularity and an intuitive guidance mechanism for the reconstruction.
 Specifically, our contributions are four-fold:
   1. 1.
-    We show that unpaired training can be used to disentangle contrast-independent and contrast-specific information as content and style, respectively, followed by a fine-tuning strategy to refine the content representation using a modest amount of paired data.
+    We show that unpaired image-domain training can be used to learn disentangled contrast-independent and contrast-specific representations, followed by a fine-tuning strategy that refines the content representation using a modest amount of paired image-domain data.
   2. 2.
-    With this content/style model as the basis, we define a content restoration operator capable of removing severe undersampling artifacts from the reconstruction image given the corresponding reference image.
+    With this content/style model as the basis, we define a content consistency operator capable of removing severe undersampling artifacts from the reconstructing image, given the corresponding reference image.
   3. 3.
-    Developing this idea further, we propose PnP-MUNIT, a modular algorithm for guided reconstruction, combining the flexibility of the plug-and-play approach with the semantic interpretability of content/style decomposition.
+    Developing this idea further and incorporating a corrective process for the content, we propose PnP-MUNIT, a modular algorithm for guided reconstruction, combining the flexibility of the plug-and-play approach with the semantic interpretability of content/style decomposition.
   4. 4.
-    Through comprehensive experimentation, we shed light on several properties of PnP-MUNIT such as interpretability and convergence as well as demonstrating its applicability on real-world raw data and its potential clinical utility for a specific radiological task.
+    Through comprehensive experiments, we shed light on several properties of PnP-MUNIT such as convergence and robustness and demonstrate its applicability on real-world raw data and its potential clinical utility for a specific radiological task.
 II Related Work
@@
 This makes them more adaptive to factors such as sampling pattern and acceleration, although at a cost of generalizability [23].
-One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke [6] introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on edge features in the reference scan.
-On the other hand, unrolled network-based guided methods [8, 12, 27, 28] supply the reference scan as an additional input to the reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proved to be more effective than conventional algorithms.
-However, they suffer from two issues limiting their practicality.
-First, they require large paired training datasets, which is too strong a constraint when working with retrospectively collected clinical data which reflects the natural inconsistencies of routine practice such as missing contrasts, differences in spatial resolutions, and the existence of inter-scan motion.
-Second, most of these methods neither explicitly represent the prior information nor provide a clear mechanism by which it is incorporated into the reconstruction.
-On the other hand, a conventional CS approach like STV precisely defines the prior information and the guidance mechanism.
-A reconstruction method combining this form of transparency with the effectiveness of deep learning-based methods, while maintaining lenient data requirement, is still needed.
+One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke [7] introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on the edge features in the reference scan.
+Later work introduced adaptive elements into the multi-contrast CS framework, e.g.
+Weizman et al.
+[8] proposed adaptive weighting-based guided CS and Song et al.
+[29] used adaptive sparse domains based on coupled dictionary learning.
+In the latter, the problem was formulated using a patch-level linear model comprising coupled and distinct sparse dictionary representations of the two contrasts.
+This model resembles a content/style model in form, although it is more restrictive.
+A general drawback of classical methods compared to deep learning-based ones is their lower flexibility.
+End-to-end learning-based methods [9, 13, 30, 31, 32, 15], on the other hand, supply the reference scan as an additional input to a deep reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proven to be more effective than conventional algorithms.
+However, end-to-end methods require large paired training datasets comprising the ground-truth image and the k-space together with aligned reference images, which is too strong a constraint when working with retrospectively collected clinical data that reflects the natural inconsistencies of routine practice such as missing contrasts, differences in spatial resolutions, and the existence of inter-scan motion.
+Additionally, by relying on end-to-end-learned features, these methods are generally less interpretable than their conventional counterparts in that they do not explicitly model the multi-contrast problem in terms of the underlying shared and non-shared information.
+MC-VarNet [15] is an important exception, which uses a simple linear decomposition of the reference contrast into common and unique components, applying the common component for guidance.
+A reconstruction method that is as effective as the learning-based methods while having more lenient data requirements and offering a high degree of interpretability is still needed.
+Our plug-and-play method relies on learning a deep non-linear content/style transform from image-domain data only, even if subject-wise paired images are not fully available.
 II-B Unpaired Image-to-Image Modeling
 Image-to-image modeling is the general problem of learning a mapping between two image domains and was first addressed by Pix2Pix [33] and CycleGAN [34] in paired and unpaired settings, respectively.
-Another line of unpaired image translation methods, the first of which was UNIT [18], assumes a shared latent space underlying the two domains to explicitly represent shared information, thereby making it more interpretable than CycleGAN which directly models the image-to-image mapping.
-However, both assume a deterministic one-to-one mapping between the two domains, ignoring the fact that an image in one domain can have multiple valid renderings in the other.
+Another line of unpaired image translation methods, the first of which was UNIT [21], assumes a shared latent space underlying the two domains to explicitly represent shared information.
+However, both CycleGAN and UNIT assume a deterministic one-to-one mapping between the two domains, ignoring the fact that an image in one domain can have multiple valid renderings in the other.
 Deterministic image translation has been widely applied to MRI contrast-to-contrast synthesis [17, 35, 18].
-However, they fundamentally do not account for the variability of the scanning setup which influence the realized contrast level and the differential visibility of pathologies in the target image.
+However, fundamentally, these methods do not account for the variability of the scanning setup that influences the realized contrast level and the differential visibility of pathologies in the target image.
 Denck et al.
@@
 sequence parameters) must be available.
-MUNIT [19] extended UNIT by modeling domain-specific variability in addition to domain-independent structure, overcoming the deterministic mapping limitation of UNIT.
+MUNIT [22] extended UNIT by modeling domain-specific variability in addition to domain-independent structure, enabling many-to-many mapping and thus overcoming the rigidity of UNIT.
 The result was a stochastic image translation model which, given an input image, generates a distribution of synthetic images sharing the same “content” but differing in “style”.
+More fundamentally, MUNIT is a learned invertible transformation between the image domains and the disentangled content/style domains.
+And unlike other content/style modeling frameworks such as [36, 37, 38], MUNIT models content and style as latent generative factors of the two image domains, providing precise distribution-level definitions for them.
 Clinical MR images of a given protocol contain multiple modes of variability, many of which are not known a priori.
-We argue that MUNIT, due to the minimal assumptions it makes, can be generally applied to learn domain-independent anatomical structure and domain-specific variability from an unpaired multi-contrast dataset as content and style, respectively.
+Compared to [19], we make a broader assumption that the contrast-independent semantic information is local in nature and that the contrast-specific variations in the dataset can include global effects of acquisition settings as well as local anatomical features unique to the contrast.
+We model these using MUNIT, referring to the shared and non-shared components as content and style, respectively.
 II-C Combining Image Translation with Reconstruction
-Acknowledging the fundamental limitations of MR contrast prediction, some prior work has attempted combining it with multi-contrast reconstruction.
-A naive form of guided reconstruction involves generating a synthetic image from the reference scan via deterministic image translation and using it in L2-regularized classical reconstruction [32].
+Acknowledging the limitation of MR cross-contrast prediction, some prior work has attempted combining it with multi-contrast reconstruction.
+A naive form of joint image synthesis and reconstruction, e.g.
+PROSIT [39] and rsGAN [40], involves generating a synthetic image from the reference scan via deterministic image translation and using it in a classical L2-regularized least-squares reconstruction.
 More recent work by Xuan et al.
-[33] proposes a joint image translation and reconstruction method based on optimal transport which additionally accounts for misalignment between the reference and reconstruction images.
+[41] proposed a joint image translation and reconstruction method which additionally accounts for misalignment between the reference and reconstructing images, with a follow-up work leveraging optimal transport theory [42].
 Levac et al.
 [16] formulate guided reconstruction as a highly general Bayesian maximum a posteriori estimation problem and solve it iteratively via Langevin update steps, using an image-domain diffusion model as the score function of the prior distribution.
-While these are promising directions, an alternative general approach leveraging semantic content/style modeling in multi-contrast reconstruction could potentially provide a connection with the well-established MR physics concepts of quantitative tissue maps, acquisition parameters, and physical models.
-Such an interpretable and unifying approach remains to be found.
+While these are promising directions, we propose an alternative approach that decomposes the multi-contrast problem in a more intuitive way – first, into two sub-problems, namely image modeling and image reconstruction; and second, within the image-domain model, the multi-contrast information is decomposed into content and style.
+This two-level decomposition results in a highly modular reconstruction algorithm with a built-in explanatory framework where (a) the guidance mechanism is a simple content-replacement operation, (b) the discrepancy between the supplied reference content and the true content of the target image represents a meaningful error term which can be minimized, and (c) the optimal content-encoding capacity of the model for a given two-contrast image dataset indicates the amount of shared structure available to be learned in this data and used in the reconstruction task.
 III Methods
@@
 $r^{k}\leftarrow f_{\theta}(x^{k-1}).$ (4)
-III-A3 Plug-and-play content restoration operator
-In guided reconstruction, a spatially aligned reference $x^{\mathrm{ref}}_{1}$ is available, which captures the same underlying semantic content as the reconstruction image.
-Inspired by the PnP design, we cast the problem of incorporating prior information from $x^{\mathrm{ref}}_{1}$ into reconstruction $x_{2}^{k-1}$ as the restoration of its semantic content, which has been corrupted by undersampling.
-We propose a content restoration operator $g_{M}(.;c)$ such that (2) becomes
+III-A3 Plug-and-play content consistency operator
+In guided reconstruction, a spatially aligned reference $x^{\mathrm{ref}}_{1}$ is available, which captures the same underlying semantic content as the target reconstruction.
+Inspired by the PnP design, we cast the problem of incorporating prior information from $x^{\mathrm{ref}}_{1}$ into the reconstruction iterate $x_{2}^{k-1}$ as enforcing a hard consistency between this image iterate and its semantic content estimated from the reference.
+We propose a content consistency operator $g_{M}(\cdot;c)$ such that
 $r_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c),$ (5)
-where a content/style model $M$ is used to decompose $x_{2}^{k-1}$ into content and style, followed by a replacement of this content with clean content $c$ derived from $x^{\mathrm{ref}}_{1}$ and composing the image $r_{2}^{k}$ from it.
+where a content/style model $M$ decomposes $x_{2}^{k-1}$ into content and style, followed by a replacement of this corrupted content with high-quality content $c$ derived from $x^{\mathrm{ref}}_{1}$ and composing the improved image $r_{2}^{k}$ from it.
 Before formally defining this operator and developing the reconstruction algorithm, we discuss the steps required to learn the content/style model $M$.
 III-B Content/Style Modeling
-Given an image dataset of two MR contrasts we make the general assumption that there exists an underlying contrast-independent spatial structure which, influenced by arbitrary contrast-specific factors, is rendered as the contrast images.
-Figure 2: A trained content/style model $M$ and the content restoration operator $g_{M}(.;c)$ based on it.
-(a) T1W/T2W pair and corresponding content maps.
-Despite general agreement between their content, there is notable discrepancy.
-(b) Synthetic T2W images generated from 16 style codes grid-sampled from $\mathcal{S}_{2}$, showing a smooth variation in image contrast and a roughly constant anatomical structure.
-(c) $g_{M}(.;c)$ is applied on 3 cases where $x_{2}^{*}$ was corrupted with $R$=4 undersampling with different center fractions.
-Sampling more low-frequency lines leads to more accurate style estimate $\hat{s}_{2}$ and thus better image $x_{2}^{\mathrm{restored}}$.
-Contours in $\mathcal{S}_{2}$ indicate the mean absolute error of the synthetic images from that region.
+Given an image dataset of two MR contrasts we make the general assumption that there exists an underlying contrast-independent structure which, influenced by arbitrary contrast-specific factors, is rendered as the contrast images.
 III-B1 Unpaired pre-training of MUNIT
 We formulate our content/style model based on the MUNIT framework [22].
-We define “domain” as the set $\mathcal{X}_{i}$ of images of a certain contrast comprising the dataset, where $i\in\{1,2\}$.
-“Content” $c\in\mathcal{C}$ is defined here as the underlying contrast-independent structure and is represented as a set of spatial maps, whereas “style” $s_{i}\in\mathcal{S}_{i}$ corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g.
-global effects of acquisition settings, contrast-specific tissue features, etc, and is thus represented as a low-dimensional vector.
-MUNIT posits the existence of functions $G_{i}^{*}:\mathcal{C}\times\mathcal{S}_{i}\rightarrow\mathcal{X}_{i}$ and their inverses $E_{i}^{*}=(G_{i}^{*})^{-1}$, and learns them jointly via unpaired training given samples from marginal distributions $p(x_{i})$.
+We define an image domain as the set $\mathcal{X}_{i}$ of images of a certain contrast comprising the dataset, where $i\in\{1,2\}$.
+“Content” $c\in\mathcal{C}$ is defined here as the underlying contrast-independent structure and is represented as a set of feature maps, whereas “style” $s_{i}\in\mathcal{S}_{i}$ corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g.
+global effects of acquisition settings, contrast-specific tissue features, etc., and is represented as a low-dimensional vector.
+MUNIT posits the existence of functions $G_{i}^{*}:\mathcal{C}\times\mathcal{S}_{i}\rightarrow\mathcal{X}_{i}$ and their inverses $E_{i}^{*}=(G_{i}^{*})^{-1}$, and learns them jointly via unpaired training, given samples from marginal distributions $p(x_{i})$.
 In practice, the encoder $E_{i}$ is split into content encoder $E_{i}^{c}$ and style encoder $E_{i}^{s}$.
 Thus, the content/style model is specified as $M=\{E_{1}^{c},E_{2}^{c},E_{1}^{s},E_{2}^{s},G_{1},G_{2}\}$.
-The MUNIT loss function comprises of 4 terms:
+The MUNIT loss function is comprised of 4 terms:
 $\mathcal{L_{\mathrm{MUNIT}}}=\mathcal{L}_{\mathrm{GAN}}+\alpha_{1}\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}+\alpha_{2}\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}+\alpha_{3}\mathcal{L}_{\mathrm{style}}^{\mathrm{self}},$ (6)
@@
 $\begin{split}\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}=\>&\mathbb{E}_{x_{1}\sim p(x_{1})}[||x_{1}-G_{1}(E_{1}^{c}(x_{1}),E_{1}^{s}(x_{1}))||_{1}]+\\ &\mathbb{E}_{x_{2}\sim p(x_{2})}[||x_{2}-G_{2}(E_{2}^{c}(x_{2}),E_{2}^{s}(x_{2}))||_{1}],\end{split}$ (8)
-$\begin{split}\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}=\>&\mathbb{E}_{c_{1}\sim p(c_{1}),s_{2}\sim q(s_{2})}[||c_{1}-E_{1}^{c}(G_{2}(c_{1},s_{2}))||_{1}]+\\ &\mathbb{E}_{c_{2}\sim p(c_{2}),s_{1}\sim q(s_{1})}[||c_{2}-E_{2}^{c}(G_{1}(c_{2},s_{1}))||_{1}],\end{split}$ (9)
+$\begin{split}\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}=\>&\mathbb{E}_{c_{1}\sim p(c_{1}),s_{2}\sim q(s_{2})}[||c_{1}-E_{2}^{c}(G_{2}(c_{1},s_{2}))||_{1}]+\\ &\mathbb{E}_{c_{2}\sim p(c_{2}),s_{1}\sim q(s_{1})}[||c_{2}-E_{1}^{c}(G_{1}(c_{2},s_{1}))||_{1}],\end{split}$ (9)
 $\begin{split}\mathcal{L}_{\mathrm{style}}^{\mathrm{self}}=\>&\mathbb{E}_{c_{1}\sim p(c_{1}),s_{2}\sim q(s_{2})}[||s_{2}-E_{2}^{s}(G_{2}(c_{1},s_{2}))||_{1}]+\\ &\mathbb{E}_{c_{2}\sim p(c_{2}),s_{1}\sim q(s_{1})}[||s_{1}-E_{1}^{s}(G_{1}(c_{2},s_{1}))||_{1}].\end{split}$ (10)
-$\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}$ is the autoencoder image recovery loss, which promotes preservation of input information in the latent space.
+$\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}$ is the image recovery loss, which promotes preservation of the image information in the latent space.
 $\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}$ and $\mathcal{L}_{\mathrm{style}}^{\mathrm{self}}$ are the content and style recovery losses, respectively, which specialize parts of the latent code into content and style components.
@@
 $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are an additional set of hyperparameters.
-III-B3 Network architecture and locality bias
+III-B3 Network architecture and content capacity
 Following the original paper [22], our content encoders $E_{i}^{c}$ consist of an input convolutional layer potentially followed by strided downsampling convolutions, and finally a series of residual blocks.
@@
 Style is introduced into the decoder via AdaIN operations which modulate the activation maps derived from the content.
-We note that in this architecture, the content-size-to-image-size ratio reflects the level of spatial structure one expects to be shared between the two domains.
-This (relative) content size is thus an inductive bias, known as the locality bias [34], built into the model.
-A model with high locality bias has a large content size, allowing for a richer content that strongly influences the output’s spatial structure, while simultaneously restricting style to learn low-level global features.
-III-C Iterative Reconstruction with Content Consistency
-Given the content/style model $M$, we consider $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ as reference and target domains, respectively and define the content restoration operator $g_{M}(.;c)$ as
-$x^{\mathrm{restored}}_{2}=g_{M}(x^{\mathrm{corrupt}}_{2};c_{1}^{\mathrm{ref}}):=G_{2}(c_{1}^{\mathrm{ref}},E^{s}_{2}(x^{\mathrm{corrupt}}_{2})),$ (14)
-where $x^{\mathrm{corrupt}}_{2}$ is the reconstruction corrupted with undersampling artifacts, $x^{\mathrm{restored}}_{2}$ is the restored image, and content $c_{1}^{\mathrm{ref}}=E^{c}_{1}(x^{\mathrm{ref}}_{1})$.
-This operation restores $x^{\mathrm{corrupt}}_{2}$ by simply replacing its aliased content with the clean content from $x^{\mathrm{ref}}_{1}$, a rule which will later be softened with (15).
-Note that this is a radical operation as it discards all structure contained in $x^{\mathrm{corrupt}}_{2}$, retaining only a compact style code $\hat{s}_{2}=E_{2}^{s}(x^{\mathrm{corrupt}}_{2})$.
-Let $x_{2}^{*}$ be the ground truth reconstruction with content $c_{2}^{*}$ and style $s_{2}^{*}$.
-The success of the content restoration operation depends on two conditions – (a) $c_{1}^{\mathrm{ref}}$ is close to $c_{2}^{*}$ and (b) $\hat{s}_{2}$ is close to $s_{2}^{*}$.
-The first condition is roughly satisfied because the model is explicitly trained to minimize content discrepancy, which can be seen in Fig.
-2a, but more on this later.
+We observe that in this architecture, the ratio between the content and image resolutions reflects the level of local structure one expects to be shared between the two domains.
+This (relative) content resolution is thus an inductive bias built into the model, which we refer to as the model’s content capacity.
+This concept is closely related to the general concept of the locality bias of image-to-image models [43].
+A model with high content capacity has a large content resolution, allowing it to learn a rich content representation that strongly influences the output’s structure, which simultaneously restricts the style spaces to learn low-level global features.
+Figure 2: A trained content/style model $M$ and the content consistency operator $g_{M}(\cdot;c)$ based on it.
+Images shown here are from the NYU brain DICOM dataset.
+Practical details are provided in Section IV.
+(a) A T1W/T2W image pair and the corresponding content maps.
+While these two contents generally agree, there is a notable discrepancy.
+(b) Synthetic T2W images generated from 16 style codes grid-sampled from $\mathcal{S}_{2}$, showing a smooth variation in image contrast and a roughly constant anatomical structure.
+(c) $g_{M}(\cdot;c)$ is applied on 3 cases where $x_{2}^{*}$ was corrupted with $R$=4 undersampling with different center fractions.
+Sampling more low-frequency lines leads to more accurate style estimate $\hat{s}_{2}$ and thus a better image $x_{2}^{\mathrm{cc}}$.
+Contours in $\mathcal{S}_{2}$ indicate the MAE of the synthetic images from that region.
+The red circle indicate a style estimation NMSE of 0.1.
+III-C Iterative Reconstruction using Content Consistency
+Given the content/style model $M$, we consider $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ as reference and target domains, respectively, in our guided reconstruction task, and define the content consistency operator $g_{M}(\cdot;c)$ from Eq.
+(5) as
+$x^{\mathrm{cc}}_{2}=g_{M}(x^{\mathrm{us}}_{2};\hat{c}):=G_{2}(\hat{c},E^{s}_{2}(x^{\mathrm{us}}_{2})),$ (14)
+where $x^{\mathrm{us}}_{2}$ is the image containing undersampling artifacts, $x^{\mathrm{cc}}_{2}$ is the content-consistent image, and $\hat{c}=E^{c}_{1}(x^{\mathrm{ref}}_{1})$ is the reference content.
+This operation improves $x^{\mathrm{us}}_{2}$ by simply replacing its aliased content with the content estimated from $x^{\mathrm{ref}}_{1}$, a rule which will later be softened with (16).
+Note that this is a radical operation as it discards all structure contained in $x^{\mathrm{us}}_{2}$, retaining only a compact style code $\hat{s}_{2}=E_{2}^{s}(x^{\mathrm{us}}_{2})$.
+Let $x_{2}^{*}$ be the ground truth reconstruction with content $c^{*}$ and style $s_{2}^{*}$.
+The success of our content consistency operation depends on two conditions – (a) $\hat{c}$ is close to $c^{*}$ and (b) $\hat{s}_{2}$ is close to $s_{2}^{*}$.
+The first condition is roughly satisfied, as seen in Fig.
+2a, because the model is explicitly trained to minimize content discrepancy, but more on this later.
 On the other hand, it is not obvious that the second condition should hold too and hence deserves a closer look.
-Assuming a high degree of shared spatial structure between the two domains, the optimal model has high locality bias (based on Section III-B3).
-Thus, style represents low-level global image features, e.g.
-contrast variations, as observed in Fig.
+Assuming a high degree of shared local structure between the two domains, the optimal model has high content capacity (see Section III-B3).
+Thus, style would represent mostly low-level global image features, e.g.
+contrast variations, as indeed observe in Fig.
 2b.
@@
 2c.
+Applying data consistency update (3) following $g_{M}(\cdot;\hat{c})$ in repetition yields an ISTA-based iterative scheme.
+Here, our content consistency update complements the data consistency update in the sense that while the latter forces the image estimate to be consistent with the given (measured) k-space data $y$, the former forces it to be consistent with the given (prior) content $\hat{c}$.
 Algorithm 1 PnP-MUNIT
@@
 2:$k\leftarrow 0$
-3:$x_{2}^{k}\leftarrow A^{H}y$
-4:$c^{k}\leftarrow E_{1}^{c}(x^{\mathrm{ref}}_{1})$
+3:$x_{2}^{k}\leftarrow A^{H}y$triangleright Initialize Reconstruction
+4:$c^{k}\leftarrow E_{1}^{c}(x^{\mathrm{ref}}_{1})$triangleright Initialize Content
 5:repeat
@@
 11:return $x_{2}^{k}$
-Applying data consistency update (3) following $g_{M}(.;c)$ in repetition yields an iterative scheme, where the latter can be viewed as a hard content consistency update.
-Fig.
-2a shows that although $c_{1}^{\mathrm{ref}}$ and $c_{2}^{*}$ are roughly similar, there is notable discrepancy between them.
-There are at least two possible sources of content discrepancy, which make this phenomenon inevitable – (a) model-related, e.g.
-practical issues like sub-optimal training or fundamental limits such as irreducible error, and (b) reference image-related, e.g.
-presence of artifacts independent of the reconstruction.
-During reconstruction, a $c_{1}^{\mathrm{ref}}$ that conflicts with $c_{2}^{*}$, and with k-space $y$ by extension, would hamper convergence.
-While model-related discrepancy is partly tackled by the PFT strategy (Section III-B2), we propose a content refinement (CR) step to correct for the remaining discrepancy in the reconstruction stage.
-The content state is initialized with $c_{1}^{\mathrm{ref}}$ and gradually refined to be consistent with the acquired k-space $y$ as
-$c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},E_{2}^{s}(x_{2}^{k}))-y||_{2}^{2},$ (15)
-where $\gamma$ is the step size.
-This update equation mirrors the form of the data consistency update (3), except here the gradient is taken with respect to content through the (non-linear) augmented forward operator $AG_{2}(.)$, which relates the latent content directly to the measured k-space.
-Hence, by aligning content with k-space explicitly, CR aligns content consistency updates with data consistency updates, correcting for any conflicts between them caused due to content discrepancy.
-We refer to the resulting algorithm as PnP-MUNIT (Algorithm 1).
-IV Experiments
-Without loss of generality, we consider the special case of reconstructing T2W scans using T1W references and focus on head applications.
-In IV-A, we present an analysis of locality bias, convergence, and tuning using BrainWeb-based simulated MR datasets [35].
-In IV-B, we benchmark against ablations and closely related methods on NYU fastMRI brain dataset [36, 37] and two in-house multi-coil raw datasets of the brain.
-Finally, in IV-C, we present a radiological evaluation comparing single- and multi-contrast reconstructions.
-All software was implemented in PyTorch, and training was performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
-IV-A Experiments on Simulated MR Dataset
+The core assumption of our idealized content/style model is that contents $\hat{c}$ and $c^{*}$ are identical.
+However, in reality, this assumption will not hold, and a discrepancy between the two contents is to be expected.
+There are two possible sources of this discrepancy – (a) model-related, e.g.
+fundamental limits such as irreducible error [^1] If perfectly zero error was possible, one could perfectly predict the (distribution of) target image from the reference without any measurement.
+and practical issues like sub-optimal training, and (b) reference image-related, e.g.
+presence of artifacts independent of the target reconstruction.
+This content discrepancy, as observed empirically, is shown in Fig.
+2a.
+During reconstruction, the error in $\hat{c}$ would limit the efficacy of the operator $g_{M}(\cdot;\hat{c})$, affecting reconstruction quality.
+While model-related discrepancy is partly tackled by PFT (Section III-B2), we now propose a content refinement (CR) process to correct for the remaining discrepancy in the reconstruction stage.
+Since we have no direct access to the true content $c^{*}$, but only to undersampled k-space measurements $y$, we aim at solving the following minimization problem in CR
+$\min_{c}||AG_{2}(c,\hat{s}_{2})-y||_{2}^{2},$ (15)
+starting from the initial point $\hat{c}=E^{c}_{1}(x^{\mathrm{ref}}_{1})$ and for a given style estimate $\hat{s}_{2}$.
+The augmented forward operator $AG_{2}(\cdot)$ is a composition of the linear MRI forward operator $A$ and the non-linear content/style decoder $G_{2}$, and it maps the content domain to the k-space domain.
+The error between the predicted k-space $AG_{2}(c,\hat{s}_{2})$ and the measured data $y$ serves as a proxy for content discrepancy, which can be computed and minimized during the reconstruction.
+We approximate the solution with a single gradient descent step
+$c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},\hat{s}_{2}^{k})-y||_{2}^{2},$ (16)
+initialized as $c^{0}\leftarrow E^{c}_{1}(x^{\mathrm{ref}}_{1})$ and updated every $k^{\text{th}}$ iteration with step size $\gamma$ following the data consistency and content consistency updates.
+Hence, by aligning the content $c$ with k-space data $y$ and correcting the discrepancy, the CR module aligns content consistency updates with data consistency updates.
+With this additional component in place, we obtain our PnP-MUNIT reconstruction algorithm (Algorithm 1).
+Note that unlike content, the style of the image estimate need not be explicitly aligned with the k-space.
+The data consistency update applied on the image implicitly “corrects” its style and as the image converges, its style would converge as well.
+IV Experimental Setup
+IV-A Experiment Design
+Without loss of generality, we considered the case of reconstructing T2W scans using T1W references and focused on head applications.
+We conducted 4 sets of experiments to comprehensively evaluate PnP-MUNIT – (a) simulations, (b) benchmark on NYU DICOM data, (c) benchmark on clinical multi-coil raw data, and (d) radiological evaluation.
+TABLE I: Overview of the sequences in the LUMC datasets.
+Sequence Params LUMC-TRA LUMC-COR
+3D T1W TFE 2D T2W TSE 3D T1W TSE 2D T2W TSE
+FA (deg) 8 90 80-90 90
+TR (ms) 9.8-9.9 4000-5000 500-800 2000-3500
+TE (ms) 4.6 80-100 6.5-16 90-100
+ETL 200 14-18 10-13 17-19
+Voxel size (mm) 0.98$\times$0.99 $\times$0.91 0.4$\times$0.54 0.59$\times$0.62 $\times$1.19 0.39$\times$0.47
+Slice thick. (mm) 3 2
+FOV (mm) 238$\times$191 $\times$218 238$\times$190 130$\times$238 $\times$38 130$\times$197
+Num slices 50 15
+TABLE II: LUMC data split. The splits were made at the subject-level, i.e. each subject belonged to exactly one subset. Splits containing only image-domain data are marked with I.
+Dataset Split Subjects Sessions Scans (T1W / T2W)
+LUMC-TRA model-trainI 295 418 360 / 415
+model-valI 16 17 17 / 17
+recon-val 18 21 21 / 21
+recon-test 20 31 31 / 31
+Total 339 487 429 / 484
+LUMC-COR model-trainI 242 277 269 / 272
+model-valI 18 18 18 / 18
+recon-val 15 18 18 / 18
+recon-test 16 17 17 / 17
+Total 291 330 322 / 325
+IV-A1 Simulation
+The goal of our simulation experiments was to empirically test several properties of our algorithm under highly controlled settings.
+To this end, we used simulated T1W and T2W images based on BrainWeb phantoms [44].
 BrainWeb provides 20 anatomical models of normal brain, each comprised of fuzzy segmentation maps of 12 tissue types.
@@
 T1W/T2W spin-echo scans were simulated using TE/TR values randomly sampled from realistic ranges.
-For testing, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling.
-IV-A1 Interpretation of content size
-TABLE I: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices) at $R$=4. Optimal locality bias is shown in bold. PSNR worse than L1-wavelet CS (16.83$\pm$0.07) is marked as .
+For testing the reconstruction, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling at various accelerations.
+We analyzed the effect of the content/style model’s disentanglement level and its content capacity on the reconstruction quality, the convergence of the PnP-MUNIT algorithm, and the effectiveness, robustness, and tuning of the CR module.
+IV-A2 Benchmark on NYU DICOM dataset
+In the NYU benchmark, we compared PnP-MUNIT against end-to-end reconstruction methods.
+The question we sought to answer was whether or not PnP-MUNIT, which requires only image-domain training data, can outperform methods that additionally require k-space training data.
+The NYU brain dataset [45, 46] includes raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR.
+Following [41], a paired DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans.
+All T1W scans were rigid-registered with the corresponding T2W scans.
+Single-coil T2W k-space data was simulated via Fourier transform with 1D Cartesian random sampling at $R\in\{2,3,4,5\}$ and added Gaussian noise of $\sigma=0.01\mathrm{max}(x_{2}^{*})$.
+The dataset was split into 4 subsets – model-train (200), model-val (27), recon-val (50), and recon-test (50).
+Images from the former two splits were used to pre-train, fine-tune, and validate the MUNIT model.
+On the other hand, end-to-end models were trained and validated on the aligned T1W, the simulated T2W k-space, and the T2W ground-truth from these two splits.
+The recon-val split was used to tune the PnP-MUNIT algorithm, and recon-test was the held-out test set.
+We benchmarked PnP-MUNIT against one well-known single-contrast network – MoDL [4] – and two recent multi-contrast networks – MTrans [32] and MC-VarNet [15].
+IV-A3 Benchmark and ablation on LUMC multi-coil dataset
+In our benchmark on clinical multi-coil data, our goal was to test PnP-MUNIT on a constrained real-world problem where only the image-domain data was available for training.
+This is often the case, as in clinical practice raw data is discarded after acquisition and only the final reconstructions are retained.
+Moreover, the T1W/T2W images were not fully subject-wise paired, representing a realistic case of data imbalance.
+Our in-house data consisted of brain scans of patients from LUMC, the use of which was approved for research purpose by the institutional review board.
+A total of 1669 brain scans were obtained from 817 clinical MR examinations of 630 patients acquired on 3T Philips Ingenia scanners.
+We focused on accelerating two T2W sequences – (a) 2D T2W TSE transversal and (b) 2D T2W TSE coronal.
+For guidance, two corresponding T1W sequences were used – (a) 3D T1W TFE transversal and (b) 3D T1W TSE coronal.
+The transversal and coronal protocols were considered as two separate datasets, namely LUMC-TRA and LUMC-COR.
+Table I shows an overview of the four sequences.
+Note that the in-plane resolution of the T1W scans was 1.3-2.5 times as low as that of the T2W scans, making these datasets more challenging for T1W-guided T2W reconstruction.
+As with the NYU dataset, the LUMC datasets were split into model-train, model-val, recon-val, and recon-test.
+The former two splits contained only image-domain data, which was used to train and validate MUNIT.
+The MUNIT models were pre-trained on the full model-train splits, ignoring any pairing between T1W and T2W scans.
+Additionally, 20 subjects from the model-train split were designated for PFT where the pairing information was used in training and the reference scans were aligned via registration.
+The recon-val and recon-test splits additionally included multi-coil T2W raw data, which comprised 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps.
+This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of $R$=1.8-2.
+We further undersampled it retrospectively to higher accelerations $R\in\{4,6,8,10\}$ by dropping subsets of the acquired lines.
+An overview of the data split is shown in Table II.
+Spatially aligned images required by all but the unpaired training set were obtained via rigid registration and the reference T1W images were resampled to the T2W resolution.
+During unpaired training, all images were resampled to the median T2W resolution.
+Given the of absence of k-space training data, end-to-end reconstruction methods were not feasible as baselines.
+Hence, we compared PnP-MUNIT with only the feasible types of baselines, i.e.
+classical, semi-classical, and plug-and-play reconstruction and image-to-image translation.
+Among classical methods, we used the unguided L1-wavelet CS (CS-WT) and the guided STV-based CS (CS-STV) [7].
+As an unguided plug-and-play baseline, we used PnP-CNN [23], and as a representative image translation baseline, we compared against MUNIT itself.
+Using MUNIT, deterministic image translation was approximated by combining the reference content with 200 randomly sampled T2W style codes and taking a pixel-wise mean of the resulting synthetic images to obtain a single synthetic image.
+As semi-classical guided baseline, we used (PROSIT) [39], which combines deterministic image translation with L2-regularized least-squares reconstruction.
+Additionally, as an ablation study for PnP-MUNIT, we ablated PFT and CR to assess their contribution, and finally, as an upper-bound for PnP-MUNIT representing zero content discrepancy, we disabled PFT and CR and used the ideal content $c^{*}$.
+In both NYU and LUMC benchmarks, we used 3 perceptual metrics for evaluation – SSIM, HaarPSI, and DISTS.
+While SSIM is used commonly, HaarPSI and DISTS are known to correlate better with visual judgment of image quality [47].
+All three metrics are bounded in $[0,1]$ where $1$ represents perfect image quality.
+We conducted paired Wilcoxon signed-rank tests to measure statistical significance when comparing pairs of algorithms.
+IV-A4 Radiological evaluation
+Finally, as an extension to the LUMC benchmark, we conducted a radiological evaluation assessing the visual and diagnostic quality of PnP-MUNIT reconstructions.
+The study was conducted on a small sample of the LUMC-TRA recon-test set.
+With the help of a junior radiologist, we selected 3 cases with brain metastases.
+The evaluation comprised two parts, namely visual quality and pathology.
+Four visual quality criteria were used, namely sharpness, noise, artifacts, and contrast between gray and white matter and CSF.
+For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases.
+The images were scored using a five-point Likert scale [48] – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality.
+A diagnostic-quality reconstruction was defined as one that showed at least 90% of the metastases and scored at least “fairly diagnostic quality” on all other criteria.
+[^2] In clinical practice, while the remaining 10% of the metastases may not be detected in the reconstructed T2W scan, they would be prominently visible in the corresponding gadolinium-enhanced T1W scan, and hence would not go undetected.
+We evaluated PnP-MUNIT and PnP-CNN at 4 clinically realistic accelerations of $R\in\{3,4,5,6\}$.
+In total, 9 images per patient, including the (4$\times$2=8) reconstructions and the clinical ground truth (acquired at $R$=2 and reconstructed by the vendor software), were presented to a senior neuroradiologist who scored each image individually, blinded to the reconstruction method and $R$.
+IV-B Implementation Details
+We used the same general residual architecture for MUNIT encoders and decoders as the original paper, except with an additional layer at the end of the content encoder to produce content maps of given number of channels.
+We used 2 or 4 channels, depending on the content downsampling factor.
+The discriminators were implemented as multi-scale PatchGAN networks enabling them to locally assess the input images for realism at different scales.
+We used 1 scale in the BrainWeb simulations and 3 scales in the NYU and LUMC benchmark.
+We additionally conditioned the discriminators on foreground masks of the images to penalize background signal.
+To stabilize GAN training, we used spectral normalization [49].
+For hyperparameters, we found that pre-training loss weight values $\alpha_{1}=\alpha_{2}=\alpha_{3}=1$ and paired fine-tuning values $\beta_{1}=\beta_{2}=\beta_{3}=1$ worked generally well for the in vivo datasets.
+In the benchmarks, the CR module parameter $\gamma$ of PnP-MUNIT was tuned per acceleration for each dataset and the number of iterations was set to 200.
+In the simulations, all MUNIT models were pre-trained for 200k and were not fine-tuned for the sake of simplicity, unless stated otherwise in Section V-A.
+In both NYU and LUMC benchmarks, the MUNIT models were pre-trained for 400k iterations and fine-tuned for 50k iterations.
+All software was implemented in PyTorch, and image registration was performed using Elastix [50].
+All training runs were performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
+V Results
+TABLE III: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices). PSNR worse than L1-wavelet CS is marked as .
+Disentanglement Strength R=2 R=4
+$\alpha_{2}=\alpha_{3}=10$ 15.14 $\pm$ 0.06 12.84 $\pm$ 0.05
+$\alpha_{2}=\alpha_{3}=1$ 28.43 $\pm$ 0.07 26.13 $\pm$ 0.07
+$\alpha_{2}=\alpha_{3}=0.1$ 29.84 $\pm$ 0.05 27.33 $\pm$ 0.05
+$\alpha_{2}=\alpha_{3}=0.01$ 25.34 $\pm$ 0.06 23.03 $\pm$ 0.05
+$\alpha_{2}=\alpha_{3}=0.001$ 18.17 $\pm$ 0.07 15.88 $\pm$ 0.05
+TABLE IV: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices) at $R$=4. Optimal content capacity is shown in bold. PSNR worse than L1-wavelet CS is marked as .
 Model Config Data Config
-RefSize-1 RefSize-2 RefSize-4 RefSize-8
-ContentSize-1 27.90$\pm$0.08 21.17$\pm$0.05 19.11$\pm$0.07 16.43$\pm$0.07
-ContentSize-2 26.19$\pm$0.08 22.80$\pm$0.08 20.05$\pm$0.08 16.57$\pm$0.08
-ContentSize-4 23.98$\pm$0.08 20.88$\pm$0.10 19.44$\pm$0.09 16.88$\pm$0.06
-The goal of this experiment was to (a) determine the relationship between the locality bias of the content/style model and the relative spatial resolution of the dataset and (b) explain the effect of reducing the reference image resolution on the reconstruction in terms of locality bias.
-We simulated 4 datasets, denoted as RefSize-$n$, where $n\in\{1,2,4,8\}$ represents the reference domain downsampling factor.
-In the $n$=1 case, T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower $1/n$ frequency components.
-For each dataset, we trained 3 models (for 200k iterations) with locality bias denoted as ContentSize-$m$, where $m\in\{1,2,4\}$ is the content downsampling factor.
-Table I compares PnP-MUNIT reconstruction quality across these configurations.
+RefRes-1 RefRes-2 RefRes-4 RefRes-8
+ContentRes-1 27.90$\pm$0.08 21.17$\pm$0.05 19.11$\pm$0.07 16.43$\pm$0.07
+ContentRes-2 26.19$\pm$0.08 22.80$\pm$0.08 20.05$\pm$0.08 16.57$\pm$0.08
+ContentRes-4 23.98$\pm$0.08 20.88$\pm$0.10 19.44$\pm$0.09 16.88$\pm$0.06
+Figure 3: (a) Convergence of PnP-MUNIT and the baseline versions at $R$=4. (b) Evolution of the reconstruction shown with its error map.
+Figure 4: Effectiveness of the CR module in resolving contrast-specific structure, which, in this case, was a simulated lesion.
+V-A Experiments on Simulated MR Datasets
+V-A1 Perturbation of disentanglement strength
+PnP-MUNIT relies heavily on a sufficient disentanglement of content and style from the images.
+Here, we evaluated the effect of perturbing the disentanglement loss weights $\alpha_{2}$ and $\alpha_{3}$ (Eq.
+(6)) of MUNIT on the PnP-MUNIT reconstruction quality.
+Table III shows the reconstruction PSNR for different weight values.
+We observed that optimal disentanglement, and thus the best reconstruction, was achieved at $\alpha_{2}=\alpha_{3}=0.1$.
+At lower values, poor disentanglement of content and style in the model resulted in poor PnP-MUNIT reconstruction quality.
+This was expected since PnP-MUNIT relies on the assumption that the content representation is (sufficiently) contrast-independent, which was less enforced at these levels.
+At higher values, the training over-emphasized disentanglement while under-emphasizing the GAN and image recovery loss terms, thereby leading to worse preservation of image information in the latent space and thus, to worse reconstructions.
+V-A2 Content resolution analysis
+PnP-MUNIT relies on a reference image of sufficiently high resolution to provide guidance to the reconstruction.
+The goal of this experiment was to determine the effect of lowering the reference image resolution on the reconstruction quality and using the content/style model’s content capacity (i.e.
+the spatial resolution of the content relative to the images) to explain this effect.
+We simulated 4 datasets, denoted as RefRes-$n$, where $n\in\{1,2,4,8\}$ represents the reference domain downsampling factor.
+In the $n$=1 case, T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower $1/n$ frequency components, while maintaining the same spatial resolution.
+For each dataset, we trained 3 content/style models with content capacity denoted as ContentRes-$m$, where $m\in\{1,2,4\}$ is the content downsampling factor, which depends on the number of downsampling blocks in the networks.
+E.g.
+ContentRes-2 model produced content maps half the spatial resolution of the ContentRes-1 model.
+Figure 5: The effect of misaligned reference on the reconstruction and the robustness of the CR module to it.
+Figure 6: Reconstruction PSNR as a function of acceleration $R$ and k-space noise level $\sigma$ for different values of the $\gamma$ parameter.
+$\gamma$ controls the consistency of the content with the measured k-space.
+Figure 7: Evaluation plots over recon-test subset of NYU DICOM dataset consisting of 550 slices.
+Figure 8: Two examples from the NYU benchmark.
+(a) Comparison of all the methods at the highest acceleration of $R$=5.
+(b) Comparison between PnP-MUNIT and MC-VarNet at $R$=2 and $R$=4 on a pathological case where many features of the edema are T2W contrast-specific.
+Table IV compares PnP-MUNIT reconstruction quality across these configurations.
 PFT and CR were disabled here for simplicity.
-We observe that (a) the optimal locality bias decreases with reference resolution and (b) the reconstruction quality decreases as well, eventually dropping below non-guided reconstruction.
-Both are explained by the fact that with decrease in reference contrast resolution, the amount of effective content in the dataset decreases, thereby limiting the value of guided reconstruction.
-IV-A2 Convergence
-Figure 3: (a) Convergence of PnP-MUNIT and the baseline versions at $R$=4. (b) Evolution of the reconstruction shown with its error map.
-We explore here the convergence of PnP-MUNIT by comparing it with two variants.
-The first uses the content $c_{2}^{*}$ of the ground truth T2W image, thus assuming zero content discrepancy and representing an upper-bound of PnP-MUNIT.
-The second uses the reference content $c_{1}^{\mathrm{ref}}$ but with CR step disabled, representing a lower-bound of PnP-MUNIT where the non-zero content discrepancy is not corrected.
-Here as well as in the following simulation experiment, we used the RefSize-1 dataset and the MUNIT-1 model additionally fine-tuned on 2 training volumes for 50k iterations.
+We observe two trends.
+First, the reconstruction quality generally decreases from left to right, eventually dropping below L1-wavelet CS reconstruction, suggesting a decrease in the amount of shared local information contained in the reference contrast.
+Second, the optimal content capacity of the model decreases with the reference resolution in accordance with the actual amount of this shared information.
+In other words, this optimal content capacity corresponds to the amount of contrast-independent structure the model discovers in the dataset, e.g.
+RefRes-1 dataset had 4 times as much shared structure (due to twice as large optimal content resolution) as RefRes-2 and RefRes-4 datasets and hence, the RefRes-1 dataset was significantly more effective in guided reconstruction.
+On the other hand, Refsize-2 and RefRes-4 datasets had similar levels of shared structure despite the lower reference resolution in the latter, suggesting a more complex relationship between the reference resolution and the shared content.
+In the RefRes-8 case, PnP-MUNIT dropped to a similar level of quality as conventional CS, suggesting a lowerbound on the amount of shared information for PnP-MUNIT to be effective, specifically 1/42=1/16 times the local information of the full-resolution case.
+V-A3 Convergence
+We explored the convergence of PnP-MUNIT by comparing it with two variants.
+The first used true content $c^{*}$ of the ground truth T2W image, thus assuming zero content discrepancy and representing an upper-bound of PnP-MUNIT.
+The second used reference content $\hat{c}$ with CR step disabled, representing a lower-bound of PnP-MUNIT where the non-zero content discrepancy is not corrected.
+Here and in the following simulation experiments, we used the RefRes-1 dataset and the ContentRes-1 model additionally fine-tuned on 2 training volumes for 50k iterations.
 Fig.
 3 shows the convergence curves and intermediate reconstructions.
-The upper-bound version, given $c_{2}^{*}$, converged in a single iteration, while with $c_{1}^{\mathrm{ref}}$, the ablated version converged equally fast but to a sub-optimal solution.
+The upper-bound version, given $c^{*}$, converged in a single iteration, while with $\hat{c}$, the ablated version converged equally fast but to a sub-optimal solution.
 Enabling the CR step closed the gap with the upper-bound, although costing convergence rate.
-Figure 4: Reconstruction PSNR as a function of acceleration $R$ and k-space noise level $\sigma$ for different values of the $\gamma$ parameter.
-$\gamma$ controls the consistency of the content with the measured k-space.
-IV-A3 Tuning the CR step size
+V-A4 Effectiveness in resolving contrast-specific structure
+In order to test the effectiveness of the CR module in resolving structure present in the target contrast but absent in the reference contrast, we simulated a lesion in the T2W image and performed reconstruction with and without the CR module.
+As shown in Fig.
+4, the CR module was able to recover the lesion fully at $R$=2 and substantially at $R$=4.
+V-A5 Sensitivity to misalignment of the reference
+While PnP-MUNIT expects an aligned reference image at reconstruction time, the CR module should, in principle, correct for small misalignments.
+To test this, we simulated a 2 rotation in the reference image and performed reconstruction with and without the CR module, comparing also with PnP-MUNIT that used an aligned reference image.
+As shown in Fig.
+5, although the reconstruction was sensitive to the misalignment in the absence of CR, it significantly improved with CR enabled.
+Figure 9: Evaluation plots over recon-test subset of LUMC-TRA and LUMC-COR datasets consisting of 1366 and 200 slices, respectively.
+Figure 10: Sample slice from the LUMC-COR recon-test set comparing PnP-MUNIT with PnP-CNN and MUNIT image translation.
+V-A6 Tuning the CR step size
 The hyperparameter $\gamma$ in Algorithm 1 controls the CR strength.
 Fig.
-4 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal $\gamma$ decreases with the amount and quality of k-space data.
+6 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal $\gamma$ decreased with the amount and quality of k-space data.
 This makes sense as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
-Figure 5: Evaluation plots over recon-test subset of NYU, LUMC-TRA, and LUMC-COR datasets, consisting of 550, 1366, and 200 slices, respectively.
-IV-B Benchmark and Ablation
-For benchmarking, we used L1-wavelet CS (CS-WT) and PnP-CNN [20] as our single-contrast baselines.
-The denoising model of PnP-CNN was defined as a 5-layer network with convolution and ReLU operations and a residual connection after the final layer.
-We additionally compared against deterministic image translation based on MUNIT defined as $x_{2}^{\mathrm{syn}}:=\mathbb{E}_{s_{2}\sim q(s_{2})}[G_{2}(E_{1}^{c}(x^{\mathrm{ref}}_{1}),s_{2})]$.
-As guided reconstruction baselines, we used STV-based CS (CS-STV) [6] and image translation combined with L2-regularized reconstruction (Translation + L2 recon) [32] given as $\mathop{\min}_{x_{2}}||Ax_{2}-y||_{2}^{2}+\lambda||x_{2}-x_{2}^{\mathrm{syn}}||_{2}^{2}$.
-For PnP-MUNIT, we ablated PFT and CR to assess their contribution, and finally, as an upper-bound representing zero content discrepancy, we disabled PFT and CR and used the ideal content $c_{2}^{*}$.
-TABLE II: Data split for benchmarking. Splits were performed at subject-level, each subject belonging to exactly one subset.
-Dataset Split Subjects Sessions Scans (T1W / T2W)
-NYU model-train 200 200 200 / 200
-model-val 27 27 27 / 27
-recon-val 50 50 50 / 50
-recon-test 50 50 50 / 50
-Total 327 327 327 / 327
-LUMC-TRA model-train 295 418 360 / 415
-model-val 16 17 17 / 17
-recon-val 18 21 21 / 21
-recon-test 20 31 31 / 31
-Total 339 487 429 / 484
-LUMC-COR model-train 242 277 269 / 272
-model-val 18 18 18 / 18
-recon-val 15 18 18 / 18
-recon-test 16 17 17 / 17
-Total 291 330 322 / 325
-The NYU brain dataset includes raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR.
-Following [33], a DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans.
-Single-coil T2W k-space data was simulated via Fourier transform with 1D Cartesian random sampling of $R\in\{2,3,4,5\}$ and added Gaussian noise of $\sigma=0.01\mathrm{max}(x_{2}^{*})$.
-Our in-house datasets consisted of brain scans of patients from LUMC, the use of which was approved for research purpose by the institutional review board.
-A total of 1669 brain scans were obtained from 817 clinical MR examinations of 630 patients acquired on 3T Philips Ingenia scanners.
-We focused on accelerating two T2W sequences – (a) 2D T2W TSE transversal (FA=90°, TR=4000-5000ms, TE=80-100ms, ETL=14-18, median voxel size=0.37$\times$0.5mm2, median FOV=238$\times$190mm2, median slice thickness=3mm, median slices=50), and (b) 2D T2W TSE coronal (FA=90°, TR=2000-3500ms, TE=90-100ms, ETL=17-19, median voxel size=0.39$\times$0.47mm2, median FOV=130$\times$197mm2, median slice thickness=2mm, median slices=15).
-For guidance, two corresponding T1W sequences were used – (a) 3D T1W TFE transversal (FA=8°, TR=9.8-9.9ms, TE=4.6ms, ETL=200, median voxel size=0.98$\times$0.99$\times$0.91mm3, median FOV=238$\times$191$\times$218mm3), and (b) 3D T1W TSE coronal (FA=80-90°, TR=500-800ms, TE=6.5-16ms, ETL=10-13, median voxel size=0.59$\times$0.62$\times$1.19mm3, median FOV=130$\times$238$\times$38mm3).
-The transversal and coronal protocols were considered as two separate datasets, namely LUMC-TRA and LUMC-COR.
-The multi-coil T2W raw data comprised of 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps.
-This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of $R$=1.8-2.
-We further undersampled it retrospectively to higher accelerations $R\in\{4,6,8,10\}$ by dropping subsets of the acquired lines.
-Each dataset was split into 4 subsets – (a) model-train, (b) model-val, (c) recon-val, and (d) recon-test.
-The former two were used to train and tune the plug-and-play models, recon-val to tune the reconstruction algorithms, and recon-test as the held-out test set.
-An overview of the data split is shown in Table II.
-Additionally, 20 subjects from model-train were designated for PFT.
-Spatially aligned images required by all except the unpaired training set were obtained via rigid registration using Elastix [38].
-We empirically found the ContentSize-1 and ContentSize-2 configurations as optimal for the NYU and LUMC datasets, respectively, which reflects the lower effective content and hence greater difficulty of the LUMC datasets.
-For each dataset, the model was pre-trained and fine-tuned for 400k and 50k iterations, respectively.
-We used 3 perceptual metrics for evaluation: SSIM, HaarPSI, and DISTS.
-While SSIM is commonly used, HaarPSI and DISTS are known to correlate better with visual judgment of image quality [39].
-All three metrics are bounded to $[0,1]$ with $1$ representing perfect image quality.
-We additionally conducted paired Wilcoxon signed-rank tests to measure statistical significance when comparing pairs of algorithms.
+V-B Benchmark on NYU DICOM Data
 Fig.
-5 plots the benchmark metrics, which can be summarized in the following three trends.
-First, pure image translation was worse compared to single-contrast reconstruction, especially at lower acceleration factors (comparing with PnP-CNN, $p$$<$0.05 for all metrics and datasets).
-Combining it with L2-regularized reconstruction improved SSIM and HaarPSI ($p$$<$0.05 for all datasets and accelerations except SSIM at $R$=3 in NYU), but not necessarily DISTS, suggesting that the the available complementary information was not fully utilized.
-PnP-MUNIT was consistently better than both image translation and naive guidance ($p$$<$0.05 throughout for both cases).
+7 shows evaluation metrics for PnP-MUNIT and the end-to-end baselines on the NYU DICOM dataset.
+Despite being trained only on image-domain data, PnP-MUNIT largely outperformed MoDL ($p<0.05$ throughout, except SSIM at $R$=2 and DISTS at $R\in${2,3,4}) and MTrans ($p<$0.05), and was roughly comparable to MC-VarNet.
+Moreover, as shown in Fig.
+8a, PnP-MUNIT could resolve certain fine details even at the high acceleration of $R$=5, which the baselines failed at.
+In the pathological case shown in Fig.
+8b, PnP-MUNIT produced sharper reconstructions preserving the structure of the edema better than MC-VarNet despite lower metrics when many features of edema were absent in the reference image.
+V-C Benchmark and Ablation on LUMC Multi-coil Data
+The LUMC datasets represent a more challenging problem with real-world data constraints.
+Additionally, in the light of the content resolution analysis (Section V-A2), we empirically found the ContentRes-2 model configuration as optimal for the LUMC datasets, compared to NYU DICOM data for which ContentRes-1 was the optimal model configuration.
+This reflects the lower effective content in LUMC image data and hence a greater difficulty for guided reconstruction.
+Fig.
+9 plots the benchmark metrics for the LUMC test sets, which can be summarized in the following three trends.
+First, pure image translation with MUNIT was worse compared to single-contrast reconstruction, especially at lower acceleration factors (comparing with PnP-CNN, $p$$<$0.05 for all metrics and both datasets).
+Combining it with L2-regularized reconstruction in PROSIT improved SSIM and HaarPSI ($p$$<$0.05 for both datasets and accelerations), but not necessarily DISTS, suggesting that the the available complementary information was not fully utilized.
+PnP-MUNIT was consistently better than both MUNIT image translation and PROSIT ($p$$<$0.05 throughout for both cases).
 It also outperformed the conventional guided CS-STV ($p$$<$0.05 throughout, except SSIM at $R$=4 in LUMC-COR).
-Second, both PnP-CNN and PnP-MUNIT performed similarly at lower acceleration.
-At higher acceleration, PnP-MUNIT outperformed PnP-CNN ($p$$<$0.05 for all metrics and datasets except SSIM at $R$=8 in LUMC-COR).
+Hence, using both measured k-space and reference scan via PnP-MUNIT was more beneficial than (a) using either one of them and (b) combining both using hand-crafted priors, suggesting that our approach maximally exploits the complementary information.
+Second, both PnP-CNN and PnP-MUNIT performed similarly at lower acceleration, except in terms of SSIM and HaarPSI at $R$=4 and $R$=6 on LUMC-COR where PnP-CNN was slightly better.
+At higher acceleration, PnP-MUNIT outperformed PnP-CNN ($p$$<$0.05 for all metrics and both datasets except SSIM at $R$=8 in LUMC-COR).
+Compared to PnP-CNN, PnP-MUNIT allowed up to 32.6% more acceleration for a given SSIM.
+[^3] Based on linear interpolation of LUMC-TRA SSIM, PnP-MUNIT and PnP-CNN allowed $R$=10 and $R$=6.7, the maximum difference in $R$, at median SSIM=0.906.
+Hence, a difference of 32.6%.
 Third, introducing PFT and CR into the ablated PnP-MUNIT improved the reconstructions ($p$$<$0.05 throughout in both cases except in DISTS for the latter at $R$=10 in LUMC-TRA).
 The contribution of PFT was almost constant across $R$, whereas that of CR decreased with $R$, which was expected since CR depends on the measured k-space data to refine the content.
-Figure 6: Sample slice from the LUMC-COR recon-test set comparing PnP-MUNIT with PnP-CNN and pure image translation.
 Fig.
-6 shows an example from LUMC-COR representative of the first two trends.
-Image translation produced severe anatomical defects and predicted structure in the lower region of the image where the ground truth contains low signal.
+10 shows an LUMC-COR example representative of the first two trends.
+MUNIT image translation produced severe anatomical defects and predicted false structure in the lower region of the image where the ground truth, in fact, contains low signal.
 At $R$=4, PnP-MUNIT was visually similar to PnP-CNN, with the main difference being a blur effect in PnP-CNN and mild texture artifacts in PnP-MUNIT.
@@
 However, an interesting failure mode of PnP-MUNIT at high acceleration was the subtle localized distortions in the anatomy (pink arrows).
-IV-C Radiological Evaluation
 Figure 11: Averaged results of the radiological evaluation.
-We conducted the radiological evaluation on a small sample of LUMC-TRA recon-test set.
-We selected 3 cases that showed brain metastases in corresponding gadolinium-enhanced T1W scans.
-The evaluation comprised two parts, namely visual quality and pathology.
-Four visual quality criteria were used, namely sharpness, noise, artifacts, and contrast between gray and white matter and CSF.
-For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases.
-The images were scored using a five-point Likert scale [40] – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality.
-We considered PnP-MUNIT and PnP-CNN at 4 clinically realistic accelerations $R\in\{3,4,5,6\}$.
-In total, 9 images per patient, including the reconstructions and the clinical ground truth of $R$=2, were presented to a senior neuroradiologist who scored each image individually, blinded to the reconstruction method and $R$.
-Figure 7 plots the evaluation result.
-At $R$=3, both algorithms produced at least “fairly diagnostic” reconstructions and allowed the detection of more than 90% of metastases.
+V-D Radiological Evaluation
+Figure 11 plots the radiological evaluation result.
+At $R$=3, both algorithms produced at least “fairly diagnostic” reconstructions and allowed the detection of more than 90% of metastases, a difference of 33.3% of k-space samples over the clinical reconstructions of $R$=2.
 PnP-CNN was better than PnP-MUNIT at $R\in\{3,4\}$ in terms of sharpness, contrast, and the pathology criteria.
@@
 At $R$=6, PnP-CNN sharply dropped to non-diagnostic level, whereas PnP-MUNIT was better, especially in terms of sharpness and contrast as well as in overall diagnostic quality.
-V Discussion
-We explored the interpretability aspect of PnP-MUNIT in terms of the effective content in a two-contrast dataset (Section IV-A1) and obtained a notion of limit to guided reconstruction in terms of the reference image resolution.
-This limit would indeed depend on more fundamental factors such as MR sequence types, which we did not investigate here.
-Our data-driven definition of content and style is loosely analogous to MR physics-based descriptions of quantitative tissue maps and acquisition-related factors.
-Hence, a further level of interpretability could be achieved by constraining the learned content to represent physically meaningful tissue parameters.
-Content discrepancy is characterized as a limiting factor in guided reconstruction (Section III-C) and plays a major role throughout this work.
-Viewed as a set of quasi tissue maps, estimating content from a single image is challenging, ideally requiring a time series of images, which captures the magnetization dynamics.
-Content discrepancy is, therefore, inevitable.
+VI Discussion
+In this work, we modeled the two-contrast MR image data in terms of two latent generative factors – content representing the contrast-independent structure and style representing the contrast-specific variations.
+Our data-driven definitions of content and style, though seemingly related to the MR physics-based representations of quantitative maps and acquisition-related factors, are rather nebulous compared to precisely defined physical concepts.
+For instance, MR quantitative maps are theoretically objective representations of tissues in terms of physical variables such as relaxation times and are ideally independent of the contrast-generating sequence parameters.
+On the other hand, our content and style representations are only defined statistically in terms of the given two contrasts and are, thus, neither as objective nor as independent in the ideal physical sense.
+That being said, we would argue that there is potential value in augmenting the data-driven content/style model with MR physical models, e.g.
+by constraining the learned content to represent physically meaningful anatomical properties, introducing elements of a physical model (e.g.
+Bloch equation) into the decoder network, etc., to enhance interpretability and reconstruction quality.
+Optimal content capacity was defined as a quantity representing the amount of shared local information contained in a two-contrast dataset.
+In terms of this quantity, we analyzed the effect of reference image resolution on reconstruction quality, obtaining a lowerbound for PnP-MUNIT to be effective (Section V-A2).
+This quantity would depend on more fundamental factors such as MR sequence types, which we did not investigate here.
+While we limited our experiments to T1W and T2W sequences, PnP-MUNIT is applicable to any pair of contrasts, with the image quality limited mainly by the amount of structural information shared between them.
+Future work could help empirically determine which sequence pairs are more amenable to guided reconstruction than others.
+Note that for a different pair of contrasts, say PD-weighted and T2W, a fresh content/style model needs to be trained since content and style are defined in the context of the specific contrast pair.
+It is, in principle, also possible to model the content and style for more than two contrasts.
+Given number of contrasts $N$, content is defined as the local structure underlying all $N$ contrasts, while style encodes intra-contrast variations of each individual contrast.
+This model would include $N$ sets of encoders and decoders and would require an (unpaired) image-domain training dataset of these $N$ contrasts.
+Subsequently, our PnP-MUNIT framework can be extended to use multiple reference contrasts to guide a given target contrast.
+The specific question to investigate then would be – given $P$ reference contrasts (where $1<P<N$), how to aggregate their contents such that the prior structural information about the target image is maximized?
+Content discrepancy was characterized as a quantity representing the gap between the true content of the target image and our estimation of this content, which is another limiting factor in PnP-MUNIT reconstruction.
+In MUNIT, the decoders model the (forward) generative process that maps the latent variables (i.e.
+content and style) to the observable variable (i.e.
+images).
+Ideally, the encoders must be a perfect inverse of this forward process.
+However, estimating the multi-channel contrast-independent content from a single image (using the content encoders) is a challenging (and perhaps ill-conditioned) inverse problem.
+Hence, some errors are to be expected in the content estimated from the reference image.
+Moreover, the reference image may contain unfavorable differences from the target image, such as misalignment or artifacts, which introduce more errors.
+Therefore, content discrepancy is expected to be a non-zero quantity.
 We proposed two complementary ways to minimize its effect on the reconstruction.
-While paired fine-tuning improves the model using a small amount of paired data, content refinement (CR) corrects for the remaining discrepancy adaptively during reconstruction using the measured k-space.
-A drawback of CR, however, is slower convergence (Section IV-A2).
-This could be mitigated by partially adopting an unrolled network design.
-The resulting tradeoff between model latency and generalizability would be interesting to explore.
-Unpaired pre-training of the content/style model boosts the practical applicability of PnP-MUNIT, e.g.
-in the case of LUMC-TRA where the T1W/T2W data imbalance is considerable.
-In this work, we limited our baselines to those algorithms that closely resemble ours in terms of design or data requirement.
-As future work, a more comprehensive study could be conducted with a broader range of methods and evaluation criteria, e.g.
-comparing generalizability across accelerations and sampling patterns with guided unrolled networks, given a fixed budget of paired training data.
-At reconstruction time, we assumed the reference and reconstruction images to be spatially aligned.
-The CR update, which can correct minor registration errors implicitly, would likely break down at the typically observed levels of patient motion.
-A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further relaxing the overall data constraints.
-In terms of reconstruction quality, we observed (Section IV-B) that using both measured k-space and reference scan via PnP-MUNIT was more beneficial than (a) using either one of them and (b) combining both in a naive way or based on hand-crafted prior, suggesting that our approach maximally exploits the complementary information.
-Compared to PnP-CNN, PnP-MUNIT allowed up to 32.6% more acceleration for a given SSIM.
-[^1] Based on linear interpolation of LUMC-TRA SSIM, PnP-MUNIT and PnP-CNN allowed $R$=10 and $R$=6.7, respectively, the maximum difference in $R$, at median SSIM=0.906.
-Hence, the maximum time saving 32.6%.
-In the radiological task (Section IV-C), PnP-MUNIT produced diagnostic-quality reconstructions at $R$=3, a difference of 33.3% over the clinical reconstructions of $R$=2.222We define diagnostic quality as visibility of at least 90% of the metastases and scoring at least “fairly diagnostic quality” on all other criteria.
-In clinical practice, while 10% of the metastases may not be detected in the reconstructed T2W scan, they would be prominently visible in the corresponding gadolinium-enhanced T1W scan, and hence would not go undetected.
-However, it was somewhat worse than PnP-CNN, while theoretically, the additional guidance should result in at least equivalent image quality.
-We argue that the likely cause is the limitations of the dataset and the specific content/style model, not of the proposed general approach.
-First, in LUMC-TRA, T1W scans had 2-3 times lower resolution than T2W scans, limiting the value of guidance (Section IV-A1).
-In the long run, this may be solved by co-designing a higher-resolution reference sequence with the target sequence to optimize guidance.
-Second, and more importantly, correcting for content discrepancy using CR relies on computing gradients through the decoder $G_{2}$.
+While paired fine-tuning improves the model using a small amount of paired image-domain training data, content refinement (CR) aims to correct the remaining discrepancy during reconstruction using the measured k-space.
+A drawback of the CR module, however, is slower convergence (Section V-A3).
+For example, the run-time for PnP-MUNIT on the GPU was around 17 seconds for a slice of matrix size 349$\times$284, as compared to 6 seconds for CS-WT, 1.4 seconds for PnP-CNN, and less than 0.1 second for MC-VarNet.
+This latency issue could be mitigated by approximating the iterative process with an unrolled network design, although perhaps at a cost of the generalizability offered by the plug-and-play design.
+The resulting trade-off between model latency and generalizability would be interesting to explore in the future.
+In terms of training data requirements, PnP-MUNIT can make use of the larger amounts of the image data available and can be applied to situations where end-to-end methods are infeasible since it does not rely on k-space training data.
+Moreover, unpaired image-domain pre-training of the content/style model boosts the practical applicability of PnP-MUNIT, e.g.
+in the case of LUMC-TRA where the T1W/T2W data imbalance was considerable.
+As indicated in our NYU DICOM benchmark, PnP-MUNIT performs similarly or better than state-of-the-art end-to-end methods.
+In our LUMC multi-coil benchmark, we limited our baselines to those algorithms that were feasible given the data constraint.
+As future work, a more comprehensive study should be conducted with a broader range of methods and evaluation criteria, e.g.
+comparing generalizability across accelerations and sampling patterns with multi-contrast unrolled networks, given a fixed budget of paired training data.
+Contrary to the training stage, we strictly assumed the spatial alignment of the reference and target images in the reconstruction stage.
+The CR update, which can implicitly correct minor registration errors, would likely break down at the typically observed levels of patient motion.
+A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further improving practical applicability.
+In terms of reconstruction quality, we observed in Section V-C that the true added value of PnP-MUNIT was at the high acceleration factors ($R$=8 and $R$=10) where the k-space data is scarce and the reference information becomes more valuable.
+However, the risk of model hallucinations increases at these accelerations, raising concerns about the accuracy of the anatomy represented in the image.
+Given the advantages of content/style decomposition, it may be possible to leverage the contrast-independence property of the content to automatically detect local hallucinations in the reconstruction and to define a corrective process to minimize it.
+This is another topic for future work.
+At lower accelerations, on the other hand, we observed that PnP-MUNIT often produced slightly lower metrics compared to PnP-CNN, e.g.
+in LUMC-COR.
+This was also observed in the radiological evaluation on LUMC-TRA samples.
+This is counter-intuitive at first glance since PnP-MUNIT has access to additional side-information and should, in principle, perform at least as good as PnP-CNN.
+However, this can be explained by the fact that at lower accelerations, the content/style model becomes the limiting factor for reconstruction quality.
+This is because unlike the CNN denoiser (Eq.
+(4)) of PnP-CNN, the content consistency operator (Eq.
+(14)) of PnP-MUNIT radically changes the image and hence, this operation is sensitive to the model’s overall performance.
+The model performance is in turn influenced by the training dataset and the network architecture.
+First, regarding data, the T1W scans in LUMC datasets had an in-plane resolution 1.3-2.5 times lower than that of the T2W scans, limiting the quality of the model (as indicated by its low optimal content capacity) and the overall value of guidance.
+In the long run, this may be solved by using 3D reference and target sequences of matched-resolution to optimize guidance.
+Second, regarding model design, correcting for content discrepancy using the CR module relies on computing gradients through the decoder $G_{2}$ (Eq.
+(16)).
 If $G_{2}$ is highly non-linear and the reference content is not sufficiently close to the optimum, CR may converge to a sub-optimal point leading to texture artifacts in the reconstruction, as seen in Fig.
 10 at $R$=4.
-In the radiological study, precisely this phenomenon is likely to have impacted the visual and diagnostic quality of PnP-MUNIT.
-Hence, to improve clinical utility, the content/style model must be improved using specialized architecture and regularization.
-VI Conclusion
-In this work, we introduced PnP-MUNIT, a novel approach to multi-contrast reconstruction combining content/style modeling with iterative reconstruction, which is interpretable in that it allows the two stages to be analyzed separately.
-By exploring the interpretability aspect, two limiting factors of guided reconstruction were identified – the amount of effective content in a two-contrast dataset and content discrepancy.
-On real-world clinical data, PnP-MUNIT provided up to 32.6% more acceleration over PnP-CNN for a given SSIM, enabling sharper reconstructions at high accelerations.
-Given the radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-MUNIT produced diagnostic-quality images at $R$=3, enabling 33.3% more acceleration over clinical reconstructions.
-To progress towards a practical implementation, future work will focus on improving the content/style model and reconstruction latency and incorporating online registration into the reconstruction.
+In our proof-of-concept, we used the same general model architecture proposed in [22].
+Better designs that are more conducive to CR may exist.
+Hence, future efforts must considerably focus on improving the content/style model using specialized architecture and regularization.
+VII Conclusion
+In this work, we introduced PnP-MUNIT, a modular approach to multi-contrast reconstruction combining content/style modeling with iterative reconstruction, which offers the reconstruction quality of end-to-end methods under stronger training data constraints.
+At its core is the content consistency operation, which provides regularization at the level of the image’s semantic content.
+We defined two quantities that determine the efficacy of this operation, namely optimal content capacity and content discrepancy, and provided several ways of maximizing this efficacy.
+On real-world clinical data, PnP-MUNIT provided up to 32.6% more acceleration over PnP-CNN for given SSIM, enabling sharper reconstructions at high accelerations.
+In the radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-MUNIT produced diagnostic-quality images at $R$=3, enabling 33.3% more acceleration over clinical reconstructions.
+To progress towards practical implementation of our proof-of-concept, future work will focus mainly on improving the content/style model and reconstruction latency, tackling the problem of model hallucinations, and incorporating online registration into the reconstruction.

### v3

@@
 Abstract
-Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can be used as a prior for guiding the reconstruction of an undersampled subsequent contrast.
-To this end, several learning-based guided reconstruction methods have been proposed.
+Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can guide the reconstruction of an undersampled subsequent contrast.
+To this end, several end-to-end learning-based guided reconstruction methods have been proposed.
 However, a key challenge is the requirement of large paired training datasets comprising raw data and aligned reference images.
-We propose a modular two-stage approach for guided reconstruction addressing this issue, which additionally provides an explanatory framework for the multi-contrast problem in terms of the shared and non-shared generative factors underlying two given contrasts.
+We propose a modular two-stage approach addressing this issue, additionally providing an explanatory framework for the multi-contrast problem based on the shared and non-shared generative factors underlying two given contrasts.
 A content/style model of two-contrast image data is learned from a largely unpaired image-domain dataset and is subsequently applied as a plug-and-play operator in iterative reconstruction.
 The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors.
-Based on this, incorporating prior information into the reconstruction reduces to simply replacing the aliased content of the image estimate with high-quality content derived from the reference scan.
+Consequently, incorporating prior information into the reconstruction reduces to a simple replacement of the aliased content of the reconstruction iterate with high-quality content derived from the reference scan.
 Combining this component with a data consistency step and introducing a general corrective process for the content yields an iterative scheme.
-We name this novel approach PnP-MUNIT.
+We name this novel approach PnP-CoSMo.
 Various aspects like interpretability and convergence are explored via simulations.
-Furthermore, its practicality is demonstrated on the NYU fastMRI DICOM dataset and two in-house multi-coil raw datasets, obtaining up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM.
-In a radiological task, PnP-MUNIT allowed 33.3% more acceleration over clinical reconstruction at diagnostic quality.
-I Introduction
+Furthermore, its practicality is demonstrated on the NYU fastMRI DICOM dataset, showing improved generalizability compared to end-to-end methods, and on two in-house multi-coil raw datasets, offering up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM.
+In a small radiological task, PnP-CoSMo allowed 33.3% more acceleration over clinical reconstruction at diagnostic quality.
+keywords:
+MRI reconstruction , multi-modal , content/style decomposition , plug-and-play
+journal: Medical Image Analysis
+[…]
+[…]
+1 Introduction
 Magnetic resonance imaging (MRI) is an invaluable medical imaging modality due to the high-quality scans it delivers, the variety of complementary information it can capture, and its lack of radiation-related risks, leading to its wide usage in clinical practice.
@@
 The raw sensor data is acquired in the frequency domain, or k-space, from which the image is reconstructed.
-Over the last 25 years, advancements such as parallel imaging [1, 2], compressed sensing (CS) [3], and deep learning reconstruction [4, 5] have enabled considerable speedups by allowing sub-Nyquist k-space sampling and relying on computationally sophisticated reconstruction.
-These techniques have subsequently been implemented on commercial MRI systems and have demonstrated (potential) improvements in clinical workflow [6].
+Over the last 25 years, advancements such as parallel imaging (Pruessmann et al., 2001; Griswold et al., 2002), compressed sensing (CS) (Lustig et al., 2007), and deep learning reconstruction (Aggarwal et al., 2018; Pezzotti et al., 2020) have enabled considerable speedups by allowing sub-Nyquist k-space sampling and relying on computationally sophisticated reconstruction.
+These techniques have subsequently been implemented on commercial MRI systems and have demonstrated (potential) improvements in clinical workflow (Seow et al., 2024).
 A clinical MRI session typically involves the acquisition of multiple scans of the same anatomy through the application of different MR pulse sequences, exhibiting different contrasts.
@@
 On the reconstruction side, multi-contrast methods have addressed this problem by introducing the shared information into the reconstruction phase to allow higher levels of k-space undersampling.
-In the simplest case of two contrasts, multi-contrast reconstruction can be classified into two types – (a) guided reconstruction, where an existing high-quality reference scan is used to guide the reconstruction of an undersampled second scan [7, 8, 9] and (b) joint reconstruction, where both contrasts are undersampled and are reconstructed simultaneously [10, 11, 12].
+In the simplest case of two contrasts, multi-contrast reconstruction can be classified into two types – (a) guided reconstruction, where an existing high-quality reference scan is used to guide the reconstruction of an undersampled second scan (Ehrhardt and Betcke, 2016; Weizman et al., 2016; Zhou and Zhou, 2020) and (b) joint reconstruction, where both contrasts are undersampled and are reconstructed simultaneously (Bilgic et al., 2011; Huang et al., 2014; Kopanoglu et al., 2020).
 In this work, we consider the problem of guided reconstruction, assuming no inter-scan motion between reference and target scans.
@@
 (a) The first stage learns a content/style model of two-contrast MR image data.
-The two image domains $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ are decomposed into a shared content domain $\mathcal{C}$ and separate style domains $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$.
-This illustration only provides an intuition about the information encoded in the content and style domains, and does not represent the actual training process, which is given in detail in Section III-B.
-(b) The reconstruction stage applies this model as a content consistency operator (bottom) within an ISTA-based iterative algorithm.
+We assume that the two image domains $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ can be decomposed into a shared content domain $\mathcal{C}$ and separate style domains $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$.
+This model is learned in two stages – an unpaired pre-training stage and a paired fine-tuning stage, both requiring only image data.
+(b) The reconstruction stage applies the content/style model as a content consistency operator (bottom) within an ISTA-based iterative scheme.
 Given an aligned reference image, guidance is introduced into the reconstruction by simply replacing its aliased content with content derived from the reference.
-CR denotes a content refinement update, which iteratively corrects for inconsistencies between the reference content and the measured k-space data, improving the effectiveness of the content consistency operator.
+The ”refine” block denotes a content refinement update, which iteratively corrects for inconsistencies between the reference content and the measured k-space data, improving the effectiveness of the content consistency operator.
 DC denotes data consistency.
-Wavelet-domain soft-thresholding (top) and CNN-based denoising (middle) used in CS-WT and PnP-CNN reconstruction, respectively, are shown for comparison.
+For comparison, two other reconstruction priors are shown, namely wavelet-domain soft-thresholding (top) and CNN-based denoising (middle) used in CS and PnP-CNN algorithms, respectively.
 The guided reconstruction problem entails using the local structure of the reference scan as a prior to complement the undersampled k-space measurements of the target scan.
-This problem has been formulated in different ways, ranging from conventional CS [7, 8] to end-to-end learning with unrolled networks [9, 13, 14, 15] and, more recently, diffusion model-based Bayesian maximum a posteriori estimation [16].
+This problem has been formulated in different ways, ranging from conventional CS (Ehrhardt and Betcke, 2016; Weizman et al., 2016) to end-to-end learning with unrolled networks (Zhou and Zhou, 2020; Yang et al., 2020; Pooja et al., 2022; Lei et al., 2023) and, more recently, diffusion model-based Bayesian maximum a posteriori estimation (Levac et al., 2023).
 Most end-to-end approaches, although more powerful than earlier hand-crafted ones, suffer from the main drawback of requiring large paired training datasets consisting of the target image and its k-space together with an aligned reference image, thereby limiting their application on real-world MR data.
@@
 To the best of our knowledge, ours is the first work that thoroughly explores content/style-based generative modeling for multi-contrast MRI reconstruction.
-In recent years, image-to-image translation has found application in the direct estimation of one MR contrast from another [17, 18, 19, 20].
+In recent years, image-to-image translation has found application in the direct estimation of one MR contrast from another (Dar et al., 2019; Yurt et al., 2021; Denck et al., 2021; Atli et al., 2024).
 While these methods are attractive due to their lack of dependence on k-space data, viewed in the light of MR physics, cross-contrast translation takes an extreme stance by not explicitly taking into account contrast-specific sensitivities and relying solely on the prior contrast.
 Hence, in the context of MR image formation, cross-contrast translation, by itself, can only provide a part of the information about the target image as data acquisition via MR sequences is necessary to obtain new information about the anatomy.
-That being said, literature on unpaired image translation provides a repertoire of useful tools such as joint generative modeling of two-domain image data [21, 22], which can be adopted to complement image reconstruction.
-We observe that methods such as MUNIT [22] can be applied to learn semantically meaningful representations of contrast-independent and contrast-specific information as content and style, respectively, without the need for paired image-domain training data.
+That being said, literature on unpaired image translation provides a repertoire of useful tools such as joint generative modeling of two-domain image data (Liu et al., 2017; Huang et al., 2018), which can be adopted to complement image reconstruction.
+We observe that methods such as MUNIT (Huang et al., 2018) can be applied to learn semantically meaningful representations of contrast-independent and contrast-specific information as content and style, respectively, without the need for paired image-domain training data.
 Plug-and-play (PnP) methods are an emerging paradigm for solving inverse problems in computational imaging.
-The main research line [23, 24] has focused on learning CNN-based denoising models on image-domain data and applying them as functions replacing proximal operators in iterative algorithms like ISTA and ADMM, demonstrating improved image recovery.
+The main research line (Ahmad et al., 2020; Kamilov et al., 2023) has focused on learning CNN-based denoising models on image-domain data and applying them as functions replacing proximal operators in iterative algorithms like ISTA and ADMM, demonstrating improved image recovery.
 An advantage of this approach is the decoupling of the learning problem of image modeling from the inverse problem of image reconstruction, thereby simplifying model training and improving generalizability across different acceleration factors, undersampling patterns, etc.
@@
 Since the content of the reconstructing image, which is the remaining piece of information and is contrast-independent, is supplied by the reference image, one can compose a de-aliased estimate of the reconstruction from the undersampled image in a single step.
-We term this the content consistency operation, which forms the basis of our iterative reconstruction algorithm PnP-MUNIT [25].
+We term this the content consistency operation, which forms the basis of our iterative reconstruction algorithm PnP-CoSMo (Plug-and-Play method based on Content/Style Modeling) (Rao et al., 2024).
 An overview of our approach is shown in Fig.
@@
   3. 3.
-    Developing this idea further and incorporating a corrective process for the content, we propose PnP-MUNIT, a modular algorithm for guided reconstruction, combining the flexibility of the plug-and-play approach with the semantic interpretability of content/style decomposition.
+    Developing this idea further and incorporating a corrective process for the content, we propose PnP-CoSMo, a modular algorithm for guided reconstruction combining the flexibility of the plug-and-play approach with the semantic interpretability of content/style decomposition.
   4. 4.
-    Through comprehensive experiments, we shed light on several properties of PnP-MUNIT such as convergence and robustness and demonstrate its applicability on real-world raw data and its potential clinical utility for a specific radiological task.
-II Related Work
-II-A Reconstruction Methods for Accelerated MRI
-Techniques for accelerating MRI by k-space undersampling go back to compressed sensing (CS) [3], which combines random sampling with sparsity-based iterative denoising, most commonly implemented based on the ISTA [26] or ADMM [27] family of algorithms.
+    Through comprehensive experiments, we shed light on several properties of PnP-CoSMo such as convergence and robustness and demonstrate its applicability on real-world raw data and its potential clinical utility with a specific radiological task.
+2 Related work
+2.1 Reconstruction methods for accelerated MRI
+Techniques for accelerating MRI by k-space undersampling go back to compressed sensing (CS) (Lustig et al., 2007), which combines random sampling with sparsity-based iterative denoising, most commonly implemented based on the ISTA (Chambolle et al., 1998) or ADMM ((Boyd et al., 2011)) family of algorithms.
 Most modern deep learning-based reconstruction methods focus on improving the denoising part.
-Plug-and-play (PnP) methods [23, 24] replace the proximal operator in ISTA and ADMM with off-the-shelf denoisers such as a learned convolutional denoising model.
-Unrolled networks [28, 4, 5] extend this idea by casting the entire iterative algorithm into one large network, trained end-to-end.
-This makes them more adaptive to factors such as sampling pattern and acceleration, although at a cost of generalizability [23].
-One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke [7] introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on the edge features in the reference scan.
+Plug-and-play (PnP) methods (Ahmad et al., 2020; Kamilov et al., 2023) replace the citep operator in ISTA and ADMM with off-the-shelf denoisers such as a learned convolutional denoising model.
+Unrolled networks (Schlemper et al., 2017; Aggarwal et al., 2018; Pezzotti et al., 2020) extend this idea by casting the entire iterative algorithm into one large network, trained end-to-end.
+This makes them more adaptive to factors such as sampling pattern and acceleration, although at a cost of generalizability (Ahmad et al., 2020).
+One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke (2016) introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on the edge features in the reference scan.
 Later work introduced adaptive elements into the multi-contrast CS framework, e.g.
 Weizman et al.
-[8] proposed adaptive weighting-based guided CS and Song et al.
-[29] used adaptive sparse domains based on coupled dictionary learning.
+(2016) proposed adaptive weighting-based guided CS and Pingfan et al.
+(2018) used adaptive sparse domains based on coupled dictionary learning.
 In the latter, the problem was formulated using a patch-level linear model comprising coupled and distinct sparse dictionary representations of the two contrasts.
@@
 A general drawback of classical methods compared to deep learning-based ones is their lower flexibility.
-End-to-end learning-based methods [9, 13, 30, 31, 32, 15], on the other hand, supply the reference scan as an additional input to a deep reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proven to be more effective than conventional algorithms.
+End-to-end learning-based methods (Zhou and Zhou, 2020; Yang et al., 2020; Liu et al., 2021a, b; Feng et al., 2022; Lei et al., 2023), on the other hand, supply the reference scan as an additional input to a deep reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proven to be more effective than conventional algorithms.
 However, end-to-end methods require large paired training datasets comprising the ground-truth image and the k-space together with aligned reference images, which is too strong a constraint when working with retrospectively collected clinical data that reflects the natural inconsistencies of routine practice such as missing contrasts, differences in spatial resolutions, and the existence of inter-scan motion.
 Additionally, by relying on end-to-end-learned features, these methods are generally less interpretable than their conventional counterparts in that they do not explicitly model the multi-contrast problem in terms of the underlying shared and non-shared information.
-MC-VarNet [15] is an important exception, which uses a simple linear decomposition of the reference contrast into common and unique components, applying the common component for guidance.
+MC-VarNet (Lei et al., 2023) is an important exception, which uses a simple linear decomposition of the reference contrast into common and unique components, applying the common component for guidance.
 A reconstruction method that is as effective as the learning-based methods while having more lenient data requirements and offering a high degree of interpretability is still needed.
 Our plug-and-play method relies on learning a deep non-linear content/style transform from image-domain data only, even if subject-wise paired images are not fully available.
-II-B Unpaired Image-to-Image Modeling
-Image-to-image modeling is the general problem of learning a mapping between two image domains and was first addressed by Pix2Pix [33] and CycleGAN [34] in paired and unpaired settings, respectively.
-Another line of unpaired image translation methods, the first of which was UNIT [21], assumes a shared latent space underlying the two domains to explicitly represent shared information.
+2.2 Unpaired image-to-image modeling
+Image-to-image modeling is the general problem of learning a mapping between two image domains and was first addressed by Pix2Pix (Isola et al., 2017) and CycleGAN (Zhu et al., 2017) in paired and unpaired settings, respectively.
+Another line of unpaired image translation methods, the first of which was UNIT (Liu et al., 2017), assumes a shared latent space underlying the two domains to explicitly represent shared information.
 However, both CycleGAN and UNIT assume a deterministic one-to-one mapping between the two domains, ignoring the fact that an image in one domain can have multiple valid renderings in the other.
-Deterministic image translation has been widely applied to MRI contrast-to-contrast synthesis [17, 35, 18].
+Deterministic image translation has been widely applied to MRI contrast-to-contrast synthesis (Dar et al., 2019; Oh et al., 2020; Yurt et al., 2021).
 However, fundamentally, these methods do not account for the variability of the scanning setup that influences the realized contrast level and the differential visibility of pathologies in the target image.
 Denck et al.
-[19] partially address this problem by proposing contrast-aware MR image translation where the acquisition sequence parameters are fed into the network to control the output’s contrast level.
+(2021) partially address this problem by proposing contrast-aware MR image translation where the acquisition sequence parameters are fed into the network to control the output’s contrast level.
 However, this model is too restrictive since it assumes a single pre-defined mode of variability (i.e.
@@
 sequence parameters) must be available.
-MUNIT [22] extended UNIT by modeling domain-specific variability in addition to domain-independent structure, enabling many-to-many mapping and thus overcoming the rigidity of UNIT.
-The result was a stochastic image translation model which, given an input image, generates a distribution of synthetic images sharing the same “content” but differing in “style”.
+MUNIT (Huang et al., 2018) extended UNIT by modeling domain-specific variability in addition to domain-independent structure, enabling many-to-many mapping and thus overcoming the rigidity of UNIT.
+The result was a stochastic image translation model which, given an input image, generates a distribution of synthetic images sharing the same ”content” but differing in ”style”.
 More fundamentally, MUNIT is a learned invertible transformation between the image domains and the disentangled content/style domains.
-And unlike other content/style modeling frameworks such as [36, 37, 38], MUNIT models content and style as latent generative factors of the two image domains, providing precise distribution-level definitions for them.
+And unlike other content/style modeling frameworks such as (Kotovenko et al., 2019; Kwon and Ye, 2021; Wu et al., 2023), MUNIT models content and style as latent generative factors of the two image domains, providing precise distribution-level definitions for them.
 Clinical MR images of a given protocol contain multiple modes of variability, many of which are not known a priori.
-Compared to [19], we make a broader assumption that the contrast-independent semantic information is local in nature and that the contrast-specific variations in the dataset can include global effects of acquisition settings as well as local anatomical features unique to the contrast.
-We model these using MUNIT, referring to the shared and non-shared components as content and style, respectively.
-II-C Combining Image Translation with Reconstruction
+Compared to Denck et al.
+(2021), we make a broader assumption that the contrast-independent semantic information is local in nature and that the contrast-specific variations in the dataset can include global effects of acquisition settings as well as local anatomical features unique to the contrast.
+We model using an approach based on the MUNIT framework, referring to the shared and non-shared components as content and style, respectively.
+2.3 Combining image translation with reconstruction
 Acknowledging the limitation of MR cross-contrast prediction, some prior work has attempted combining it with multi-contrast reconstruction.
 A naive form of joint image synthesis and reconstruction, e.g.
-PROSIT [39] and rsGAN [40], involves generating a synthetic image from the reference scan via deterministic image translation and using it in a classical L2-regularized least-squares reconstruction.
+PROSIT (Mattern et al., 2020) and rsGAN (Dar et al., 2020), involves generating a synthetic image from the reference scan via deterministic image translation and using it in a classical L2-regularized least-squares reconstruction.
 More recent work by Xuan et al.
-[41] proposed a joint image translation and reconstruction method which additionally accounts for misalignment between the reference and reconstructing images, with a follow-up work leveraging optimal transport theory [42].
+(2022) proposed a joint image translation and reconstruction method which additionally accounts for misalignment between the reference and reconstructing images, with a follow-up work leveraging optimal transport theory (Wang et al., 2024).
 Levac et al.
-[16] formulate guided reconstruction as a highly general Bayesian maximum a posteriori estimation problem and solve it iteratively via Langevin update steps, using an image-domain diffusion model as the score function of the prior distribution.
+(2023) formulate guided reconstruction as a highly general Bayesian maximum a posteriori estimation problem and solve it iteratively via Langevin update steps, using an image-domain diffusion model as the score function of the prior distribution.
 While these are promising directions, we propose an alternative approach that decomposes the multi-contrast problem in a more intuitive way – first, into two sub-problems, namely image modeling and image reconstruction; and second, within the image-domain model, the multi-contrast information is decomposed into content and style.
-This two-level decomposition results in a highly modular reconstruction algorithm with a built-in explanatory framework where (a) the guidance mechanism is a simple content-replacement operation, (b) the discrepancy between the supplied reference content and the true content of the target image represents a meaningful error term which can be minimized, and (c) the optimal content-encoding capacity of the model for a given two-contrast image dataset indicates the amount of shared structure available to be learned in this data and used in the reconstruction task.
-III Methods
-III-A Reconstruction Problem
-III-A1 Undersampled MRI reconstruction
+This two-level decomposition results in a highly modular reconstruction algorithm with a built-in explanatory framework where (a) the guidance mechanism is a simple content-replacement operation, (b) the discrepancy between the supplied reference content and the true content of the target image represents a meaningful error term which can be minimized, and (c) the optimal content-encoding capacity of the model for a given two-contrast image dataset indicates the amount of shared structure that is available to be learned in this data and utilized in the reconstruction task.
+3 Methods
+3.1 Reconstruction problem
+3.1.1 Undersampled MRI reconstruction
 Given a set of $P$ acquired k-space samples $y\in\mathbb{C}^{P}$ and the MRI forward operator $A\in\mathbb{C}^{P\times Q}$, CS reconstruction of the image $x\in\mathbb{C}^{Q}$ with $Q$ voxels is given as
@@
 (2) performs soft-thresholding in the transform domain, thereby reducing the incoherent undersampling artifacts in image $x^{k-1}$, whereas (3) enforces soft data consistency on image $r^{k}$ by taking a single gradient descent step over the least-squares term, controlled by step size $\eta$.
-III-A2 Plug-and-play denoiser
+3.1.2 Plug-and-play denoiser
 Plug-and-play methods replace the analytical operation of (2) with off-the-shelf denoisers.
@@
 Given a CNN model $f_{\theta}$ trained to remove i.i.d.
-Gaussian noise from an image, PnP-CNN [23] modifies Eq.
+Gaussian noise from an image, PnP-CNN (Ahmad et al., 2020) modifies Eq.
 (2) to
-$r^{k}\leftarrow f_{\theta}(x^{k-1}).$ (4)
-III-A3 Plug-and-play content consistency operator
+$z^{k}\leftarrow f_{\theta}(x^{k-1}).$ (4)
+3.1.3 Plug-and-play content consistency operator
 In guided reconstruction, a spatially aligned reference $x^{\mathrm{ref}}_{1}$ is available, which captures the same underlying semantic content as the target reconstruction.
@@
 We propose a content consistency operator $g_{M}(\cdot;c)$ such that
-$r_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c),$ (5)
-where a content/style model $M$ decomposes $x_{2}^{k-1}$ into content and style, followed by a replacement of this corrupted content with high-quality content $c$ derived from $x^{\mathrm{ref}}_{1}$ and composing the improved image $r_{2}^{k}$ from it.
-Before formally defining this operator and developing the reconstruction algorithm, we discuss the steps required to learn the content/style model $M$.
-III-B Content/Style Modeling
+$z_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c),$ (5)
+where a content/style model $M$ decomposes $x_{2}^{k-1}$ into content and style, followed by a replacement of this corrupted content with high-quality content $c$ derived from $x^{\mathrm{ref}}_{1}$ and composing the improved image $z_{2}^{k}$ from it.
+Before formally defining this operator and developing the reconstruction algorithm, we discuss in the following section the design and learning of the content/style model $M$.
+3.2 Content/Style modeling
 Given an image dataset of two MR contrasts we make the general assumption that there exists an underlying contrast-independent structure which, influenced by arbitrary contrast-specific factors, is rendered as the contrast images.
-III-B1 Unpaired pre-training of MUNIT
-We formulate our content/style model based on the MUNIT framework [22].
+We formulate our content/style model based on the MUNIT framework (Huang et al., 2018).
+3.2.1 Unpaired MUNIT pre-training
 We define an image domain as the set $\mathcal{X}_{i}$ of images of a certain contrast comprising the dataset, where $i\in\{1,2\}$.
-“Content” $c\in\mathcal{C}$ is defined here as the underlying contrast-independent structure and is represented as a set of feature maps, whereas “style” $s_{i}\in\mathcal{S}_{i}$ corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g.
+”Content” $c\in\mathcal{C}$ is defined as the underlying contrast-independent structure and is represented as a set of feature maps, whereas ”style” $s_{i}\in\mathcal{S}_{i}$ corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g.
 global effects of acquisition settings, contrast-specific tissue features, etc., and is represented as a low-dimensional vector.
@@
 Thus, the content/style model is specified as $M=\{E_{1}^{c},E_{2}^{c},E_{1}^{s},E_{2}^{s},G_{1},G_{2}\}$.
-The MUNIT loss function is comprised of 4 terms:
+The MUNIT loss function is comprised of four terms:
 $\mathcal{L_{\mathrm{MUNIT}}}=\mathcal{L}_{\mathrm{GAN}}+\alpha_{1}\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}+\alpha_{2}\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}+\alpha_{3}\mathcal{L}_{\mathrm{style}}^{\mathrm{self}},$ (6)
@@
 $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are hyperparameters.
-III-B2 Paired fine-tuning
-While the pre-training stage learns useful content/style representations, the model can be adapted to our task by improving its content preservation.
-To this end, we propose a paired fine-tuning (PFT) stage, leveraging a modest amount of paired data.
-Given samples from the joint distribution $p(x_{1},x_{2})$, our fine-tuning objective is given as
+For our purpose, we consider the MUNIT training process as the pre-training stage of our content/style model.
+3.2.2 Paired fine-tuning
+Figure 2: Our proposed paired fine-tuning (PFT) stage for improving the alignment of the content representations of the two domains given a small dataset of paired images.
+While the pre-training stage learns useful content/style representations, we adapt the model to our task by improving its content preservation.
+To this end, we propose a paired fine-tuning (PFT) stage which more strongly aligns the content representations of the two domains given a modest amount of paired data.
+Given samples from the joint distribution $p(x_{1},x_{2})$, our fine-tuning objective is expressed as
 $\mathcal{L}_{\mathrm{PFT}}=\mathcal{L}_{\mathrm{GAN}}+\beta_{1}\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}+\beta_{2}\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}+\beta_{3}\mathcal{L}_{\mathrm{content}}^{\mathrm{cross}},$ (11)
 where
-$\begin{split}\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}=\>&\mathbb{E}_{\{x_{1},x_{2}\}\sim p(x_{1},x_{2})}[\\ &||x_{2}-G_{2}(E_{1}^{c}(x_{1}),E_{2}^{s}(x_{2}))||_{1}+\\ &||x_{1}-G_{1}(E_{2}^{c}(x_{2}),E_{1}^{s}(x_{1}))||_{1}],\end{split}$ (12)
+$\begin{split}\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}=\mathbb{E}_{\{x_{1},x_{2}\}\sim p(x_{1},x_{2})}[\>&||x_{2}-G_{2}(E_{1}^{c}(x_{1}),E_{2}^{s}(x_{2}))||_{1}+\\ &||x_{1}-G_{1}(E_{2}^{c}(x_{2}),E_{1}^{s}(x_{1}))||_{1}],\end{split}$ (12)
 $\begin{split}\mathcal{L}_{\mathrm{content}}^{\mathrm{cross}}=\mathbb{E}_{\{x_{1},x_{2}\}\sim p(x_{1},x_{2})}[||E_{1}^{c}(x_{1})-E_{2}^{c}(x_{2})||_{1}].\end{split}$ (13)
@@
 $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are an additional set of hyperparameters.
-III-B3 Network architecture and content capacity
-Following the original paper [22], our content encoders $E_{i}^{c}$ consist of an input convolutional layer potentially followed by strided downsampling convolutions, and finally a series of residual blocks.
+Note that our PFT preserves the bi-directionality of the content/style model, a property which will later be essential for generalization over different reconstruction tasks.
+3.2.3 Network architecture and content capacity
+Following the original MUNIT paper (Huang et al., 2018), our content encoders $E_{i}^{c}$ consist of an input convolutional layer potentially followed by strided downsampling convolutions, and finally a series of residual blocks.
 Style encoders $E_{i}^{s}$ consist of input and downsampling convolutions followed by adaptive average pooling and a fully-connected layer that outputs the latent vector.
@@
 Style is introduced into the decoder via AdaIN operations which modulate the activation maps derived from the content.
-We observe that in this architecture, the ratio between the content and image resolutions reflects the level of local structure one expects to be shared between the two domains.
-This (relative) content resolution is thus an inductive bias built into the model, which we refer to as the model’s content capacity.
-This concept is closely related to the general concept of the locality bias of image-to-image models [43].
-A model with high content capacity has a large content resolution, allowing it to learn a rich content representation that strongly influences the output’s structure, which simultaneously restricts the style spaces to learn low-level global features.
-Figure 2: A trained content/style model $M$ and the content consistency operator $g_{M}(\cdot;c)$ based on it.
-Images shown here are from the NYU brain DICOM dataset.
-Practical details are provided in Section IV.
+We observe that in this architecture, the ratio between the content resolution and the image resolution reflects the level of local structure one expects to be shared between the two domains.
+This (relative) content resolution is thus an inductive bias built into the model, which we refer to as the model’s content capacity and define as
+$J_{M}:=\frac{H_{c}W_{c}}{H_{x}W_{x}},$ (14)
+where $H_{x}$ and $W_{x}$ are the height and width of the input image and $H_{c}$ ($\leq H_{x}$) and $W_{c}$ ($\leq W_{x}$) are the those of the corresponding content maps.
+This ratio is controlled via the number of up/downsampling blocks in the networks.
+Our notion of content capacity is closely related to the more general concept of the locality bias of image-to-image models (Richardson and Weiss, 2021).
+A content/style model with a high content capacity has a large content resolution, allowing it to learn a rich, highly expressive content representation that strongly influences the output’s structure.
+Consequently.
+the model’s style spaces are restricted to learn mostly low-level global features.
+Figure 3: Demonstration of the content consistency operator $g_{M}(\cdot;c)$ based on a content/style model $M$.
+The image pair shown here are NYU brain DICOMs, practical details regarding which are provided in Section 4.
 (a) A T1W/T2W image pair and the corresponding content maps.
-While these two contents generally agree, there is a notable discrepancy.
+While these two content maps generally agree, there is a notable discrepancy.
 (b) Synthetic T2W images generated from 16 style codes grid-sampled from $\mathcal{S}_{2}$, showing a smooth variation in image contrast and a roughly constant anatomical structure.
@@
 Contours in $\mathcal{S}_{2}$ indicate the MAE of the synthetic images from that region.
-The red circle indicate a style estimation NMSE of 0.1.
-III-C Iterative Reconstruction using Content Consistency
+The red circle indicates a style estimation NMSE of 0.1.
+3.3 Iterative reconstruction based on content consistency
 Given the content/style model $M$, we consider $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ as reference and target domains, respectively, in our guided reconstruction task, and define the content consistency operator $g_{M}(\cdot;c)$ from Eq.
@@
 Let $x_{2}^{*}$ be the ground truth reconstruction with content $c^{*}$ and style $s_{2}^{*}$.
-The success of our content consistency operation depends on two conditions – (a) $\hat{c}$ is close to $c^{*}$ and (b) $\hat{s}_{2}$ is close to $s_{2}^{*}$.
-The first condition is roughly satisfied, as seen in Fig.
-2a, because the model is explicitly trained to minimize content discrepancy, but more on this later.
-On the other hand, it is not obvious that the second condition should hold too and hence deserves a closer look.
-Assuming a high degree of shared local structure between the two domains, the optimal model has high content capacity (see Section III-B3).
-Thus, style would represent mostly low-level global image features, e.g.
-contrast variations, as indeed observe in Fig.
-2b.
+The success of our content consistency operation depends on three conditions – (a) $\hat{c}$ and $c^{*}$ are rich, high-resolution representations of the underlying anatomy, (b) $\hat{c}$ is close to $c^{*}$, and (c) $\hat{s}_{2}$ is close to $s_{2}^{*}$.
+The first condition is satisfied provided that there is sufficient amount of shared structure between the reference and target domains and that the model has high content capacity $J_{M}$ for capturing a highly expressive content.
+The second condition is roughly satisfied, as seen in Fig.
+3a, because the model is explicitly trained to minimize content error, but more on this later.
+On the other hand, it is not obvious that the third condition should hold too and it hence deserves a closer look.
+If the first condition is satisfied, the optimal model will have a high content capacity and consequently, style would represent mostly low-level global image features, e.g.
+contrast variations (see Section 3.2.3).
+This is observed in practice as shown in Fig.
+3b.
 It is a well known fact that image contrast is contained prominently in the center of the k-space.
-Hence, the estimate $\hat{s}_{2}$ can be made arbitrarily close to $s_{2}^{*}$ by sufficiently sampling the k-space center, as shown empirically in Fig.
-2c.
+Hence, the estimate $\hat{s}_{2}$ can be made arbitrarily close to $s_{2}^{*}$ by sufficiently sampling the k-space center, as demonstrated empirically in Fig.
+3c.
 Applying data consistency update (3) following $g_{M}(\cdot;\hat{c})$ in repetition yields an ISTA-based iterative scheme.
 Here, our content consistency update complements the data consistency update in the sense that while the latter forces the image estimate to be consistent with the given (measured) k-space data $y$, the former forces it to be consistent with the given (prior) content $\hat{c}$.
-Algorithm 1 PnP-MUNIT
+Algorithm 1 PnP-CoSMo iterative guided reconstruction
 1:$y$, $A$, $x^{\mathrm{ref}}_{1}$, $M$, $\eta$, $\gamma$triangleright where $M=\{E_{1}^{c},E_{1}^{s},E_{2}^{c},E_{2}^{s},G_{1},G_{2}\}$
@@
 6:     $k\leftarrow k+1$
-7:     $r_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c^{k-1})$triangleright Content consistency
-8:     $x_{2}^{k}\leftarrow r_{2}^{k}-\eta A^{H}(Ar_{2}^{k}-y)$triangleright Data consistency
-9:     $c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},E_{2}^{s}(x_{2}^{k}))-y||_{2}^{2}$triangleright CR
+7:     $z_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c^{k-1})$triangleright Content consistency update
+8:     $x_{2}^{k}\leftarrow z_{2}^{k}-\eta A^{H}(Az_{2}^{k}-y)$triangleright Data consistency update
+9:     $c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},E_{2}^{s}(x_{2}^{k}))-y||_{2}^{2}$triangleright Content refinement (CR) update
 10:until convergence
 11:return $x_{2}^{k}$
-The core assumption of our idealized content/style model is that contents $\hat{c}$ and $c^{*}$ are identical.
+The core assumption of an idealized content/style model is that contents $\hat{c}$ and $c^{*}$ are identical.
 However, in reality, this assumption will not hold, and a discrepancy between the two contents is to be expected.
@@
 fundamental limits such as irreducible error [^1] If perfectly zero error was possible, one could perfectly predict the (distribution of) target image from the reference without any measurement.
-and practical issues like sub-optimal training, and (b) reference image-related, e.g.
+and practical issues like sub-optimal network design and training, and (b) reference image-related, e.g.
 presence of artifacts independent of the target reconstruction.
-This content discrepancy, as observed empirically, is shown in Fig.
-2a.
-During reconstruction, the error in $\hat{c}$ would limit the efficacy of the operator $g_{M}(\cdot;\hat{c})$, affecting reconstruction quality.
-While model-related discrepancy is partly tackled by PFT (Section III-B2), we now propose a content refinement (CR) process to correct for the remaining discrepancy in the reconstruction stage.
+We refer to this gap between $\hat{c}$ and $c^{*}$ as content discrepancy.
+Fig.3a shows content discrepancy as observed empirically.
+During reconstruction, the error in $\hat{c}$ would limit the efficacy of the operator $g_{M}(\cdot;\hat{c})$, affecting the reconstruction quality.
+While model-related discrepancy is partly tackled by PFT (Section 3.2.2), we now propose a content refinement (CR) process to correct for the remaining discrepancy in the reconstruction stage.
 Since we have no direct access to the true content $c^{*}$, but only to undersampled k-space measurements $y$, we aim at solving the following minimization problem in CR
@@
 Hence, by aligning the content $c$ with k-space data $y$ and correcting the discrepancy, the CR module aligns content consistency updates with data consistency updates.
-With this additional component in place, we obtain our PnP-MUNIT reconstruction algorithm (Algorithm 1).
+With this additional component in place, we obtain our PnP-CoSMo reconstruction algorithm (Algorithm 1).
 Note that unlike content, the style of the image estimate need not be explicitly aligned with the k-space.
-The data consistency update applied on the image implicitly “corrects” its style and as the image converges, its style would converge as well.
-IV Experimental Setup
-IV-A Experiment Design
-Without loss of generality, we considered the case of reconstructing T2W scans using T1W references and focused on head applications.
-We conducted 4 sets of experiments to comprehensively evaluate PnP-MUNIT – (a) simulations, (b) benchmark on NYU DICOM data, (c) benchmark on clinical multi-coil raw data, and (d) radiological evaluation.
-TABLE I: Overview of the sequences in the LUMC datasets.
+Since the style information is assumed to be derived largely from the k-space data, the data consistency update applied on the image implicitly ”corrects” its style.
+Thus, as the image converges, we hypothesize that style would converge as well.
+4 Experimental setup
+4.1 Experiment design
+We conducted 4 sets of experiments to comprehensively evaluate PnP-CoSMo – (a) simulations, (b) benchmark against end-to-end methods on NYU DICOM data, (c) benchmark and ablation on clinical multi-coil raw data, and (d) radiological evaluation.
+Without loss of generality, we considered the case of reconstructing T2W scans using T1W references and focused on head applications, except in the NYU benchmark where the reverse case was considered as well.
+4.1.1 Simulation
+The goal of our simulation experiments was to empirically test several properties of our algorithm under highly controlled settings.
+To this end, we used simulated T1W and T2W images based on BrainWeb phantoms (Collins et al., 1998).
+BrainWeb provides 20 anatomical models of normal brain, each comprised of fuzzy segmentation maps of 12 tissue types.
+The 20 volumes were first split into the model training (18), model validation (1), and reconstruction testing (1) subsets.
+T1W/T2W spin-echo scans were simulated using TE/TR values randomly sampled from realistic ranges.
+For testing the reconstruction, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling at various accelerations.
+We analyzed the effect of the content/style model’s disentanglement level and its content capacity on the reconstruction quality, the convergence of the PnP-CoSMo algorithm, and the effectiveness, robustness, and tuning of the CR module.
+4.1.2 Benchmark on the NYU DICOM dataset
+In the NYU benchmark, we compared PnP-CoSMo against end-to-end reconstruction methods.
+The question we sought to answer was whether or not PnP-CoSMo, which requires only image-domain training data, can outperform methods that additionally require k-space training data.
+Furthermore, end-to-end methods are task-specific where the reference and target contrast are fixed – e.g.
+T1W and T2W, respectively.
+On the other hand, the content/style model used in PnP-CoSMo does not possess this kind of uni-directionality.
+Hence, we sought to assess the generalizability of PnP-CoSMo across two different reconstruction tasks, namely T1W-guided T2W reconstruction and T2W-guided T1W reconstruction.
+The NYU brain dataset (Knoll et al., 2020; Zbontar et al., 1811) includes raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR.
+For multi-contrast experiments, however, the NYU raw dataset does not provide subject labels for pairing the contrasts.
+Therefore, instead of the NYU raw data, we used the NYU DICOMs in this study, following Xuan et al.
+(2022).
+A paired DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans.
+All T1W scans were rigid-registered with the corresponding T2W scans.
+Single-coil k-space data was simulated via Fourier transform with 1D Cartesian random sampling at $R\in\{2,3,4,5\}$ and added Gaussian noise of $\sigma=0.01\mathrm{max}(x_{2}^{*})$.
+The dataset was split into 4 subsets – model-train (200), model-val (27), recon-val (50), and recon-test (50).
+Images from the former two splits were used to pre-train, fine-tune, and validate the content/style model.
+End-to-end models were trained and validated on the aligned T1W, the simulated T2W k-space, and the T2W ground-truth from these two splits.
+The recon-val split was used to tune the PnP-CoSMo algorithm, and recon-test was the held-out test set.
+We benchmarked PnP-CoSMo against one well-known single-contrast network – MoDL (Aggarwal et al., 2018) – and two recent multi-contrast networks – MTrans (Feng et al., 2022) and MC-VarNet (Lei et al., 2023).
+These networks were evaluated on the aforementioned two reconstruction tasks, assessing both their in-distribution and out-of-distribution performance – e.g.
+for the T1W-guided T2W reconstruction task, in-distribution means that the end-to-end model was trained on this task, whereas out-of-distribution means the model was trained of the T2W-guided T1W reconstruction task.
+Table 1: Overview of the sequences in the LUMC datasets.
 Sequence Params LUMC-TRA LUMC-COR
@@
 Num slices 50 15
-TABLE II: LUMC data split. The splits were made at the subject-level, i.e. each subject belonged to exactly one subset. Splits containing only image-domain data are marked with I.
+Table 2: LUMC data split.
+The splits were made at the subject-level, i.e.
+each subject belonged to exactly one subset.
+Splits containing only image-domain data are marked with I.
+The recon-val and recon-test splits additionally include raw T2W multi-coil k-space and coil sensitivities.
 Dataset Split Subjects Sessions Scans (T1W / T2W)
@@
 Total 291 330 322 / 325
-IV-A1 Simulation
-The goal of our simulation experiments was to empirically test several properties of our algorithm under highly controlled settings.
-To this end, we used simulated T1W and T2W images based on BrainWeb phantoms [44].
-BrainWeb provides 20 anatomical models of normal brain, each comprised of fuzzy segmentation maps of 12 tissue types.
-The 20 volumes were first split in 18:1:1 ratio for model training, validation, and reconstruction testing, respectively.
-T1W/T2W spin-echo scans were simulated using TE/TR values randomly sampled from realistic ranges.
-For testing the reconstruction, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling at various accelerations.
-We analyzed the effect of the content/style model’s disentanglement level and its content capacity on the reconstruction quality, the convergence of the PnP-MUNIT algorithm, and the effectiveness, robustness, and tuning of the CR module.
-IV-A2 Benchmark on NYU DICOM dataset
-In the NYU benchmark, we compared PnP-MUNIT against end-to-end reconstruction methods.
-The question we sought to answer was whether or not PnP-MUNIT, which requires only image-domain training data, can outperform methods that additionally require k-space training data.
-The NYU brain dataset [45, 46] includes raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR.
-Following [41], a paired DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans.
-All T1W scans were rigid-registered with the corresponding T2W scans.
-Single-coil T2W k-space data was simulated via Fourier transform with 1D Cartesian random sampling at $R\in\{2,3,4,5\}$ and added Gaussian noise of $\sigma=0.01\mathrm{max}(x_{2}^{*})$.
-The dataset was split into 4 subsets – model-train (200), model-val (27), recon-val (50), and recon-test (50).
-Images from the former two splits were used to pre-train, fine-tune, and validate the MUNIT model.
-On the other hand, end-to-end models were trained and validated on the aligned T1W, the simulated T2W k-space, and the T2W ground-truth from these two splits.
-The recon-val split was used to tune the PnP-MUNIT algorithm, and recon-test was the held-out test set.
-We benchmarked PnP-MUNIT against one well-known single-contrast network – MoDL [4] – and two recent multi-contrast networks – MTrans [32] and MC-VarNet [15].
-IV-A3 Benchmark and ablation on LUMC multi-coil dataset
-In our benchmark on clinical multi-coil data, our goal was to test PnP-MUNIT on a constrained real-world problem where only the image-domain data was available for training.
+4.1.3 Benchmark and ablation on the LUMC multi-coil dataset
+In our benchmark on clinical multi-coil data, the goal was to test PnP-CoSMo on a constrained real-world problem where only the image-domain data was available for training.
 This is often the case, as in clinical practice raw data is discarded after acquisition and only the final reconstructions are retained.
@@
 The transversal and coronal protocols were considered as two separate datasets, namely LUMC-TRA and LUMC-COR.
-Table I shows an overview of the four sequences.
+Table 1 shows an overview of the four sequences.
 Note that the in-plane resolution of the T1W scans was 1.3-2.5 times as low as that of the T2W scans, making these datasets more challenging for T1W-guided T2W reconstruction.
 As with the NYU dataset, the LUMC datasets were split into model-train, model-val, recon-val, and recon-test.
-The former two splits contained only image-domain data, which was used to train and validate MUNIT.
-The MUNIT models were pre-trained on the full model-train splits, ignoring any pairing between T1W and T2W scans.
+The former two splits contained only image-domain data, which was used to train and validate the content/style models.
+The models were pre-trained on the full model-train splits, ignoring any pairing between T1W and T2W scans.
 Additionally, 20 subjects from the model-train split were designated for PFT where the pairing information was used in training and the reference scans were aligned via registration.
@@
 We further undersampled it retrospectively to higher accelerations $R\in\{4,6,8,10\}$ by dropping subsets of the acquired lines.
-An overview of the data split is shown in Table II.
+An overview of the data split is shown in Table 2.
 Spatially aligned images required by all but the unpaired training set were obtained via rigid registration and the reference T1W images were resampled to the T2W resolution.
@@
 Given the of absence of k-space training data, end-to-end reconstruction methods were not feasible as baselines.
-Hence, we compared PnP-MUNIT with only the feasible types of baselines, i.e.
+Hence, we compared PnP-CoSMo with only the feasible types of baselines, i.e.
 classical, semi-classical, and plug-and-play reconstruction and image-to-image translation.
-Among classical methods, we used the unguided L1-wavelet CS (CS-WT) and the guided STV-based CS (CS-STV) [7].
-As an unguided plug-and-play baseline, we used PnP-CNN [23], and as a representative image translation baseline, we compared against MUNIT itself.
+Among classical methods, we used the unguided L1-wavelet CS (CS-WT) and the guided STV-based CS (CS-STV) (Ehrhardt and Betcke, 2016).
+As an unguided plug-and-play baseline, we used PnP-CNN (Ahmad et al., 2020), and as a representative image translation baseline, we compared against MUNIT itself.
 Using MUNIT, deterministic image translation was approximated by combining the reference content with 200 randomly sampled T2W style codes and taking a pixel-wise mean of the resulting synthetic images to obtain a single synthetic image.
-As semi-classical guided baseline, we used (PROSIT) [39], which combines deterministic image translation with L2-regularized least-squares reconstruction.
-Additionally, as an ablation study for PnP-MUNIT, we ablated PFT and CR to assess their contribution, and finally, as an upper-bound for PnP-MUNIT representing zero content discrepancy, we disabled PFT and CR and used the ideal content $c^{*}$.
-In both NYU and LUMC benchmarks, we used 3 perceptual metrics for evaluation – SSIM, HaarPSI, and DISTS.
-While SSIM is used commonly, HaarPSI and DISTS are known to correlate better with visual judgment of image quality [47].
+As semi-classical guided baseline, we used (PROSIT) (Mattern et al., 2020), which combines deterministic image translation with L2-regularized least-squares reconstruction.
+Additionally, as an ablation study for PnP-CoSMo, we ablated PFT and CR to assess their contribution, and finally, as an upperbound for PnP-CoSMo representing zero content discrepancy, we disabled PFT and CR and used the ideal content $c^{*}$.
+In both NYU and LUMC benchmarks, we used three perceptual metrics for evaluation – SSIM, HaarPSI, and DISTS.
+While SSIM is used commonly, HaarPSI and DISTS are known to correlate better with visual judgment of image quality (Kastryulin et al., 2023).
 All three metrics are bounded in $[0,1]$ where $1$ represents perfect image quality.
 We conducted paired Wilcoxon signed-rank tests to measure statistical significance when comparing pairs of algorithms.
-IV-A4 Radiological evaluation
-Finally, as an extension to the LUMC benchmark, we conducted a radiological evaluation assessing the visual and diagnostic quality of PnP-MUNIT reconstructions.
+4.1.4 Radiological evaluation
+Finally, as an extension to the LUMC benchmark, we conducted a radiological evaluation assessing the visual and diagnostic quality of PnP-CoSMo reconstructions.
 The study was conducted on a small sample of the LUMC-TRA recon-test set.
 With the help of a junior radiologist, we selected 3 cases with brain metastases.
-The evaluation comprised two parts, namely visual quality and pathology.
+The evaluation was performed by a senior neuroradiologist and comprised two sets of criteria, namely visual quality and pathology.
 Four visual quality criteria were used, namely sharpness, noise, artifacts, and contrast between gray and white matter and CSF.
 For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases.
-The images were scored using a five-point Likert scale [48] – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality.
-A diagnostic-quality reconstruction was defined as one that showed at least 90% of the metastases and scored at least “fairly diagnostic quality” on all other criteria.
-[^2] In clinical practice, while the remaining 10% of the metastases may not be detected in the reconstructed T2W scan, they would be prominently visible in the corresponding gadolinium-enhanced T1W scan, and hence would not go undetected.
-We evaluated PnP-MUNIT and PnP-CNN at 4 clinically realistic accelerations of $R\in\{3,4,5,6\}$.
-In total, 9 images per patient, including the (4$\times$2=8) reconstructions and the clinical ground truth (acquired at $R$=2 and reconstructed by the vendor software), were presented to a senior neuroradiologist who scored each image individually, blinded to the reconstruction method and $R$.
-IV-B Implementation Details
-We used the same general residual architecture for MUNIT encoders and decoders as the original paper, except with an additional layer at the end of the content encoder to produce content maps of given number of channels.
+The images were scored using a five-point Likert scale (Mason et al., 2019) – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality.
+A diagnostic-quality reconstruction was defined as one that scored at least ”fairly diagnostic” on the overall diagnostic quality criterion.
+We evaluated PnP-CoSMo and PnP-CNN at 4 clinically realistic accelerations of $R\in\{3,4,5,6\}$.
+In total, 9 images per patient, including the (4$\times$2=8) reconstructions and the clinical ground truth (acquired at $R$=2 and reconstructed by the vendor software), were presented to the senior neuroradiologist who scored each image individually, blinded to the reconstruction method and $R$.
+4.2 Implementation details
+We used the same general residual architecture for MUNIT encoders and decoders as the original paper, except with an additional layer at the end of the content encoder to produce content maps of specifies number of channels.
 We used 2 or 4 channels, depending on the content downsampling factor.
@@
 We additionally conditioned the discriminators on foreground masks of the images to penalize background signal.
-To stabilize GAN training, we used spectral normalization [49].
+To stabilize GAN training, we used spectral normalization (Miyato et al., 2018).
 For hyperparameters, we found that pre-training loss weight values $\alpha_{1}=\alpha_{2}=\alpha_{3}=1$ and paired fine-tuning values $\beta_{1}=\beta_{2}=\beta_{3}=1$ worked generally well for the in vivo datasets.
-In the benchmarks, the CR module parameter $\gamma$ of PnP-MUNIT was tuned per acceleration for each dataset and the number of iterations was set to 200.
-In the simulations, all MUNIT models were pre-trained for 200k and were not fine-tuned for the sake of simplicity, unless stated otherwise in Section V-A.
-In both NYU and LUMC benchmarks, the MUNIT models were pre-trained for 400k iterations and fine-tuned for 50k iterations.
-All software was implemented in PyTorch, and image registration was performed using Elastix [50].
+In the benchmarks, the step size $\gamma$ of the CR module of PnP-CoSMo was tuned per acceleration for each dataset and the number of iterations was set to 200.
+In the simulations, all content/style models were pre-trained for 200k and were not fine-tuned for the sake of simplicity, unless stated otherwise in Section 5.1.
+In both NYU and LUMC benchmarks, the content/style models were pre-trained for 400k iterations and fine-tuned for 50k iterations.
+All software was implemented in PyTorch, and image registration was performed using Elastix (Klein et al., 2009).
 All training runs were performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
-V Results
-TABLE III: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices). PSNR worse than L1-wavelet CS is marked as .
+5 Results
+5.1 Experiments on simulated MR datasets
+5.1.1 Perturbation of disentanglement strength
+PnP-CoSMo assumes that the content and style aspects of the images are sufficiently disentangled.
+Here, we evaluated the effect of perturbing the disentanglement loss weights $\alpha_{2}$ and $\alpha_{3}$ (Eq.
+(6)) of MUNIT on the PnP-CoSMo reconstruction quality.
+Table 3 shows the reconstruction PSNR for different weight values.
+We observed that optimal disentanglement, and thus the best reconstruction, was achieved at $\alpha_{2}=\alpha_{3}=0.1$.
+At lower values, poor disentanglement of content and style in the model resulted in poor PnP-CoSMo reconstruction quality.
+This was expected since PnP-CoSMo relies on the assumption that the content representation is (sufficiently) contrast-independent, which was less enforced at these levels.
+At higher values, the training over-emphasized disentanglement while under-emphasizing the GAN and image recovery loss terms, thereby leading to worse preservation of image information in the latent space and thus, to worse reconstructions.
+Table 3: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices). PSNR worse than the conventional L1-wavelet CS reconstruction is marked as .
 Disentanglement Strength R=2 R=4
@@
 $\alpha_{2}=\alpha_{3}=0.001$ 18.17 $\pm$ 0.07 15.88 $\pm$ 0.05
-TABLE IV: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices) at $R$=4. Optimal content capacity is shown in bold. PSNR worse than L1-wavelet CS is marked as .
+Table 4: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices) at $R$=4.
+Optimal content capacity is shown in bold.
+PSNR worse than the conventional L1-wavelet CS reconstruction is marked as .
 Model Config Data Config
-RefRes-1 RefRes-2 RefRes-4 RefRes-8
-ContentRes-1 27.90$\pm$0.08 21.17$\pm$0.05 19.11$\pm$0.07 16.43$\pm$0.07
-ContentRes-2 26.19$\pm$0.08 22.80$\pm$0.08 20.05$\pm$0.08 16.57$\pm$0.08
-ContentRes-4 23.98$\pm$0.08 20.88$\pm$0.10 19.44$\pm$0.09 16.88$\pm$0.06
-Figure 3: (a) Convergence of PnP-MUNIT and the baseline versions at $R$=4. (b) Evolution of the reconstruction shown with its error map.
-Figure 4: Effectiveness of the CR module in resolving contrast-specific structure, which, in this case, was a simulated lesion.
-V-A Experiments on Simulated MR Datasets
-V-A1 Perturbation of disentanglement strength
-PnP-MUNIT relies heavily on a sufficient disentanglement of content and style from the images.
-Here, we evaluated the effect of perturbing the disentanglement loss weights $\alpha_{2}$ and $\alpha_{3}$ (Eq.
-(6)) of MUNIT on the PnP-MUNIT reconstruction quality.
-Table III shows the reconstruction PSNR for different weight values.
-We observed that optimal disentanglement, and thus the best reconstruction, was achieved at $\alpha_{2}=\alpha_{3}=0.1$.
-At lower values, poor disentanglement of content and style in the model resulted in poor PnP-MUNIT reconstruction quality.
-This was expected since PnP-MUNIT relies on the assumption that the content representation is (sufficiently) contrast-independent, which was less enforced at these levels.
-At higher values, the training over-emphasized disentanglement while under-emphasizing the GAN and image recovery loss terms, thereby leading to worse preservation of image information in the latent space and thus, to worse reconstructions.
-V-A2 Content resolution analysis
-PnP-MUNIT relies on a reference image of sufficiently high resolution to provide guidance to the reconstruction.
-The goal of this experiment was to determine the effect of lowering the reference image resolution on the reconstruction quality and using the content/style model’s content capacity (i.e.
-the spatial resolution of the content relative to the images) to explain this effect.
-We simulated 4 datasets, denoted as RefRes-$n$, where $n\in\{1,2,4,8\}$ represents the reference domain downsampling factor.
-In the $n$=1 case, T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower $1/n$ frequency components, while maintaining the same spatial resolution.
-For each dataset, we trained 3 content/style models with content capacity denoted as ContentRes-$m$, where $m\in\{1,2,4\}$ is the content downsampling factor, which depends on the number of downsampling blocks in the networks.
+RefRes-1 RefRes-1/2 RefRes-1/4 RefRes-1/8
+$J_{M}$=1 27.90$\pm$0.08 21.17$\pm$0.05 19.11$\pm$0.07 16.43$\pm$0.07
+$J_{M}$=(1/2)2 26.19$\pm$0.08 22.80$\pm$0.08 20.05$\pm$0.08 16.57$\pm$0.08
+$J_{M}$=(1/4)2 23.98$\pm$0.08 20.88$\pm$0.10 19.44$\pm$0.09 16.88$\pm$0.06
+5.1.2 Content capacity analysis
+PnP-CoSMo relies on a reference image of sufficiently high resolution to provide guidance to the reconstruction.
+The goal of this experiment was to determine the effect of lowering the reference image resolution on the reconstruction quality and using the content/style model’s content capacity (Eq.
+(14)) to explain this effect.
+We simulated 4 datasets, denoted as RefRes-$1/n$, where $n\in\{1,2,4,8\}$ represents the reference domain downsampling factor.
+In the $n$=1 case, the T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower $1/n$ frequency components, while maintaining the same spatial resolution.
+For each dataset, we trained 3 content/style models with content capacity $J_{M}$=$(1/m)^{2}$, where $m\in\{1,2,4\}$ is the content downsampling factor, which depends on the number of up/downsampling blocks in the networks.
 E.g.
-ContentRes-2 model produced content maps half the spatial resolution of the ContentRes-1 model.
-Figure 5: The effect of misaligned reference on the reconstruction and the robustness of the CR module to it.
-Figure 6: Reconstruction PSNR as a function of acceleration $R$ and k-space noise level $\sigma$ for different values of the $\gamma$ parameter.
-$\gamma$ controls the consistency of the content with the measured k-space.
-Figure 7: Evaluation plots over recon-test subset of NYU DICOM dataset consisting of 550 slices.
-Figure 8: Two examples from the NYU benchmark.
-(a) Comparison of all the methods at the highest acceleration of $R$=5.
-(b) Comparison between PnP-MUNIT and MC-VarNet at $R$=2 and $R$=4 on a pathological case where many features of the edema are T2W contrast-specific.
-Table IV compares PnP-MUNIT reconstruction quality across these configurations.
+the $J_{M}$=(1/2)2 model produced content maps half the spatial resolution of the $J_{M}$=1 model.
+Figure 4: (a) Convergence of PnP-CoSMo and its two variants at $R$=2. (b) Evolution of the reconstruction shown with the corresponding error map.
+Table 4 compares PnP-CoSMo reconstruction quality across these configurations.
 PFT and CR were disabled here for simplicity.
 We observe two trends.
-First, the reconstruction quality generally decreases from left to right, eventually dropping below L1-wavelet CS reconstruction, suggesting a decrease in the amount of shared local information contained in the reference contrast.
+First, the reconstruction quality generally decreases from left to right, eventually dropping below the conventional L1-wavelet CS reconstruction, suggesting a decrease in the amount of shared local information contained in the reference contrast.
 Second, the optimal content capacity of the model decreases with the reference resolution in accordance with the actual amount of this shared information.
 In other words, this optimal content capacity corresponds to the amount of contrast-independent structure the model discovers in the dataset, e.g.
-RefRes-1 dataset had 4 times as much shared structure (due to twice as large optimal content resolution) as RefRes-2 and RefRes-4 datasets and hence, the RefRes-1 dataset was significantly more effective in guided reconstruction.
-On the other hand, Refsize-2 and RefRes-4 datasets had similar levels of shared structure despite the lower reference resolution in the latter, suggesting a more complex relationship between the reference resolution and the shared content.
-In the RefRes-8 case, PnP-MUNIT dropped to a similar level of quality as conventional CS, suggesting a lowerbound on the amount of shared information for PnP-MUNIT to be effective, specifically 1/42=1/16 times the local information of the full-resolution case.
-V-A3 Convergence
-We explored the convergence of PnP-MUNIT by comparing it with two variants.
-The first used true content $c^{*}$ of the ground truth T2W image, thus assuming zero content discrepancy and representing an upper-bound of PnP-MUNIT.
-The second used reference content $\hat{c}$ with CR step disabled, representing a lower-bound of PnP-MUNIT where the non-zero content discrepancy is not corrected.
-Here and in the following simulation experiments, we used the RefRes-1 dataset and the ContentRes-1 model additionally fine-tuned on 2 training volumes for 50k iterations.
+RefRes-1 dataset had 4 times as much shared structure (due to twice as large optimal content resolution) as RefRes-1/2 and RefRes-1/4 datasets and hence, the RefRes-1 dataset was significantly more effective in guided reconstruction.
+On the other hand, RefRes-1/2 and RefRes-1/4 datasets had similar levels of shared structure (i.e.
+$J_{M}$=1/22=1/4 times the full-resolition case) despite the lower reference resolution in RefRes-1/4, suggesting a more complex relationship between the reference resolution and the shared content.
+In the RefRes-1/8 case, PnP-CoSMo dropped to a similar level of quality as conventional CS, suggesting a lowerbound on the amount of shared information for PnP-CoSMo to be effective, specifically $J_{M}$=1/42=1/16 times the local information of the full-resolution case.
+5.1.3 Convergence
+We explored the convergence of PnP-CoSMo by comparing it with two variants.
+The first used true content $c^{*}$ of the ground truth T2W image, thus assuming zero content discrepancy and representing an upperbound of PnP-CoSMo.
+The second used reference content $\hat{c}$ with CR module disabled, representing a lowerbound of PnP-CoSMo where the non-zero content discrepancy is not corrected.
+Here and in the following simulation experiments, we used the RefRes-1 dataset and the $J_{M}$=1 model which was additionally fine-tuned on 2 training volumes for 50k iterations.
 Fig.
 4 shows the convergence curves and intermediate reconstructions.
-The upper-bound version, given $c^{*}$, converged in a single iteration, while with $\hat{c}$, the ablated version converged equally fast but to a sub-optimal solution.
-Enabling the CR step closed the gap with the upper-bound, although costing convergence rate.
-V-A4 Effectiveness in resolving contrast-specific structure
+The upperbound version, given $c^{*}$, converged in a single iteration, while with $\hat{c}$, the ablated version converged equally fast but to a sub-optimal solution.
+Enabling the CR module closed the gap with the upperbound, although at the cost of slower convergence.
+It is also observed that the content discrepancy reduced when using CR and that the style estimate always converged within the first few iterations, as we hypothesized in Section 3.3.
+5.1.4 Effectiveness in resolving contrast-specific structure
+Figure 5: Effectiveness of the CR module in resolving contrast-specific structure, which, in this case, was a simulated lesion.
 In order to test the effectiveness of the CR module in resolving structure present in the target contrast but absent in the reference contrast, we simulated a lesion in the T2W image and performed reconstruction with and without the CR module.
 As shown in Fig.
-4, the CR module was able to recover the lesion fully at $R$=2 and substantially at $R$=4.
-V-A5 Sensitivity to misalignment of the reference
-While PnP-MUNIT expects an aligned reference image at reconstruction time, the CR module should, in principle, correct for small misalignments.
-To test this, we simulated a 2 rotation in the reference image and performed reconstruction with and without the CR module, comparing also with PnP-MUNIT that used an aligned reference image.
+5, the CR module was able to recover the lesion fully at $R$=2 and substantially at $R$=4.
+5.1.5 Sensitivity to misalignment of the reference
+Figure 6: A misaligned reference image severely affects the reconstruction when CR is disabled. Enabling CR makes the reconstruction process robust to small misalignments.
+While PnP-CoSMo expects an aligned reference image at reconstruction time, the CR module should, in principle, correct for small misalignments.
+To test this, we simulated a 2 rotation in the reference image and performed reconstruction at $R$=4 with and without the CR module, comparing also with PnP-CoSMo that used an aligned reference image.
 As shown in Fig.
-5, although the reconstruction was sensitive to the misalignment in the absence of CR, it significantly improved with CR enabled.
-Figure 9: Evaluation plots over recon-test subset of LUMC-TRA and LUMC-COR datasets consisting of 1366 and 200 slices, respectively.
-Figure 10: Sample slice from the LUMC-COR recon-test set comparing PnP-MUNIT with PnP-CNN and MUNIT image translation.
-V-A6 Tuning the CR step size
-The hyperparameter $\gamma$ in Algorithm 1 controls the CR strength.
+6, although the reconstruction was sensitive to the misalignment in the absence of CR, it significantly improved with CR enabled.
+5.1.6 Influence of the CR step size
+Figure 7: Reconstruction PSNR as a function of acceleration $R$ and k-space noise level $\sigma$ for different values of the hyperparameter $\gamma$.
+$\gamma$ is the CR step size controlling the consistency of the content with the measured k-space.
+The hyperparameter $\gamma$ in Algorithm 1 controls the CR step size.
 Fig.
@@
 This makes sense as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
-V-B Benchmark on NYU DICOM Data
+5.2 Benchmark on the NYU DICOM data
+In the NYU benchmarking setup considered here, let the T1W-guided T2W reconstruction task be denoted as Task-1 and the T2W-guided T1W reconstruction task as Task-2.
 Fig.
-7 shows evaluation metrics for PnP-MUNIT and the end-to-end baselines on the NYU DICOM dataset.
-Despite being trained only on image-domain data, PnP-MUNIT largely outperformed MoDL ($p<0.05$ throughout, except SSIM at $R$=2 and DISTS at $R\in${2,3,4}) and MTrans ($p<$0.05), and was roughly comparable to MC-VarNet.
-Moreover, as shown in Fig.
-8a, PnP-MUNIT could resolve certain fine details even at the high acceleration of $R$=5, which the baselines failed at.
+8 shows evaluation metrics for PnP-CoSMo and the end-to-end baselines on these two tasks.
+We observe three main trends.
+First, the end-to-end methods generally dropped in performance in the out-of-distribution cases ($p$$<$0.05 throughout, with the exception of SSIM at $R$=5 and HaarPsi at $R\in${2,4,5} for MoDL on Task-2).
+Second, PnP-CoSMo outperformed out-of-distribution models ($p$$<$0.05, except DISTS of MoDL and MCVarNet at $R$=2 on Task-2).
+Since PnP-CoSMo does not depend on any task-specific k-space training data owing to both its plug-and-play design and because of the bi-directionality of the content/style model it uses, it is not susceptible to the distribution shift of this kind.
+Third, considering the in-distribution model performance, PnP-CoSMo largely outperformed MoDL ($p$$<$0.05 throughout, except SSIM at $R$=2 and DISTS at $R\in${2,3,4} on Task-1 and DISTS at $R$=2 on Task-2) and MTrans ($p$$<$0.05 throughout, except SSIM at $R\in${3,4} and DISTS at $R\in${2,3}), and was roughly comparable to MC-VarNet, despite having lenient training data requirements.
+As shown in Fig.
+9a, PnP-CoSMo could resolve certain fine details even at the high acceleration of $R$=5, which the baselines failed at.
 In the pathological case shown in Fig.
-8b, PnP-MUNIT produced sharper reconstructions preserving the structure of the edema better than MC-VarNet despite lower metrics when many features of edema were absent in the reference image.
-V-C Benchmark and Ablation on LUMC Multi-coil Data
+9b, PnP-CoSMo produced sharper reconstructions while preserving the structure of the edema better than (the in-distribution) MC-VarNet, despite lower metrics, when many features of edema were absent in the reference image.
+Figure 8: Evaluation plots over recon-test subset of NYU DICOM dataset consisting of 550 slices for the two reconstruction tasks.
+ID and OOD refer to in-distribution and out-of-distribution models, respectively.
+Note that MoDL is a single-contrast baseline and hence did not use any reference images.
+Unlike other methods, PnP-CoSMo, by design, required only a single content/style model for both tasks.
+Figure 9: Two examples from the NYU benchmark for Task-1 (i.e.
+T1W-guided T2W reconstruction).
+(a) Comparison of PnP-CoSMo with the in-distribution end-to-end models at the highest acceleration of $R$=5.
+(b) Comparison between PnP-CoSMo and MC-VarNet at $R$=2 and $R$=4 on a pathological case where many features of the edema are T2W contrast-specific.
+5.3 Benchmark and ablation on the LUMC multi-coil data
+Figure 10: Evaluation plots over recon-test subset of LUMC-TRA and LUMC-COR datasets consisting of 1366 and 200 slices, respectively.
+Figure 11: Sample slice from the LUMC-COR recon-test set comparing PnP-CoSMo with PnP-CNN and MUNIT image translation.
 The LUMC datasets represent a more challenging problem with real-world data constraints.
-Additionally, in the light of the content resolution analysis (Section V-A2), we empirically found the ContentRes-2 model configuration as optimal for the LUMC datasets, compared to NYU DICOM data for which ContentRes-1 was the optimal model configuration.
-This reflects the lower effective content in LUMC image data and hence a greater difficulty for guided reconstruction.
+Additionally, in the light of the content capacity analysis (Section 5.1.2), we empirically found that the content capacity $J_{M}$=1/4 was optimal for the LUMC datasets, compared to NYU DICOM data where $J_{M}$=1 was the optimal content capacity.
+This reflects the lower effective content in LUMC image data and hence a greater difficulty of the guided reconstruction problem.
+Figure 12: Scores from the radiological evaluation averaged over the three subjects.
+Figure 13: Example slice from the LUMC-TRA subset used in the radiological evaluation.
+Several relevant features are marked including a brain metastasis-related hyperintense feature (salmon-pink arrows), reconstruction artifacts (cyan arrows), and blurring (purple arrows).
 Fig.
@@
 Combining it with L2-regularized reconstruction in PROSIT improved SSIM and HaarPSI ($p$$<$0.05 for both datasets and accelerations), but not necessarily DISTS, suggesting that the the available complementary information was not fully utilized.
-PnP-MUNIT was consistently better than both MUNIT image translation and PROSIT ($p$$<$0.05 throughout for both cases).
+PnP-CoSMo was consistently better than both MUNIT image translation and PROSIT ($p$$<$0.05 throughout for both cases).
 It also outperformed the conventional guided CS-STV ($p$$<$0.05 throughout, except SSIM at $R$=4 in LUMC-COR).
-Hence, using both measured k-space and reference scan via PnP-MUNIT was more beneficial than (a) using either one of them and (b) combining both using hand-crafted priors, suggesting that our approach maximally exploits the complementary information.
-Second, both PnP-CNN and PnP-MUNIT performed similarly at lower acceleration, except in terms of SSIM and HaarPSI at $R$=4 and $R$=6 on LUMC-COR where PnP-CNN was slightly better.
-At higher acceleration, PnP-MUNIT outperformed PnP-CNN ($p$$<$0.05 for all metrics and both datasets except SSIM at $R$=8 in LUMC-COR).
-Compared to PnP-CNN, PnP-MUNIT allowed up to 32.6% more acceleration for a given SSIM.
-[^3] Based on linear interpolation of LUMC-TRA SSIM, PnP-MUNIT and PnP-CNN allowed $R$=10 and $R$=6.7, the maximum difference in $R$, at median SSIM=0.906.
+Hence, using both measured k-space and reference scan via PnP-CoSMo was more beneficial than (a) using either one of them and (b) combining both using hand-crafted priors, suggesting that our approach maximally exploits the complementary information.
+Second, both PnP-CNN and PnP-CoSMo performed similarly at lower acceleration, except in terms of SSIM and HaarPSI at $R$=4 and $R$=6 on LUMC-COR where PnP-CNN was slightly better.
+At higher acceleration, PnP-CoSMo outperformed PnP-CNN ($p$$<$0.05 for all metrics and both datasets except SSIM at $R$=8 in LUMC-COR).
+Compared to PnP-CNN, PnP-CoSMo allowed up to 32.6% more acceleration for a given SSIM.
+[^2] Based on linear interpolation of LUMC-TRA SSIM, PnP-CoSMo and PnP-CNN allowed $R$=10 and $R$=6.7, which is the maximum difference in $R$, at median SSIM=0.906.
 Hence, a difference of 32.6%.
-Third, introducing PFT and CR into the ablated PnP-MUNIT improved the reconstructions ($p$$<$0.05 throughout in both cases except in DISTS for the latter at $R$=10 in LUMC-TRA).
-The contribution of PFT was almost constant across $R$, whereas that of CR decreased with $R$, which was expected since CR depends on the measured k-space data to refine the content.
+Third, introducing PFT and CR into the ablated PnP-CoSMo improved the reconstructions ($p$$<$0.05 throughout in both cases except in DISTS for the latter at $R$=10 in LUMC-TRA).
+The contribution of PFT was almost constant across $R$, whereas that of CR decreased as $R$ increased, which was expected since CR depends on the measured k-space data to refine the content.
 Fig.
@@
 MUNIT image translation produced severe anatomical defects and predicted false structure in the lower region of the image where the ground truth, in fact, contains low signal.
-At $R$=4, PnP-MUNIT was visually similar to PnP-CNN, with the main difference being a blur effect in PnP-CNN and mild texture artifacts in PnP-MUNIT.
+At $R$=4, PnP-CoSMo was visually similar to PnP-CNN, with the main difference being a blur effect in PnP-CNN and mild texture artifacts in PnP-CoSMo.
 At $R$=10, PnP-CNN reconstruction was non-viable, containing severe blur and artifacts (green arrows).
-On the other hand, PnP-MUNIT did not degrade much from the $R$=4 case and most fine structures were sharply resolved (green arrows).
-However, an interesting failure mode of PnP-MUNIT at high acceleration was the subtle localized distortions in the anatomy (pink arrows).
-Figure 11: Averaged results of the radiological evaluation.
-V-D Radiological Evaluation
+On the other hand, PnP-CoSMo did not degrade much from the $R$=4 case and most fine structures were sharply resolved (green arrows).
+However, an interesting failure mode of PnP-CoSMo at high acceleration was the subtle localized distortions in the anatomy (magenta arrows).
+5.4 Radiological evaluation
 Figure 12 plots the radiological evaluation result.
-At $R$=3, both algorithms produced at least “fairly diagnostic” reconstructions and allowed the detection of more than 90% of metastases, a difference of 33.3% of k-space samples over the clinical reconstructions of $R$=2.
-PnP-CNN was better than PnP-MUNIT at $R\in\{3,4\}$ in terms of sharpness, contrast, and the pathology criteria.
-At $R$=5, PnP-MUNIT matched or exceeded PnP-CNN in terms of visual quality, although being worse in terms of pathology criteria.
-At $R$=6, PnP-CNN sharply dropped to non-diagnostic level, whereas PnP-MUNIT was better, especially in terms of sharpness and contrast as well as in overall diagnostic quality.
-VI Discussion
-In this work, we modeled the two-contrast MR image data in terms of two latent generative factors – content representing the contrast-independent structure and style representing the contrast-specific variations.
+At $R$=3, both algorithms produced diagnostic-quality reconstructions, a difference of 33.3% of k-space samples over the clinical reconstructions of $R$=2.
+PnP-CNN was slightly better than PnP-CoSMo at $R\in\{3,4\}$ in terms of sharpness, contrast, and the pathology criteria.
+At $R$=5, PnP-CoSMo matched or exceeded PnP-CNN in terms of visual quality, although being worse in terms of pathology criteria.
+At $R$=6, PnP-CNN sharply dropped to the ”non-diagnostic” level, whereas PnP-CoSMo was better, especially in terms of sharpness and contrast as well as in overall diagnostic quality.
+This trend is illustrated with a representative example in Figure 13.
+While the $R$=3 reconstructions of PnP-CNN and PnP-CoSMo are of a similarly high quality, reconstruction artifacts (cyan arrows) become more pronounced in PnP-CNN reconstructions with increasing acceleration.
+At $R$=6, compared to PnP-CoSMo, the PnP-CNN reconstruction has severe artifacts, blurrier and more distorted features (purple arrows), and a poorly visible hyperintense area of a brain metastasis (salmon-pink arrows).
+6 Discussion
+In this work, we modeled two-contrast MR image data in terms of two latent generative factors – content representing the contrast-independent structure and style representing the contrast-specific variations.
 Our data-driven definitions of content and style, though seemingly related to the MR physics-based representations of quantitative maps and acquisition-related factors, are rather nebulous compared to precisely defined physical concepts.
@@
 by constraining the learned content to represent physically meaningful anatomical properties, introducing elements of a physical model (e.g.
-Bloch equation) into the decoder network, etc., to enhance interpretability and reconstruction quality.
+Bloch equations) into the decoder network, etc., to enhance interpretability and reconstruction quality.
 Optimal content capacity was defined as a quantity representing the amount of shared local information contained in a two-contrast dataset.
-In terms of this quantity, we analyzed the effect of reference image resolution on reconstruction quality, obtaining a lowerbound for PnP-MUNIT to be effective (Section V-A2).
+In terms of this quantity, we analyzed the effect of reference image resolution on reconstruction quality, obtaining a lowerbound for PnP-CoSMo to be effective (Section 5.1.2).
 This quantity would depend on more fundamental factors such as MR sequence types, which we did not investigate here.
-While we limited our experiments to T1W and T2W sequences, PnP-MUNIT is applicable to any pair of contrasts, with the image quality limited mainly by the amount of structural information shared between them.
+While we limited our experiments to T1W and T2W sequences, PnP-CoSMo is, in principle, applicable to any pair of contrasts, with the image quality limited mainly by the amount of structural information shared between them.
 Future work could help empirically determine which sequence pairs are more amenable to guided reconstruction than others.
-Note that for a different pair of contrasts, say PD-weighted and T2W, a fresh content/style model needs to be trained since content and style are defined in the context of the specific contrast pair.
+A consequence of our MUNIT-based model design was the model’s lack of dependence on the uni-directionality of the guided reconstruction problem, which allowed greater generalizability by design (Section 5.2).
 It is, in principle, also possible to model the content and style for more than two contrasts.
-Given number of contrasts $N$, content is defined as the local structure underlying all $N$ contrasts, while style encodes intra-contrast variations of each individual contrast.
-This model would include $N$ sets of encoders and decoders and would require an (unpaired) image-domain training dataset of these $N$ contrasts.
-Subsequently, our PnP-MUNIT framework can be extended to use multiple reference contrasts to guide a given target contrast.
+This would enable not only the generalizability of our approach to any combinatorial set of guided reconstruction tasks, but also better estimation of the target contrast’s content by utilizing multiple reference contrasts in the reconstruction.
+Given number of contrasts $N$, content could be defined as the local structure underlying all these contrasts, while style would encode intra-contrast variations of each individual contrast.
+This model would then require an (unpaired) image-domain training dataset of the $N$ contrasts.
+An open question is how to effectively scale the model and training complexity to $N$ contrasts.
+The PnP-CoSMo algorithm can subsequently be extended to utilize multiple reference contrasts to guide a given target contrast.
 The specific question to investigate then would be – given $P$ reference contrasts (where $1<P<N$), how to aggregate their contents such that the prior structural information about the target image is maximized?
-Content discrepancy was characterized as a quantity representing the gap between the true content of the target image and our estimation of this content, which is another limiting factor in PnP-MUNIT reconstruction.
-In MUNIT, the decoders model the (forward) generative process that maps the latent variables (i.e.
+Content discrepancy was characterized as a quantity representing the gap between the true content of the target image and our estimation of this content, which is another limiting factor in PnP-CoSMo reconstruction.
+In our MUNIT-based network design, the decoders model the (forward) generative process that maps the latent variables (i.e.
 content and style) to the observable variable (i.e.
@@
 While paired fine-tuning improves the model using a small amount of paired image-domain training data, content refinement (CR) aims to correct the remaining discrepancy during reconstruction using the measured k-space.
-A drawback of the CR module, however, is slower convergence (Section V-A3).
-For example, the run-time for PnP-MUNIT on the GPU was around 17 seconds for a slice of matrix size 349$\times$284, as compared to 6 seconds for CS-WT, 1.4 seconds for PnP-CNN, and less than 0.1 second for MC-VarNet.
-This latency issue could be mitigated by approximating the iterative process with an unrolled network design, although perhaps at a cost of the generalizability offered by the plug-and-play design.
+A drawback of the CR module, however, is slower convergence (Section 5.1.3).
+For example, the run-time for PnP-CoSMo on the GPU was around 17 seconds for a slice of matrix size 349$\times$284, as compared to 6 seconds for CS-WT, 1.4 seconds for PnP-CNN, and less than 0.1 second for MC-VarNet.
+This latency issue could be mitigated by approximating the iterative process with an unrolled network design, although potentially at a cost of the generalizability offered by the plug-and-play design.
 The resulting trade-off between model latency and generalizability would be interesting to explore in the future.
-In terms of training data requirements, PnP-MUNIT can make use of the larger amounts of the image data available and can be applied to situations where end-to-end methods are infeasible since it does not rely on k-space training data.
-Moreover, unpaired image-domain pre-training of the content/style model boosts the practical applicability of PnP-MUNIT, e.g.
+In terms of training data requirements, PnP-CoSMo can make use of the larger amounts of the image data available and can be applied to situations where end-to-end methods are infeasible since it does not rely on k-space training data.
+In situations where end-to-end methods are indeed feasible, PnP-CoSMo is comparable to state-of-the-art methods in terms of metrics and superior in terms of visual quality, especially at higher accelerations (Section 5.2 ).
+Moreover, unpaired image-domain pre-training of the content/style model boosts the practical applicability of PnP-CoSMo, e.g.
 in the case of LUMC-TRA where the T1W/T2W data imbalance was considerable.
-As indicated in our NYU DICOM benchmark, PnP-MUNIT performs similarly or better than state-of-the-art end-to-end methods.
+As indicated in our NYU DICOM benchmark, PnP-CoSMo performs similarly or better than state-of-the-art end-to-end methods.
 In our LUMC multi-coil benchmark, we limited our baselines to those algorithms that were feasible given the data constraint.
@@
 Contrary to the training stage, we strictly assumed the spatial alignment of the reference and target images in the reconstruction stage.
-The CR update, which can implicitly correct minor registration errors, would likely break down at the typically observed levels of patient motion.
+The CR module, which can implicitly correct minor registration errors, would likely break down at the typically observed levels of patient motion.
 A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further improving practical applicability.
-In terms of reconstruction quality, we observed in Section V-C that the true added value of PnP-MUNIT was at the high acceleration factors ($R$=8 and $R$=10) where the k-space data is scarce and the reference information becomes more valuable.
+In terms of reconstruction quality, we observed in Section 5.3 that the true added value of PnP-CoSMo was at the high acceleration factors ($R$=8 and $R$=10) where the k-space data is scarce and the reference information becomes more valuable.
 However, the risk of model hallucinations increases at these accelerations, raising concerns about the accuracy of the anatomy represented in the image.
-Given the advantages of content/style decomposition, it may be possible to leverage the contrast-independence property of the content to automatically detect local hallucinations in the reconstruction and to define a corrective process to minimize it.
-This is another topic for future work.
-At lower accelerations, on the other hand, we observed that PnP-MUNIT often produced slightly lower metrics compared to PnP-CNN, e.g.
+We hypothesize that as a consequence of content/style decomposition, it may be possible to leverage the contrast-independence property of the content to – (a) automatically detect local hallucinations in the reconstruction and (b) define a corrective process in combination with an adaptive sampling strategy to minimize it.
+This is yet another topic for future work.
+At lower accelerations, on the other hand, we observed that PnP-CoSMo often produced slightly lower metrics compared to PnP-CNN, e.g.
 in LUMC-COR.
 This was also observed in the radiological evaluation on LUMC-TRA samples.
-This is counter-intuitive at first glance since PnP-MUNIT has access to additional side-information and should, in principle, perform at least as good as PnP-CNN.
-However, this can be explained by the fact that at lower accelerations, the content/style model becomes the limiting factor for reconstruction quality.
+This is counter-intuitive at first glance since PnP-CoSMo has access to additional side information and should, in principle, perform at least as well as PnP-CNN.
+However, this can be explained by the fact that at lower accelerations, the content/style model becomes the main limiting factor for reconstruction quality.
 This is because unlike the CNN denoiser (Eq.
 (4)) of PnP-CNN, the content consistency operator (Eq.
-(14)) of PnP-MUNIT radically changes the image and hence, this operation is sensitive to the model’s overall performance.
+(15)) of PnP-CoSMo radically alters the image and hence, this operation is sensitive to the model’s overall performance.
 The model performance is in turn influenced by the training dataset and the network architecture.
@@
 11 at $R$=4.
-In our proof-of-concept, we used the same general model architecture proposed in [22].
+In our proof-of-concept, we used the same general model architecture proposed in Huang et al.
+(2018).
 Better designs that are more conducive to CR may exist.
 Hence, future efforts must considerably focus on improving the content/style model using specialized architecture and regularization.
-VII Conclusion
-In this work, we introduced PnP-MUNIT, a modular approach to multi-contrast reconstruction combining content/style modeling with iterative reconstruction, which offers the reconstruction quality of end-to-end methods under stronger training data constraints.
-At its core is the content consistency operation, which provides regularization at the level of the image’s semantic content.
+7 Conclusion
+In this work, we introduced PnP-CoSMo, a modular approach to multi-contrast reconstruction which combines content/style modeling with iterative reconstruction, offering the reconstruction quality of state-of-the-art end-to-end methods with improved generalizability despite lenient training data requirements.
+At its core is the content consistency operation, which provides regularization at the semantic level of the image content.
 We defined two quantities that determine the efficacy of this operation, namely optimal content capacity and content discrepancy, and provided several ways of maximizing this efficacy.
-On real-world clinical data, PnP-MUNIT provided up to 32.6% more acceleration over PnP-CNN for given SSIM, enabling sharper reconstructions at high accelerations.
-In the radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-MUNIT produced diagnostic-quality images at $R$=3, enabling 33.3% more acceleration over clinical reconstructions.
-To progress towards practical implementation of our proof-of-concept, future work will focus mainly on improving the content/style model and reconstruction latency, tackling the problem of model hallucinations, and incorporating online registration into the reconstruction.
+On real-world clinical data, PnP-CoSMo provided up to 32.6% more acceleration over PnP-CNN for given SSIM, enabling sharper reconstructions at high accelerations.
+In the radiological task of visual quality assessment and brain metastasis diagnosis at realistic accelerations, PnP-CoSMo produced diagnostic-quality images at $R$=3, enabling 33.3% more acceleration over clinical reconstructions.
+To progress towards practical implementation of our proof-of-concept, future work will focus mainly on improving and extending the content/style model, reducing reconstruction latency, tackling the problem of model hallucinations, and incorporating online registration into the reconstruction.
+CRediT authorship contribution statement
+Chinmay Rao: Conceptualization, Methodology, Data curation, Software, Investigation, Validation, Formal analysis, Visualization, Writing – original draft.
+Matthias van Osch: Funding acquisition, Supervision, Conceptualization, Validation, Writing – review & editing.
+Nicola Pezzotti: Funding acquisition, Supervision, Conceptualization, Methodology, Validation.
+Jeroen de Bresser: Investigation, Validation.
+Laurens Beljaards: Data curation, Software.
+Jakob Meineke: Resources, Software.
+Elwin de Weerdt: Conceptualization, Resources.
+Huangling Lu: Investigation.
+Mariya Doneva: Supervision, Conceptualization, Methodology, Validation, Writing – review & editing.
+Marius Staring: Funding acquisition, Project administration, Supervision, Conceptualization, Methodology, Validation, Writing – review & editing.
+Declaration of competing interest
+The authors declare the following financial interests or personal relationships which may be considered as potential competing interests: Chinmay Rao and Laurens Beljaards declare the grant and research support they receive received from Philips.
+Nicola Pezzotti, Jakob Meineke, Elwin de Weerdt, and Mariya Doneva declare their employment at Philips.
+Given his role as Associate Editor, Marius Staring had no involvement in the peer review of this article and had no access to information regarding its peer review.
+Full responsibility for the editorial process for this article was delegated to another journal editor.
+If there are other authors, they declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Acknowledgments
+This work is part of the project ROBUST: Trustworthy AI-based Systems for Sustainable Growth with project number KICH3.LTP.20.006, which is (partly) financed by the Dutch Research Council (NWO), Philips Research, and the Dutch Ministry of Economic Affairs and Climate Policy (EZK) under the program LTP KIC 2020-2023.
+Data availability
+The authors do not have permission to share data.